### v1

Abstract
Trustworthiness is essential for the precise and interpretable application of artificial intelligence (AI) in medical imaging. Traditionally, precision and interpretability have been addressed as separate tasks, namely medical image analysis and explainable AI, each developing its own models independently. In this study, for the first time, we investigate the development of a unified medical vision-language pre-training model that can achieve both accurate analysis and interpretable understanding of medical images across various modalities. To build the model, we construct MedConcept-23M, a large-scale dataset comprising 23 million medical image-text pairs extracted from 6.2 million scientific articles, enriched with concepts from the Unified Medical Language System (UMLS). Based on MedConcept-23M, we introduce ConceptCLIP, a medical AI model utilizing concept-enhanced contrastive language-image pre-training. The pre-training of ConceptCLIP involves two primary components: image-text alignment learning (IT-Align) and patch-concept alignment learning (PC-Align). Specifically, IT-Align facilitates the global alignment of medical image and text representations, while PC-Align uses UMLS knowledge for detailed alignment between image patches and their corresponding conceptual representations. This dual alignment strategy enhances the model’s capability to associate specific image regions with relevant concepts, thereby improving both the precision of analysis and the interpretability of the AI system. We conducted extensive experiments on 5 diverse types of medical image analysis tasks, spanning 51 subtasks across 10 image modalities, with the broadest range of downstream tasks. The results demonstrate the effectiveness of the proposed vision-language pre-training model. Further explainability analysis across 6 modalities reveals that ConceptCLIP achieves superior performance, underscoring its robust ability to advance explainable AI in medical imaging. These findings highlight ConceptCLIP’s capability in promoting trustworthy AI in the field of medicine.
keywords:
Artificial Intelligence, Trustworthiness, Medical Image Analysis, Explainable AI, Image-Language Pre-training, Concept Enhancement
1 Introduction
Trustworthiness in medical image analysis refers to the accuracy and interpretability of AI systems, which are essential for the effective deployment in clinical settings [1]. Ensuring that AI systems deliver precise and interpretable results is crucial for reliable clinical decision-making. Traditionally, the development of trustworthy medical AI systems involves two distinct components: a medical image analysis module [2] and an explainable AI module [3]. The medical image analysis module is designed to produce outputs that meet clinical requirements, such as providing accurate classification labels for diagnosis or generating coherent textual outputs for report generation and visual question answering. Meanwhile, the explainable AI module audits and elucidates the decision-making process of medical image analysis, offering insights into the model’s decision. Despite advancements in these areas, there remains a gap in achieving both tasks using a unified system.
In this study, we propose that a single medical vision-language pre-training model can effectively address both analysis and explanation tasks across various modalities. This proposition is substantiated by two key aspects. First, vision-language models pre-trained with large-scale image-text pairs have demonstrated promising performance across a wide spectrum of image analysis tasks in both general [4, 5, 6] and medical domains [7, 8]. Second, these models exhibit robust zero-shot classification capabilities [4], which provide a robust foundation for zero-shot concept annotation. As stated in  [3], stronger zero-shot concept annotation ability could result in higher-quality explanations within AI systems. Thus, this capability serves as a bridge, motivating the integration of analysis and explanation modules in a unified framework.
Figure 1: Overview of ConceptCLIP and MedConcept-23M. (a) Bar plot depicting the number of image-text pairs for each image type in MedConcept-23M. (b) Evaluation tasks categorized by modality and subtask, along with the number of images in each subtask. (c) ConceptCLIP demonstrates state-of-the-art performance across various medical imaging tasks. (d) Pre-training strategy for ConceptCLIP, involving joint pre-training with global image-caption alignment and local patch-concept alignment. (e) Examples of tasks that ConceptCLIP is capable of handling.
Implementing a vision-language pre-training model capable of handling various modalities offers several key advantages for enhancing trustworthiness in medical AI systems. First, such a model simplifies the construction of these systems by integrating two traditionally separate modules—analysis and explainability—into a single framework. Second, it addresses the challenge of explanation data [9] scarcity, which is a significant barrier in developing explainable AI systems. Traditionally, creating explanations for specific types of medical images requires substantial resources. This includes getting help from medical experts in different fields to annotate explanations for these images [10, 9, 11, 12, 13]. In scenarios where there are few explanation annotations available, conducting explainable AI becomes nearly impossible. However, a robust vision-language model capable of operating across diverse modalities can significantly alleviate this issue by enabling high-quality zero-shot concept annotation. This capability makes explainable AI feasible even in medical domains with limited explanation annotations, thereby advancing the development of trustworthy medical AI systems.
To support this endeavor, we introduce MedConcept-23M, a large-scale dataset comprising 23 million image-text-concept triples derived from 6.2 million scientific articles and enriched with medical concepts from the Unified Medical Language System (UMLS) [14]. Unlike previous datasets [7, 8], MedConcept-23M is enhanced with the knowledge of medical concept, providing fine-grained textual information paired with various image modalities. As shown in Figure 1 (a), MedConcept-23M includes a diverse range of image types within the field of biomedicine, such as radiology, microscopy, photography, 3D reconstructions, and signal images, covering a broad spectrum of visual data.
Building on this dataset, we develop ConceptCLIP, the first medical vision-language pre-training model designed for both accurate analysis and interpretable understanding of medical images across multiple modalities. As shown in Figure 1 (d), given the nuanced information inherent in medical images, which is critical for precise diagnosis and treatment, ConceptCLIP employs two main pre-training components: image-text alignment (IT-Align) and patch-concept alignment (PC-Align). IT-Align facilitates the global alignment of medical image and text representations, while PC-Align refines this alignment by incorporating concepts from UMLS [14], enabling local alignment between image patches and their corresponding concepts. This approach enhances the precision of medical image analysis by combining global and fine-grained local alignment across image and text modalities.
To evaluate the performance of ConceptCLIP, we develop the most comprehensive benchmark on 5 types of medical image analysis tasks, spanning 51 medical image analysis subtasks across 10 image modalities (Figure 1 (b)). Our evaluation includes 36 subtasks for medical image diagnosis, incorporating zero-shot, linear probes, and full fine-tuning settings. Additionally, it includes 2 subtasks for cross-modal retrieval, 2 subtasks for report generation, 2 subtasks for visual question answering, and 9 subtasks for whole-slide image analysis. As depicted in Figure 1, ConceptCLIP consistently achieves state-of-the-art results. The outcomes in cross-modal retrieval, zero-shot classification, and visual question answering demonstrate that ConceptCLIP is the leading vision-language pre-training model in the medical domain. The results from linear probes, full fine-tuning classification, medical report generation, and whole-slide image analysis underscore the robustness of its vision encoder, highlighting its capability and potential applications across various medical imaging modalities. Further explainability analysis across 6 modalities (Figure 1(b) and Figure 1(c)) confirms ConceptCLIP superior performance, highlighting its strong potential to advance explainable AI in medical imaging and enhance the trustworthiness of AI in medicine.
In summary, our contributions are threefold:
  • We curate MedConcept-23M, a large-scale medical dataset comprising 23 million medical image-text pairs derived from 6.2 million PubMed articles. Utilizing the UMLS system, each image-text pair is supplemented with UMLS concept information, resulting in a fine-grained knowledge-enhanced dataset with 23 million image-text-concept triplets.
  • We develop ConceptCLIP, the state-of-the-art medical vision-language pre-training model capable of accurate analysis and interpretable understanding of medical images across diverse modalities, leveraging both image-text and patch-concept alignments.
  • We conduct the most extensive evaluation on 5 types of tasks spanning 51 medical image analysis subtasks, across 10 image modalities. The results demonstrate that ConceptCLIP achieves state-of-the-art performance, establishing it as the leading vision-language pre-training model in the medical domain and featuring the most advanced vision encoder, which has potential applications across various image modalities. Further explainability analysis across 6 modalities highlights the superior performance of the proposed model, showcasing the strong capability to advance trustworthy AI in the field of medicine.
Results
Large-scale pre-training dataset with image-text-concept triplets and comprehensive evaluation benchmark
This section outlines the creation and application of MedConcept-23M, a pioneering dataset that serves as the foundational pre-training resource for ConceptCLIP, a state-of-the-art medical vision-language model. Additionally, we detail the comprehensive evaluation benchmark employed to assess ConceptCLIP across a diverse array of downstream tasks.
Overview of the MedConcept-23M Dataset
The MedConcept-23M dataset is designed to facilitate the pre-training of ConceptCLIP, offering a large-scale resource of 23 million medical image-text pairs, each augmented with UMLS concept information. Specifically, derived from the extensive PubMed Central Open Access Subset (PMC-OA) [15], MedConcept-23M provides a rich treasure of medical knowledge across various domains. By leveraging advanced concept extraction techniques [^1] https://github.com/allenai/scispacy, each image-text pair is enriched with relevant UMLS concepts, resulting in a dataset that is both voluminous and semantically rich. This comprehensive dataset forms the backbone for pre-training ConceptCLIP, enabling it to capture intricate medical relationships and semantics.
Comprehensive evaluation benchmark
Following its pre-training on MedConcept-23M, ConceptCLIP is subjected to a rigorous evaluation across 51 distinct medical image analysis subtasks, as detailed in Table LABEL:tab:datasets. This benchmark is the most exhaustive to date, encompassing 10 different image modalities and a wide range of tasks, including medical image diagnosis, cross-modal retrieval, report generation, visual question answering, and whole-slide image analysis. Importantly, there is no overlap between MedConcept-23M and the datasets used for evaluation, ensuring an unbiased assessment of ConceptCLIP’s capabilities.
The evaluation benchmark measures the performance of ConceptCLIP across various medical image analysis tasks. Through explainable analyses conducted across six modalities, we highlight ConceptCLIP’s ability to provide interpretable insights, reinforcing its utility in developing trustworthy AI systems in medicine.
Superior diagnostic capabilities of ConceptCLIP in medical image analysis
In this section, we present a comprehensive analysis of the performance of ConceptCLIP in medical image diagnosis across various imaging modalities and experimental settings. The results demonstrate the superior diagnostic capabilities of ConceptCLIP, highlighting its potential as a robust tool in medical image classification.
Zero-shot classification performance
In this experiment, we conduct a systematic evaluation of ConceptCLIP’s ability in the zero-shot setting, where our model is tested on classifying novel classes without requiring additional fine-tuning. Table 3 illustrates the AUC scores for zero-shot classification across different medical imaging modalities, including X-Ray, Fundus, Pathology, Endoscopy, Mammography, Dermoscopy, Ultrasound, OCT, MRI, and CT. ConceptCLIP consistently achieves the highest AUC scores across most datasets, outperforming competing models such as CLIP, SigLIP-400M, PMC-CLIP, and BiomedCLIP. Notably, in the X-Ray modality, ConceptCLIP achieves an average AUC score of 70.82, significantly surpassing the second-best model, BiomedCLIP, which scores 61.05. This trend of superior performance is consistent across other modalities, underscoring the model’s effectiveness in zero-shot settings.
Linear probes with varying training data
To further explore the capabilities of the image encoder of ConceptCLIP, we take 36 subtasks in medical image analysis to evaluate the ability of image representation. The experimental results are demonstrated in the linear probes setting with 1%, 10%, and 100% training data, as detailed in Tables 4, 5, and 6, respectively. ConceptCLIP consistently outperforms other models across all modalities and training data percentages. For instance, with 1% training data, ConceptCLIP achieves an average AUC of 67.95 in the X-Ray category, compared to the second-best average of 64.75 by SigLIP-400M. This performance advantage is maintained and even amplified as the percentage of training data increases, demonstrating the model’s ability to leverage additional data effectively to improve diagnostic accuracy.
Full fine-tuning performance
In the full fine-tuning setting, as depicted in Table 7, ConceptCLIP continues to exhibit outstanding performance across all imaging modalities. The model achieves near-perfect scores in several datasets, such as the Breast Tumor MRI and Brain Tumor CT, where it reaches an AUC of 100.00. This indicates ConceptCLIP’s capability to fully exploit the available data, achieving exceptional diagnostic precision and reliability.
ConceptCLIP shows strong performances in advanced medical tasks
This section delves into the capabilities of ConceptCLIP beyond traditional medical image diagnosis, exploring its performance in cross-modal retrieval, visual question answering, medical report generation, and whole-slide image analysis. The results underscore ConceptCLIP’s versatility and effectiveness across a range of complex tasks, highlighting its potential to enhance various facets of medical data analysis.
Cross-modal retrieval
ConceptCLIP is capable of identifying and retrieving the most relevant textual information given an image input (image-to-text retrieval) and conversely, retrieving images based on a text input (text-to-image retrieval). To assess this capability, we use the PMC-9K and QUILT-1M datasets, as detailed in Tables 8 and 9. ConceptCLIP achieves the highest Recall scores in both Image-to-Text and Text-to-Image tasks. Specifically, in the PMC-9K dataset, ConceptCLIP achieves a Recall@1 of 82.85 for Image-to-Text retrieval, significantly surpassing the second-best model, BiomedCLIP, which scores 73.41. This superior performance demonstrates ConceptCLIP’s proficiency in understanding and linking visual and textual information, a crucial capability for developing integrated medical information systems.
Visual question answering
The visual question answering performance of ConceptCLIP is analyzed using the SLAKE and VQA-RAD datasets. ConceptCLIP achieves the highest overall accuracy in both datasets, with an overall accuracy of 83.91 on SLAKE and 70.78 on VQA-RAD. These results indicate that ConceptCLIP can effectively comprehend and respond to complex medical queries, demonstrating its potential as a decision-support tool in clinical environments where quick and accurate medical decision-making is essential.
Medical report generation
Table 11 presents the performance of ConceptCLIP on the medical report generation task using the MIMIC-CXR and IU X-Ray datasets. ConceptCLIP consistently outperforms or reaches competitive performance with other models, achieving the highest scores in most metrics, including Micro Precision, Micro Recall, Micro F1, BLEU scores, ROUGE-L, METEOR, and CIDEr. Notably, in MIMIC-CXR, ConceptCLIP achieves a Macro F1 score of 4.43 and a BLEU-4 score of 8.36, indicating its superior ability to generate accurate and coherent medical reports. This performance suggests that ConceptCLIP can effectively understand and summarize complex medical information, providing valuable support in clinical documentation.
Whole-slide image analysis
In the realm of whole-slide image analysis, as shown in Table 12, ConceptCLIP demonstrates exceptional performance across various tasks, including cancer diagnosis, molecular subtyping, and survival prediction. For instance, in the Cancer Diagnosis task on the BRACS-3 dataset, ConceptCLIP achieves an AUC score of 91.65, outperforming all other models, especially pathology vision-language models like QuiltNet [16] and PLIP [17]. Similarly, in the Molecular Subtyping task for BRCA, ConceptCLIP records the highest AUC of 74.36. These results highlight ConceptCLIP’s capability to process and analyze high-dimensional pathology images, making it a promising tool for enhancing diagnostic accuracy in histopathology.
ConceptCLIP excels in advancing explainable AI
Medical concept annotation
Figure 2: Visualization of zero-shot medical concept annotation. For each example, the focused regions are highlighted based on the gradient of image-concept similarities, where the predicted probabilities of concepts are also presented.
Figure 3: Concept-based data analysis on datasets of various modalities. We calculate the concept presence difference to obtain which concepts are more likely to appear in which image set using the zero-shot concept annotation capability of our model.
Figure 4: Examples of concept-class association. We present the top three relevant concepts for selected diseases, ranked by their weights in the concept linear layer, paired with an example image of the specific disease from each considered dataset.
Figure 5: Disease-level concept inspection. For a specific disease, disease-level concept inspection conducts concept-disease association analysis from multiple datasets of different institutions to get more accurate and consistent concept-based explanations.
Trustworthiness is essential for AI models to be deployed and used in healthcare, where doctors and patients tend to trust the models with potent reliability and interpretability. This stringent trustworthiness requirement of the medical field has catalyzed research into Explainable Artificial Intelligence (XAI) for medical image analysis [18, 19, 20], with concept-based methods as one of the representative explainable approaches [20, 21, 22]. These methods integrate human-understandable concepts into the model development and evaluation, offering concept-based explanations and improving model interpretability. Medical concepts can be utilized to make models and their predictions more trustworthy because they provide a meaningful reference for human users by predicting the clinical-relevant attributes and symptoms instead of only giving the diagnosis results. However, most existing generalist medical vision language models ignore the importance of this valuable medical information. Furthermore, fine-grained concept annotations are highly valuable and require human experts’ efforts. Some methods propose to utilize general pre-trained foundation models (e.g., CLIP) to offer concept information [23, 24] but still suffer from low performance in specific domains such as healthcare. Hence building a medical model that is able to annotate concepts in a zero-shot setting holds the potential to facilitate further research for XAI.
In this work, we utilize the fine-grained concept knowledge from the textual data using the UMLS terminology [14]. We evaluate the performance of zero-shot medical concept annotation in various modalities, including dermoscopy, hematology, breast ultrasound, and chest CTs. The results are shown in Table 13. Specifically, our model achieves state-of-the-art performance, outperforming other compared vision-language foundation models. The average performance of ConceptCLIP achieves 10.9% and 10.4% relative performance boost over the results of PMC-CLIP and BiomedCLIP, respectively. Additionally, our proposed model outperforms the domain-specific foundation model MONET on the dermoscopic image datasets. In Fig. 2, we present the image-concept visualization to help better understand the model decision by highlighting the focus region of our model for specific concepts. From the experimental results, it can be observed that our model has the capability to identify clinical concepts across various modalities, thereby holding the potential to facilitate the development of trustworthy models and fine-grained datasets.
Concept-based data analysis
To facilitate trustworthy AI, the quality of the data used is crucial, and the correctness of the correlations within the data impacts model performance. Therefore, it is essential to have a foundation model capable of analyzing correlations between human-understandable concepts and target classes across various image sets. Previous research MONET [3] conducts data auditing to validate the correlations within data. However, it can only be applied to analyze dermoscopic images, which is insufficient for trustworthy AI in general medicine. In contrast, ConceptCLIP can analyze medical images of various modalities by measuring the concept presence difference of different image sets (Method) for medical datasets without fine-grained concept labels. Specifically, to demonstrate the concept-based data analysis ability of our model, we conduct experiments on datasets including modalities of chest X-rays, mammography, pathology, and ultrasound. The results are shown in Fig. 3. We observe that ConceptCLIP successfully identifies differentially present concepts in different datasets. For example, in the analysis for NLM-TB dataset [25], our model concludes that the concepts cavitation, consolidation, opacity, lymphadenopathy, etc. are more likely to present in the image set of tuberculosis while clear lung fields, normal diaphragmatic contours, etc., present more in normal chest X-rays, which are consistent to the human expectation.
Inherently interpretable model employment and evaluation
Within explainable AI, various approaches have been proposed to explain neural networks. Some methods utilize saliency maps [26, 27, 28] to highlight the contribution of each pixel or region in the model’s predictions, while some focus on feature interactions [29], and influence functions [30] to explain the model. However, these post-hoc XAI techniques may not be faithful enough to truly reflect the model decision-making process [31]. Therefore, building inherently interpretable models such as Concept Bottleneck Models (CBMs) [22] gains increasing attention since these models allow doctors and patients to easily understand the model decisions, hence holding the potential to advance trustworthy AI in medicine. Specifically, CBM, as an inherently interpretable model, first predicts the concepts present in the given images and then makes the final prediction based on the concepts through a linear layer. For the concept bottleneck layer, each neuron represents one pre-defined human-understandable concept, thus the weights of the layer can be regarded as the contribution of each concept to the final class (i.e., concept-class association). In this paper, we report the results of inherently interpretable models built upon ConceptCLIP and other vision-language models on disease diagnosis, as shown in Table 14. Our proposed model significantly outperforms all other compared vision-language models, achieving a 6.13% improvement in AUC score over the second-best method. Additionally, the results of concept-class association are shown in Fig. 4, where we present the top 3 relevant concepts for each selected disease across four image modalities. The results of inherently interpretable model employment demonstrate that ConceptCLIP achieves promising diagnosis performance while offering concept-based explanations.
Disease-level concept inspection
In healthcare, concept-based explanations can help medical experts and patients understand the decision process of AI models. In addition to the analysis for a single dataset, a global concept inspection for a disease beyond the granularity of the dataset level is essential for human users to better understand how the model deciphers the disease in terms of human-interpretable concepts. The inspection results may be biased when only one dataset is considered, due to potential spurious correlations within data and the limited number of data samples. The disease-level concept inspection result of our model is shown in Fig. 5, with pneumonia as the selected disease. It can be observed that when only one dataset (e.g., RSNA) is inspected, the output top relevant concepts for pneumonia include consolidation, infiltrates, atelectasis, and effusion. The result changes to the concepts of consolidation, infiltrates, atelectasis, and opacities when more images of pneumonia extracted from different datasets and institutions are considered (e.g., RSNA, VinDr-CXR, VinDr-PCXR). The results of disease-level concept inspection across multiple datasets are more consistent with the ground-truth concepts given by a certified radiologist collaborator. The observations demonstrate the potential of our model to inspect specific diseases with more accurate and consistent global concept-based explanations beyond a single dataset.
Discussion
The introduction of ConceptCLIP, a medical vision-language pre-training model, marks a significant advancement in the field of medical AI, particularly in the realms of image analysis and explainability. By integrating image-text and patch-concept alignments, ConceptCLIP has demonstrated superior performance across a diverse range of medical imaging tasks. However, alongside these accomplishments, the experimental results have also highlighted certain limitations that provide valuable insights for future research and development.
One notable strength of ConceptCLIP is its ability to perform zero-shot medical image diagnosis and cross-modal retrieval with high accuracy. This capability underscores the model’s robustness and adaptability across various medical image modalities without requiring extensive task-specific fine-tuning. However, despite these strengths, the model’s performance in certain complex tasks, such as whole-slide image analysis and visual question answering, suggests room for improvement. These tasks often involve intricate visual and contextual information that might not be fully captured by the current alignment mechanisms. Enhancing the model’s ability to understand and integrate complex visual patterns and contextual cues could further improve its effectiveness in these areas.
Moreover, while ConceptCLIP excels in leveraging UMLS concepts for enhanced interpretability, the reliance on predefined medical concepts may limit the model’s flexibility in adapting to novel or rare medical conditions that are not well-represented in existing medical ontologies. Future iterations of the model could benefit from incorporating more dynamic and adaptive learning strategies that allow for the integration of new concepts as they emerge in medical literature and practice. This could be achieved by employing continual learning frameworks or by integrating external knowledge bases that are frequently updated.
Another limitation arises from the data used for pre-training ConceptCLIP. Although the MedConcept-23M dataset provides a comprehensive resource of medical image-text pairs, it is derived from a specific subset of scientific articles, which may not fully capture the diversity and variability of real-world clinical data. This constraint could affect the generalizability of the model to different clinical settings or populations. Expanding the dataset to include more diverse sources of medical images and texts, such as electronic health records or data from different geographical regions, could enhance the model’s applicability and robustness.
In terms of explainability, while ConceptCLIP offers promising capabilities for concept annotation and inherently interpretable model construction, the evaluation of these features is largely dependent on the accuracy and relevance of the UMLS concepts. The current evaluation metrics may not fully capture the nuanced understanding required for clinical decision-making. Developing more sophisticated evaluation frameworks that consider clinical utility and relevance could provide a more comprehensive assessment of the model’s explainability.
In conclusion, ConceptCLIP represents a significant step forward in the integration of analysis and explainability in medical AI. However, addressing the identified limitations through methodological enhancements and broader data integration will be crucial for realizing the full potential of ConceptCLIP in advancing trustworthy AI in medicine. Future research should focus on refining the model’s adaptability, expanding its data sources, and developing more comprehensive evaluation metrics to ensure its effectiveness and reliability in diverse clinical contexts.
Method
In this section, we first provide a detailed description of the construction of medical image-text-concept triples. Subsequently, we explore how these triples can be utilized for model pre-training.
MedConcept-23M: a dataset of 23 million medical image-text-concept triples
In this section, we detail the construction of the MedConcept-23M dataset, which encompasses two primary stages: dataset collection and concept extraction. Following this, we examine the statistical properties of the proposed MedConcept-23M.
Dataset collection
To collect large-scale medical image-text pairs for pre-training a robust medical vision-language model, we focus on the PubMed Central Open Access Subset (PMC-OA)[15], which contains 6,246,351 articles under licenses permitting reuse (as of August 19, 2024). The articles are downloaded from https://ftp.ncbi.nlm.nih.gov/pub/pmc/oa_package/. Each article includes original images and parsed XML files. Utilizing this information, we employ PubMed Parser[32] to extract image-text pairs from each article, yielding a total of 23,289,898 pairs. We allocate 23,000,000 pairs for model pre-training and 289,898 pairs for validation and testing. For the latter set, we apply an image classifier [33] and InternVL [34] to rigorously filter out images containing non-medical information. This process results in the formation of PMC-9K, a cross-modal retrieval dataset in the medical domain for evaluation purposes.
Concept extraction
To leverage the knowledge embedded in UMLS for enhancing model pre-training, we extract UMLS concepts from each image caption, thereby coupling the image-text pair with its corresponding UMLS concepts. Specifically, for each caption, we employ SciSpacy [35] linked to the UMLS embedding database to extract entities directly associated with specific UMLS concepts. We filter out recognized UMLS concepts with a matching similarity below 0.8, ultimately extracting relevant UMLS concepts for each caption. This process results in the formation of the MedConcept-23M dataset, comprising 23 million medical image-text-concept triples.
Dataset statistics
As illustrated in Figure 1 (a), the curated dataset encompasses medical images from various domains, including Radiology, Photography, Microscopy, etc. These images form the foundation for pre-training a robust medical vision-language model capable of handling diverse image modalities.
ConceptCLIP: large-scale medical vision-language model with concept-enhanced pre-training
In this section, we first introduce the model architecture employed for pre-training. Building on this model, we then describe the training objectives in detail, including global image-caption alignment and local patch-concept alignment.
Consider a dataset comprising $N$ image-caption-concept triples, denoted as $\mathcal{D}=\{(I_{1},T_{1},G_{1}),(I_{2},T_{2},G_{2}),\ldots,(I_{N},T_{N},G_{N})\}$, where $I_{m}$ represents the $m$-th image, $T_{m}$ is the $m$-th caption, and $G_{m}$ consists of the concepts extracted from the $m$-th caption. Our objective is to develop a medical vision-language model utilizing an image encoder $f(\cdot)$ and a text encoder $g(\cdot)$. Specifically, for a given image-text pair $(I_{m},T_{m})$, the Vision Transformer-based image encoder $f(\cdot)$ and the BERT-based text encoder $g(\cdot)$ are employed to encode the pair as follows:
$\mathbf{i}_{m}=f(I_{m})$ (1)
$\mathbf{t}_{m}=g(T_{m})$ (2)
Here, $\mathbf{i}_{m}=\{\mathbf{i}_{\text{cls}},\mathbf{i}_{1},\mathbf{i}_{2},...,\mathbf{i}_{r}\}$ represents the encoded representation of the $m$-th image, where $\mathbf{i}_{m}\in\mathbb{R}^{(r+1)\times h}$. $\mathbf{t}_{m}=\{\mathbf{t}_{\text{cls}},\mathbf{t}_{1},\mathbf{t}_{2},...,\mathbf{t}_{s}\}$, where $\mathbf{t}_{m}\in\mathbb{R}^{(s+1)\times h}$ represents the encoded representation of the $m$-th text, $r$ is the number of image patches, $s$ is the number of text tokens, $h$ is the hidden size, $\mathbf{i}_{\text{cls}},\mathbf{t}_{\text{cls}}$ are the global representation of an image and a text sequence, respectively.
Training strategy
In this section, we introduce a dual alignment training strategy, i.e., image-text alignment learning and patch-concept alignment learning.
Image-text alignment learning (IT-Align).
We take sigmoid loss following the idea of SigLIP [5] to conduct the image-text alignment:
$\mathcal{L}_{\text{IT-Align}}=-\frac{1}{|\mathcal{B}|}\sum^{|\mathcal{B}|}_{i=1}\sum^{|\mathcal{B}|}_{j=1}\text{log}\frac{1}{1+e^{{z^{ij}(-t\mathbf{x}_{i}\cdot\mathbf{y}_{j}+b)}}}$ (3)
where $|\mathcal{B}|$ is the size of the mini-batch of image-text pairs, $t$ and $b$ are the logit scale and logit bias for the image-text alignment. $z_{ij}=1$ if the image and text are pairs, and $z_{ij}=-1$ otherwise. $\mathbf{x}_{i}=\frac{\mathbf{i}_{i,\text{cls}}}{||\mathbf{i}_{i,\text{cls}}||}$ and $\mathbf{y}_{j}=\frac{\mathbf{t}_{j,\text{cls}}}{||\mathbf{t}_{j,\text{cls}}||}$ are the $i$-th and the $j$-th normalized image and text representations respectively. These representation vectors are obtained following the attentive pooling in [5] and mean pooling in [36]. The two vectors are then used to calculate the similarity between them.
Patch-concept alignment learning (PC-Align).
To conduct fine-grained alignment between the image and text, the matching between image patches and medical concepts are explored. As shown in Figure 1 (d), image patches are encoded via the image encoder into $\mathbf{i}_{m}$:
$\mathbf{i}_{m}=f(I_{m})$ (4)
For each text $T_{m}=\{t_{1},t_{2},...,t_{N}\}$, where there are $N$ tokens in $T_{m}$, the text encoder encodes the text into $\mathbf{t}_{m}=\{\mathbf{t}_{m,1},\mathbf{t}_{m,2},...,\mathbf{t}_{m,s}\}$:
$\mathbf{t}_{m}=g(T_{m})$ (5)
Considering there is the $b$-th concept in the $m$-th text is $G_{m,b}$, which has the in-text indices: $\{v_{1},v_{2},...,v_{U}\}$, where the length of the concept is $U$, then through the encoded text $\mathbf{t}_{m}$, the concept representation can be derived:
$\mathbf{g}_{m,b}=\text{mean\_pooling}(\mathbf{t}_{m,v_{1}},\mathbf{t}_{m,v_{2}},...,\mathbf{t}_{m,v_{U}})$ (6)
For the $m$-th image-text pair, a patch-concept similarity matrix $\mathbf{A}\in\mathbb{R}^{r\times w}$, where $w$ is the number of concepts in the $m$-th image-text pair. Then, for the $i$-th patch and the $j$-th concept, the alignment score is:
$(\mathbf{a})_{ij}=\frac{(\mathbf{g}_{m,j})^{T}\mathbf{t}_{m,i}}{||\mathbf{g}_{m,j}||\cdot||\mathbf{t}_{m,i}||}$ (7)
$(\mathbf{A})_{ij}=\text{log}\frac{1}{1+e^{z^{m,n}(-t_{g}\mathbf{a}_{i,j}+b_{g})}}$ (8)
where $z^{m,n}=1$ if the $m$-th image and the $n$-th text are paired. Otherwise, $z^{m,n}=-1$. $t_{g}$ and $b_{g}$ are the logit scale and logit bias for the patch-concept alignment.
For each image $I_{m}$ and text $T_{n}$, the similarity score is:
$S(I_{m},T_{n})=\frac{1}{w}\sum_{j=1}^{w}\max_{i}(\mathbf{A}_{ij})$ (9)
Thus, the PC-Align loss in a mini-batch of $|\mathcal{B}|$ image-text-concept triples is:
$\mathcal{L}_{\text{PC-Align}}=-\frac{1}{|\mathcal{B}|}\sum_{m=1}^{|\mathcal{B}|}\sum_{n=1}^{|\mathcal{B}|}S(I_{m},T_{n})$ (10)
Overall, the total training loss is:
$\mathcal{L}=\mathcal{L}_{\text{IT-Align}}+\alpha\mathcal{L}_{PC-Align}$ (11)
where $\alpha$ is a hyper-parameter to decide the weight of the PC-Align loss during the pre-training.
Inference with local and global information integration
In this section, we elucidate the inference strategy employed in our model, which integrates both local and global information to enhance the zero-shot classification performance [37]. This approach is particularly advantageous in the medical domain, where nuanced and context-specific details often play a crucial role in accurate diagnosis.
The integration of local and global information is inspired by the need to capture both the holistic context of an image and the specific details that may be critical for identifying particular medical conditions. While global image-caption alignment provides a broad understanding of the visual and textual data, local patch-concept alignment allows for a more detailed examination of specific regions within an image that correspond to medical concepts.
Our inference strategy employs a dual-classifier system consisting of a global classifier and a local classifier. The global classifier leverages the overall image features, while the local classifier focuses on the alignment between image patches and text-derived concepts. By combining these classifiers, our model can make more informed decisions that account for both general and specific information.
During inference, the model processes input images through an image encoder to extract image features, denoted as $\mathbf{i}_{m}$. Simultaneously, the text encoder processes the associated concepts to derive text features, $\mathbf{t}_{m}$. The global classifier computes logits based on the similarity between the normalized global image features and the global class representations obtained during pre-training. For the local classifier, the model evaluates patch-level features against concept representations. This involves computing a patch-concept similarity matrix, which quantifies the alignment between image patches and text-derived concepts. The top-k alignment scores are aggregated to form the local logits, reflecting the degree of correspondence between image patches and medical concepts.
The final classification decision is derived by combining the outputs of the global and local classifiers. A weighted sum of the global and local logits is computed, controlled by a hyper-parameter $\beta$, which balances the contribution of each classifier:
$=\beta\cdot\text{local-logits}+(1-\beta)\cdot\text{global-logits}$-logits (12)
This integration allows the model to harness the strengths of both global and fine-grained local information, thereby improving the robustness and accuracy of zero-shot classification in medical imaging tasks.
Experimental settings
In this section, we first describe the dataset used to pre-train the medical vision-language model. Based on the pre-trained model, we conduct extensive experiments on downstream tasks. Besides, we describe the baseline medical vision-language models in detail. Finally, we cover the implementation details in pre-training the proposed ConceptCLIP.
Pre-training dataset
In our pre-training process, the proposed MedConcept-23M is used. In detail, it contains 23 million image-text-concept triples, where the image-text pairs are collected from PMC-OA and concepts are aligned to the UMLS concepts and extracted from the text.
Downstream tasks
Medical image diagnosis
We assess the zero-shot, linear probe (using 1%, 10%, and 100% of the training data), and full fine-tuning capabilities of ConceptCLIP and previous state-of-the-art medical CLIP models across 36 datasets spanning 10 image modalities. In the zero-shot setting, no ground truth labels are provided. We follow the zero-shot protocol of the CLIP model [4], where the text encoder and image encoder are used to encode the medical images and labels (using a prompt template to form a sentence), calculate the similarity, and assign labels to the images. For linear probes, we first extract features from each image using a specific image encoder from a pre-trained model and employ these image features and their labels to perform logistic regression. The trained logistic regression models are subsequently used to test model performance. This experiment aims to evaluate each model’s ability to extract image features. In the full fine-tuning setting, each model is concatenated with a classification head and fine-tuned on a specific dataset with all parameters adjustable. We use the AUC score for model evaluation to ensure that the performance is not influenced by thresholds.
Cross-modal image-text retrieval
Cross-modal image-text retrieval includes tasks such as medical image-to-text and text-to-image retrieval. To evaluate the performance of ConceptCLIP across various medical image modalities, we utilize two datasets: PMC-9K and Quilt-1M [16], to assess the models’ cross-modal retrieval capabilities in a mix of image modalities and pathology images. We use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to evaluate each model’s performance.
Medical report generation
Medical report generation aims to alleviate the workload of doctors by automatically generating a report given a medical image. In our study, following the shallow alignment in the R2GenGPT framework [38], we replace the vision encoder with that of each medical CLIP model and use the same training scheme to fine-tune the report generation model. We conduct these experiments on the MIMIC-CXR [39] and IU X-Ray [40] datasets. To comprehensively evaluate each model’s performance, we use natural language generation metrics: BLEU-1,2,3,4, METEOR, ROUGE-L, CIDEr, and clinical efficacy metrics: Micro Precision, Micro Recall, and Micro F1.
Medical visual question answering
Medical Visual Question Answering (VQA) serves as a bridge between AI systems and humans. Given a medical image, this task involves providing a question about the image to the system, which should then answer the question based on the image. In our experiments, we use the METER framework [41] to evaluate different models on the VQA task, employing the SLAKE [42] and VQA-RAD [43] datasets. Following METER’s evaluation metrics, we use Closed Accuracy, Open Accuracy, and Overall Accuracy to assess each model’s performance.
Whole-slide image analysis
To explore our model’s ability to understand whole-slide images, we conduct experiments on 6 related datasets for 9 subtasks in Cancer Diagnosis, Molecular Subtyping, and Survival Prediction, each of which can be regarded as a classification task, with a whole-slide image as input to the vision encoder and a predicted label as output. Given the high resolution of whole-slide images, we follow the conventional approach by segmenting the whole-slide image into smaller patches and using each pre-trained vision encoder to extract features from each image patch. These features are input into the ABMIL model [44] for multiple instance learning (MIL). The model outputs a predicted label. Similar to the Medical Image Diagnosis task, we use the AUC score to evaluate different models in Cancer Diagnosis and Molecular Subtyping, and concordance index (C-index) for survival analysis.
Medical concept annotation
We evaluate to what extent can the proposed ConceptCLIP annotate fine-grained concepts for medical images, hence facilitating the development of concept-based explainable artificial intelligence. The performance of medical concept annotation is evaluated using the concept labels of the datasets. For each concept, two positive and negative prompts are designed to calculate the similarity with input images to get the predicted output. Specifically, we adopt five datasets with clinical concept annotations to evaluate the zero-shot concept annotation performance, including dermoscopic images, blood cell images (hematology), breast ultrasound, and chest CT.
Derm7pt [10]: A dermoscopic image dataset containing 1,011 images with clinical concepts for melanoma skin lesions in dermatology. The test set is adopted for evaluation. Only the dermoscopic images are considered in this paper for concept annotations. The seven clinical concepts include “Pigment network”, “Dots and globules”, “Pigmentation”, “Streaks”, “Regression structures”, “Blue-whitish veil”, and “Vascular structures”.
SkinCon [9]: A skin disease dataset densely annotated by experts for fine-grained model debugging and analysis. We choose 22 concepts that have at least 50 images representing the concept from the Fitzpatrick 17k (F17k) [45] subset. 350 images are sampled for evaluation. The clinical concepts include “Papule”, “Plaque”, “Pustule”, “Bulla”, “Patch”, “Nodule”, “Ulcer”, “Crust”, “Erosion”, “Atrophy”, “Exudate”, “Telangiectasia”, “Scale”, “Scar”, “Friable”, “Dome-shaped”, “Brown (Hyperpigmentation)”, “White (Hypopigmentation)”, “Purple”, “Yellow”, “Black”, and “Erythema”.
WBCAtt [11]: This dataset is a densely annotated dataset for white blood cell (WBC) recognition, containing 11 morphological attributes for 10,298 cell images. The test set which contains 3,099 images is used for evaluation. There are 11 concepts, including “Cell size”,“Cell shape”,“Nucleus shape”,“Nuclear cytoplasmic ratio”,“Chromatin density”, “Cytoplasm vacuole”, “Cytoplasm texture”, “Cytoplasm colour”, “Granule type”, “Granule colour”, and “Granularity”.
BrEaST [12]: A breast ultrasound dataset containing 256 images with fine-grained concepts for breast lesion analysis. Seven BI-RADS descriptors are adopted as the evaluated concepts, including “Shape”, “Magrin”, “Echogenicity”, “Posterior features”, “Halo”, “Calcifications”, and “Skin thickening”.
LUNA16 [13]: A curated version of LIDC-IDRI [46] dataset, which contains lung images with clinical concepts. 1,185 slices of lesions from the chest CTs with six clinical concepts are used for evaluation. The concepts include “Calcification”, “Lobulation”, “Margin”, “Sphericity”, “Spiculation”, and “Texture”.
Concept-based data analysis
To conduct concept-based data analysis, concept presence difference is calculated for datasets of four modalities, including NLM-TB (chest X-ray), CBIS-DDSM (mammography), UBIBC (breast ultrasound), and MedFMC-Colon (pathology). Specifically, for each dataset, assume there are two image sets of different classes (e.g., tuberculosis and normal in NLM-TB dataset), denoted as $I_{+}=\{I_{1+},I_{2+},...,I_{M+}\}$ and $I_{-}=\{I_{1-},I_{2-},...,I_{N-}\}$, and the pre-defined concept list $C$ (e.g., cavitation, consolidation, etc.), where $M$ and $N$ are the numbers of images in the positive and negative image sets, respectively, and $C$ are generated by a large language model (i.e., GPT-4 [47]) since the dataset does not have concept labels. We adopt the proposed ConceptCLIP to annotate the presence of the concepts for each given image. The concept presence proportion of each image set is computed by $\frac{N_{c_{i}}}{M}$ and $\frac{N_{c_{i}}}{N}$, respectively, here $N_{c_{i}}$ is the number of images where a specific concept $c_{i}$ is present. Then the concept presence difference of concept $c_{i}$ is defined as the difference between the two concept presence proportions, i.e., $D_{c_{i}}=\frac{N_{c_{i}}}{M}-\frac{N_{c_{i}}}{N}$. A positive value of $D_{c_{i}}$ means concept $c_{i}$ is more likely to appear in the positive image set $I_{+}$, and vice versa.
Inherently interpretable model employment and evaluation
Building inherently interpretable models is essential for deploying AI in medicine since it is more trustworthy for doctors and patients during clinical diagnosis. In this paper, we employ the Concept Bottleneck Model (CBM) [22] to build the inherently interpretable model based on our method. Inspired by previous works [23, 24], we use ConceptCLIP to calculate the cosine similarity between input images and medical concepts, then the similarity is used as the input to the concept bottleneck layer, which maps the concept similarity to the final prediction. The concept bottleneck layer is trained using the class labels. For concept-class association, the weights of each neuron within the linear layer can be regarded as the contribution to the final prediction. To evaluate the effectiveness of our method, we compare ConceptCLIP with other vision-language models of both the natural and medical domains on four datasets of different modalities, including dermoscopy, hematology, ultrasound, and CT.
Disease-level concept inspection
As shown in Fig. 5, pneumonia is selected as the inspected disease. For concept inspection on a single dataset, RSNA [48] dataset is adopted, which only includes classes of No finding and Pneumonia. For disease-level concept inspection, we collect datasets that also include the two classes, i.e., VinDr-CXR [49] and VinDr-PCXR [50], where VinDr-PCXR focus on the chest X-rays of children. It is noteworthy that VinDr-CXR and VinDr-PCXR contain other classes such as Lung tumor and Tuberculosis, but we only consider the images of No Finding and Pneumonia to analyze the pneumonia disease. For each dataset, we use ConceptCLIP to obtain the concept-disease association. Then we average the weights of the concept bottleneck layer across all datasets to derive the final top relevant concepts for pneumonia.
General domain models and the state-of-the-art medical models
CLIP [4] is a vision-language model pre-trained on large-scale image-text pairs via contrastive learning in the general domain, which showcases strong ability in zero-shot classification. We take the version of CLIP-ViT-B-16 pre-trained on LAION-2B [51] under the framework of OpenCLIP [52].
SigLIP [5] is a variant of the CLIP model, which is pre-trained via contrastive learning using the sigmoid loss. In our experiments, we use the version of ViT-SO400M-14-SigLIP (SigLIP-400M).
PMC-CLIP [7] is a vision-language model tailored for medicine. The authors collected PubMed open access articles and conducted a series of data processing operations, producing a dataset with 1.6 million medical image-text pairs. The dataset is used to pre-train a CLIP model, thus resulting in the PMC-CLIP model. We use the origin checkpoint [^2] https://huggingface.co/datasets/axiong/pmc_oa_beta/blob/main/checkpoint.pt that the authors provided.
BiomedCLIP [8] is a medical vision-language model pre-trained on PMC-15M, a dataset with 15 million medical image-text pairs sourced from PMC-OA. We take the released checkpoint [^3] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224 for evaluation.
Implementation details
We develop ConceptCLIP by utilizing SigLIP-ViT-400M-16444https://huggingface.co/timm/ViT-SO400M-14-SigLIP-384 as the visual encoder and PubMedBERT [^5] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as the text encoder. Each input image is resized to $336\times 336$ pixels. During pre-training, we employ the AdamW optimizer [53] with a learning rate of $5\times 10^{-4}$. The model is initially pre-trained on H800 GPUs with a batch size of 12,288 for 32 epochs on the MedConcept-23M dataset, without the PC-Align loss. Subsequently, we conduct further pre-training with the PC-Align loss on this dataset with a batch size of 6,144 and learning rate of $3\times 10^{-4}$ for 20 epochs. The weight $\alpha$ during the pre-training is set to 0.5.
In medical image diagnosis, for zero-shot classification, we set $\beta$ to 0.5 so that the prediction can consider the information from both global information and local information. For different datasets, we choose different “Top-K” values and prompts as described in LABEL:tab:zero_shot_settings. For linear probes, for all experiments, we use the LogisticRegression function provided in scikit-learn [^6] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, with inverse of regularization strength is 0.316, maximum iteration is 1,000, random state is 1, and all other parameters are default. For full fine-tuning experiments, for all experiments, the batch size is 64, epochs is 20, learning rate is 1e-4. We use AdamW optimizer for training the model.
In the visual question answering task, we use AdamW as the optimized. For SigLIP-400M and ConceptCLIP, the batch size is 32, learning rate is 1e-6. For other models, the batch size is 4, learning rate is 5e-6. All experiments have the same epoch number 50.
In whole-slide image analysis, the batch size is 1, learning rate is 2e-4, epochs are 30, and the cosine learning rate scheduler is used.
Metrics
AUC
The Area Under the Receiver Operating Characteristic Curve (AUC) is a performance measurement for classification problems at various threshold settings. It is widely used in medical image diagnosis and explainable AI tasks to evaluate the ability of a model to distinguish between classes.
Accurary
Accuracy (ACC) is the proportion of true results among the total number of cases examined. In the context of visual question answering tasks, we use closed accuracy, open-accuracy, and overall accuracy to assess the model’s ability to select the correct answer from a candidate answer set. The formula is:
$\text{ACC}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$
C-Index
The concordance index (C-Index) measures the predictive accuracy of a survival model. It is particularly used in survival prediction tasks in whole-slide image analysis to evaluate how well the model predicts the order of events. It is calculated as the proportion of all usable patient pairs whose predictions and outcomes are concordant.
Recall
Recall, or sensitivity, measures the ability of a model to correctly identify all relevant instances. In cross-modal retrieval tasks, we use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to assess the model’s retrieval performance at different levels. The formula is:
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
BLEU
The Bilingual Evaluation Understudy (BLEU) score [54] is a metric for evaluating the quality of text generated by a machine, such as translations or reports. It is used for medical report generation to assess the accuracy and fluency of the generated text. The formula for BLEU is:
$\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$
where $p_{n}$ is the precision for n-grams, $w_{n}$ are weights, and BP is the brevity penalty.
METEOR
The Metric for Evaluation of Translation with Explicit ORdering (METEOR) [55] is another metric for evaluating text generation quality, focusing on precision, recall, and alignment of phrases. It is used for medical report generation to provide a more nuanced assessment of generated text quality.
ROUGE-L
The Recall-Oriented Understudy for Gisting Evaluation (ROUGE-L) [56] measures the longest common subsequence (LCS) between the generated text and reference texts. It is used for evaluating medical report generation to capture the overlap in content. The formula for ROUGE-L is:
[…]
CIDEr
Consensus-based Image Description Evaluation (CIDEr) [57] is a metric designed to assess the similarity of a generated text to multiple reference texts, emphasizing consensus. It is used for medical report generation to evaluate the relevance and quality of the generated content.
Clinical efficacy
Clinical Efficacy measures the practical effectiveness of a generated medical report in a clinical setting, focusing on its impact on clinical outcomes and decision-making. It includes precision, recall, and F1, by annotating the generated report and comparing it with the ground truth report via 14 disease classification labels in CheXbert [58].
Micro precision calculates the precision of the model by considering all true positives and false positives across all classes, providing a detailed view of the model’s performance:
Micro $\text{Micro precision}=\frac{\sum\text{TP}}{\sum\text{TP}+\sum\text{FP}}$
Micro recall aggregates the true positives and false negatives across all classes to compute recall, offering a comprehensive assessment of the model’s ability to identify relevant instances:
Micro $\text{Micro recall}=\frac{\sum\text{TP}}{\sum\text{TP}+\sum\text{FN}}$
Micro F1 is the harmonic mean of Micro Precision and Micro Recall, providing a balanced measure of the model’s accuracy and completeness across all classes:
[…]
Data availability
The dataset used for pre-training is described in Section Method. The datasets used for evaluation are listed in Extended Data Table LABEL:tab:datasets.
Code availability
The implementation of ConceptCLIP will be available at https://github.com/JerrryNie/ConceptCLIP. The weights of ConceptCLIP will be released at https://huggingface.co/JerrryNie/ConceptCLIP.
Author contribution
Y.N., S.H., Y.B., and H.C. conceived and designed the work. Y.N., S.H., Y.B., Y.W., and Z.C. collected the data for pre-training and downstream task evaluation. Y.N. and S.H. contributed to the pre-training technical implementation. Y.B. contributed to the technical implementation of explainable AI. Y.N. and S.H. evaluated medical image diagnosis tasks. Y.N. evaluated cross-modal retrieval tasks. Y.B. evaluated explainable AI tasks. Y.W. evaluated whole-slide image analysis tasks. Z.C. evaluated medical report generation tasks. S.Y. evaluated medical visual question answering tasks. Y.N., Y.B., and S.H. wrote the manuscript with inputs from all authors. All authors reviewed and approved the final paper. H.C. supervised the research.
Declarations
The authors have no conflicts of interest to declare.
Appendix A Extended Data
Table 1: Medical image analysis datasets.
Dataset Website Description
Binary Classification
SIIM-ACR [59] https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation A binary classification dataset containing chest radiographic images.
Covid-CXR2 [60] https://www.kaggle.com/datasets/andyczhao/covidx-cxr2 A dataset of over 16,000 chest X-ray images from more than 15,100 patients across 51 countries, including 2,300 positive COVID-19 images, designed to differentiate between no pneumonia, non-COVID-19 pneumonia, and COVID-19 pneumonia in the COVIDx V8A dataset, while the COVIDx V8B dataset focuses on detecting COVID-19 positive and negative cases.
NLM-TB [25] https://www.ncbi.nlm.nih.gov/pmc/articles/PMC4256233/ A radiology image dataset with 800 total images with the label of normal or tuberculosis.
LC25000 (Colon) [61] https://github.com/tampapath/lung_colon_image_set A dataset comprises 10,000 pathology images from colon tissue, featuring benign tissues and adenocarcinomas.
UBIBC [62] https://www.kaggle.com/datasets/vuppalaadithyasairam/ultrasound-breast-images-for-breast-cancer This dataset consists of ultrasound images related to breast cancer, both benign and malignant.
PCam200 [63] https://github.com/enigmanx20/patchtcga A dataset of pathological H&E images created from the Camelyon2016 challenge dataset [64].
RSNA [48] https://www.rsna.org/rsnai/ai-image-challenge/rsna-pneumonia-detection-challenge-2018 A dataset of about 30,000 frontal view chest X-ray images, each labeled for binary classification to indicate the presence of pneumonia.
Brain Tumor CT [65] https://www.kaggle.com/datasets/murtozalikhon/brain-tumor-multimodal-image-ct-and-mri A dataset of CT scans for brain tumor detection and analysis. It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
Brain Tumor MRI [65] https://www.kaggle.com/datasets/murtozalikhon/brain-tumor-multimodal-image-ct-and-mri A dataset of MRI scans, for brain tumor detection and analysis. It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
DDSM [66] https://www.kaggle.com/datasets/skooch/ddsm-mammography A dataset containing 55,890 pre-processed images from the DDSM [67] and CBIS-DDSM [68] datasets, formatted as 299x299 pixel tfrecords for TensorFlow, with an imbalance of 14% positive and 86% negative examples, incorrectly split into test and validation sets.
Breast Cancer [69] https://www.kaggle.com/datasets/hayder17/breast-cancer-detection/ A dataset of 3,383 annotated mammogram images focused on breast tumors, exported from Roboflow, designed for building and testing deep learning models for tumor detection.
Multi-Label Classification
RFMiD2 [70] https://www.mdpi.com/2306-5729/8/2/29 A multi-label dataset of fundus images annotated by three eye specialists.
MedFMC (Colon) [71] https://github.com/openmedlab/MedFM A dataset focused on facilitating the detection of early-stage cancer cells in tissue slides, enabling pathologists to classify and quantify cancerous regions for lesion and non-lesion classes.
VinDr-Mammo [72] https://vindr.ai/datasets/mammo A large-scale benchmark dataset for computer-aided detection and diagnosis in full-field digital mammography.
VinDr-PCXR [50] https://vindr.ai/datasets/pediatric-chest-x-ray A dataset focuses on pediatric chest X-rays for the interpretation of common thoracic diseases.
VinDr-CXR [49] https://vindr.ai/datasets/cxr A dataset of chest x-ray images containing 18,000 high-quality postero-anterior scans with detailed localization of 22 critical findings and classification of 6 thoracic diseases, annotated by experienced radiologists from major hospitals in Vietnam. To make the official training and testing label consistent, we filter out part of inconsistent training data, thus resulting in 32,976 training images and 3,000 testing images.
VinDr-SpineXR [73] https://vindr.ai/datasets/spinexr A dataset of annotated medical images for detecting and classifying spinal lesions from radiographs.
NIH ChestX-ray14 [74] https://nihcc.app.box.com/v/ChestXray-NIHCC A dataset comprising 112,120 frontal-view X-ray images from 30,805 unique patients, featuring fourteen disease labels identified through NLP techniques, collected between 1992 and 2015.
CheXpert [58] https://stanfordmlgroup.github.io/competitions/chexpert/ A dataset of chest radiographs from 65,240 patients.
Multi-Class Classification
DRD [75] https://www.kaggle.com/competitions/diabetic-retinopathy-detection A dataset consists of high-resolution retina images labeled by subject ID and eye side, with each image rated for diabetic retinopathy severity on a scale of 0 (no DR) to 4 (proliferative DR) by a clinician.
LC25000 (Lung) [61] https://github.com/tampapath/lung_colon_image_set A dataset consists of 15,000 pathology images from lung tissue, including benign tissues and various carcinomas.
MedFMC (Chest) [71] https://github.com/openmedlab/MedFM A dataset for screening thoracic diseases encompasses 19 common thoracic abnormalities.
MedFMC (Endo) [71] https://github.com/openmedlab/MedFM A dataset aimed at enhancing the automatic detection and classification of four different lesion types in colonoscopy images, facilitating early diagnosis of colorectal cancer.
HAM10000 [76] https://www.kaggle.com/datasets/kmader/skin-cancer-mnist-ham10000 This dataset is composed of dermatoscopic images of pigmented lesions.
BUSBRA [77] https://zenodo.org/records/8231412 A dataset of anonymized breast ultrasound images from 1,064 patients, including biopsy-proven tumors, BI-RADS annotations, and ground truth delineations of tumoral and normal regions.
WCE [78] https://www.kaggle.com/datasets/francismon/curated-colon-dataset-for-deep-learning A dataset of colon disease images containing curated samples for training and testing, derived from the Kvasir [79] and ETIS-Larib-Polyp DB [80] datasets.
Fundus JSIEC [81] https://www.kaggle.com/datasets/linchundan/fundusimage1000 A dataset of 1,000 fundus images belonging to 39 classes, sourced from the Joint Shantou International Eye Centre in Shantou city, Guangdong province, China.
HyperKvasir [82] https://datasets.simula.no/hyper-kvasir/ A dataset consisting of 10,662 labeled images in JPEG format, categorized into 23 classes of medical findings, with each class represented by a specific folder, highlighting the imbalance in the number of images per class.
Kvasir [79] https://datasets.simula.no/kvasir/ A dataset of annotated and verified gastrointestinal tract images, featuring various classes of anatomical landmarks and pathological findings, organized in separate folders with resolutions ranging from 720x576 to 1920x1072 pixels, suitable for tasks such as image retrieval and machine learning.
ODIR [83] https://www.kaggle.com/datasets/andrewmvd/ocular-disease-recognition-odir5k A dataset of 5,000 patients featuring age, color fundus photographs from both left and right eyes, along with doctors’ diagnostic keywords, with 6,392 samples used for training.
BUID [84] https://www.kaggle.com/datasets/aryashah2k/breast-ultrasound-images-dataset A dataset of 780 breast ultrasound images from 600 female patients aged 25 to 75, collected in 2018, categorized into three classes: normal, benign, and malignant, with ground truth images included.
PAD-UFES-20 [85] https://www.notion.so/c1e0ff8f36814b94b205ee919e2e3ff8?pvs=21 A dataset of skin lesion images labeled into six categories: Basal Cell Carcinoma, Squamous Cell Carcinoma, Actinic Keratosis, Seborrheic Keratosis, Melanoma, and Nevus.
OCTMNIST [86] https://medmnist.com/ A dataset for multi-class classification of retinal OCT images.
Breast Tumor MRI [87] https://www.kaggle.com/datasets/masoudnickparvar/brain-tumor-mri-dataset A combined dataset (including three datasets: figshare [88], SARTAJ dataset [89]. and Br35H dataset [90]) of 7,023 human brain MRI images classified into four classes: glioma, meningioma, no tumor, and pituitary. The no tumor class images are from the Br35H dataset.
Retinal OCT [91] https://www.kaggle.com/datasets/obulisainaren/retinal-oct-c8 A dataset of high-quality retinal OCT images categorized into 8 classes of retinal diseases, designed for research and model training in classification using machine learning and deep learning.
COVIDxCT [92] https://www.kaggle.com/datasets/hgunraj/covidxct/data A dataset consisting of two variants, ”A” with confirmed COVID-19 cases and ”B” which includes weakly verified cases, designed for training, validation, and testing of CT scans. We choose the “A” set in our experiment.
Retrieval
PMC-9K - This newly curated dataset consists of a cleaned, held-out collection of 9,222 image-text pairs in MedConcept-23M. It is designed to comprehensively evaluate the cross-modal retrieval capabilities of ConceptCLIP across various modalities.
Quilt-1M [16] https://quilt1m.github.io/ This dataset is a large-scale histopathology collection containing 768,826 image-text pairs. For evaluation purposes, we filter out the data sourced from PubMed Central to avoid data leakage and utilize a held-out subset comprising pathology image-text pairs to assess the retrieval performance.
Medical Report Generation
MIMIC-CXR [39] https://physionet.org/content/mimic-cxr/2.1.0/ A dataset contains 371,920 chest X-rays linked to 227,943 imaging studies from 65,079 patients.
IU X-Ray [40] https://www.kaggle.com/datasets/raddar/chest-xrays-indiana-university/data A dataset consists of 7,470 pairs of chest X-ray images and their corresponding diagnostic reports.
Visual Question Answering
VQA-RAD [43] https://osf.io/89kps/ A dataset called VQA-RAD is manually constructed, featuring 3,064 question-answer pairs where clinicians ask natural questions about radiology images and provide reference answers.
SLAKE [42] https://www.med-vqa.com/slake/ A bilingual radiology VQA dataset that includes 642 images and 14,000 questions. We only use the English part.
Cancer Diagnosis
BRACS-3 [93] https://www.bracs.icar.cnr.it/ A dataset contains 6 different subtypes of lesions including also images representing atypical lesions. Histological images representing normal tissue samples are also included. In this setting, a breast tumor according to a pathology image can be classified into “benign”, “atypical”, or “malignant”.
BRACS-7 [93] https://www.bracs.icar.cnr.it/ A dataset contains 6 different subtypes of lesions including also images representing atypical lesions. Histological images representing normal tissue samples are also included. In this setting, a breast tumor according to a pathology image can be classified into “normal”, “pathological benign”, “usual ductal hyperplasia”, “flat epithelial atypia”, “atypical ductal hyperplasia”, “ductal carcinoma in situ”, or “invasive carcinoma”.
BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset features 985 histopathology whole slide images of Breast Invasive Carcinoma, including 787 Invasive Ductal Carcinoma and 198 Invasive Lobular Carcinoma cases.
NSCLC [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset contains 1,053 histopathology slides of Non-Small Cell Lung Cancer, including 541 lung adenocarcinoma and 512 lung squamous cell carcinoma cases.
Camelyon [64, 95] https://camelyon16.grand-challenge.org/https://camelyon17.grand-challenge.org/ A dataset based on CAMELYON16 [64] and CAMELYON17 [95], which evaluates new and existing algorithms for automated detection and classification of breast cancer metastases in whole-slide images of histological lymph node sections.
Molecular Subtyping
BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset includes 985 histopathology whole slide images of breast invasive carcinoma, specifically 787 Invasive Ductal Carcinoma and 198 Invasive Lobular Carcinoma cases.
Survival Prediction
BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis. We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
LUAD [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis. We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
LUSC [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis. We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
Table 2: Zero-shot experimental settings for medical image analysis datasets.
Dataset Classes Prompts Top-K
SIIM-ACR [59] No Finding, Pneumothorax a chest radiology presents {} 256
RFMiD2 [70] Neovascularization, Macular Edema, Myopia, Retinal Traction, Coloboma, Choroidal folds, Tortuous Vessels, Retinitis Pigmentosa, Retinal pigment epithelium changes, Optic Disc Pallor, Media Haze, Retinitis, Preretinal Hemorrhage, Asteroid Hyalosis, Drusens, Hemorrhagic Pigment Epithelial Detachment, Branch Retinal Vein Occlusion, Optic Disc Edema, Exudation, Haemorrhagic Retinopathy, Tilted Disc, Tessellation, Retinal Tears, Retinal Detachment, Optic Disc Cupping, Macular Hole, Silicone Oil-Filled Eye, Cotton Wool Spots, Vasculitis, Microaneurysm, Macular Scar, Age-Related Macular Degeneration, Optic neuritis, Anterior Ischemic Optic Neuropathy, Laser Scar, Chorioretinitis, Within Normal Limit, Epiretinal Membrane, Central retinal vein occlusion, Central Serous Retinopathy, Optociliary Shunt an image of {} 4
DRD [75] No diabetic retinopathy, Mild diabetic retinopathy, Moderate diabetic retinopathy, Severe diabetic retinopathy, Proliferative diabetic retinopathy a detailed view of a retina indicating {};
a close-up of a retina highlighting {}
32
Covid-CXR2 [60] No Finding, Covid-19 a radiographic representation assessing for {} 256
NLM-TB [25] Normal, Tuberculosis a close-up view of a chest x-ray presenting evidence of {};
a visual analysis of a chest x-ray with signs of {};
a radiographic scan highlighting the presence of {};
an annotated image from a chest x-ray showing signs of {}
32
LC25000 (Colon) [61] normal colonic tissue, colon adenocarcinomas the histopathological image illustrates that of {};
a histopathological image featuring {} in colon;
a pathological image highlighting a {}
256
LC25000 (Lung) [61] lung squamous cell carcinomas, lung adenocarcinomas, normal lung tissue histopathological image contains {};
a pathological image highlighting a {};
an annotated histopathological image representing a {}
256
MedFMC (Chest) [71] pleural effusion, nodule, pneumonia, cardiomegaly, hilar enlargement, fracture old, fibrosis, aortic calcification, tortuous aorta, thickened pleura, TB, pneumothorax, emphysema, atelectasis, calcification, pulmonary edema, increased lung markings, elevated diaphragm, consolidation lung situation of {} 2
MedFMC (Colon) [71] tumor, normal this slide features an annotated section indicating a {};
the image illustrates a {} in the context of colon tissue
32
MedFMC (Endo) [71] ulcer, erosion, polyp, tumor a diagnostic endoscopy showing features of {};
an endoscopic finding suggestive of {}
32
VinDr-Mammo [72] birads negative, breast heterogeneously density, No Finding, breast scattered areas of fibroglandular, birads suspicious malignant, Mass, breast extremely density, birads highly suggestive of malignant, Suspicious Calcification, birads benign, breast almost entirely fatty, Suspicious Lymph Node, Focal Asymmetry, birads probably benign, Asymmetry, Architectural Distortion, Skin Thickening, Global Asymmetry, Nipple Retraction, Skin Retraction an image of {} 32
HAM10000 [76] actinic keratoses and intraepithelial carcinoma, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanoma, melanocytic nevi, vascular lesions a dermatology has that of {} presented 32
VinDr-PCXR [50] Bronchiolitis, Pneumonia, Other disease, Bronchitis, Brocho-pneumonia, Tuberculosis, Mediastinal tumor, nan, Situs inversus, Hyaline membrane disease, CPAM, No finding an image of {} 4
BUSBRA [77] metaplasia apocrina, ductal carcinoma in situ, lipoma, papillary carcinoma, hyperplasia, hamartoma, ductal hyperplasia, cyst, lymphoma, invasive ductal carcinoma, fibroadenoma, lymph node, sclerosing adenosis, medullary carcinoma, adenocarcinoma, intraductal papilloma, duct ectasia, mastitis, fibrosis, phyllodes tumor, lobular atrophy, lobular carcinoma in situ, fibrocystic changes, mucinous carcinoma, galactocele, invasive lobular carcinoma, proliferative lesions, fat necrosis this is an image of {};
{} presented in image
32
VinDr-CXR [49] Atelectasis, Lung tumor, Lung cavity, Tuberculosis, Pulmonary fibrosis, Clavicle fracture, Lung cyst, Pneumonia, Calcification, No finding, Rib fracture, Pleural thickening, Other disease, Mediastinal shift, Enlarged PA, Nodule/Mass, ILD, COPD, Pneumothorax, Consolidation, Infiltration, Pleural effusion, Other lesion, Cardiomegaly, Emphysema, Lung Opacity an image of {} 2
UBIBC [62] benign, malignant {} show in an ultrasound image of breast cancer;
an ultrasound image of breast cancer showing signs of {}
32
WCE [78] normal, ulcerative colitis, polyps, esophagitis an endoscopic finding suggestive of {} 32
PCam200 [63] normal, tumor histopathological image contains {};
an annotated histopathological image representing a {}
32
Fundus JSIEC [81] Massive hard exudates, Vitreous particles, Possible glaucoma, Yellow-white spots-flecks, CRVO, BRVO, Large optic cup, Blur fundus without PDR, Bietti crystalline dystrophy, Rhegmatogenous RD, CSCR, Fibrosis, Congenital disc abnormality, Laser Spots, Vessel tortuosity, Maculopathy, RAO, Pathological myopia, DR2, Tessellated fundus, Retinitis pigmentosa, ERM, VKH disease, MH, Cotton-wool spots, Optic atrophy, Severe hypertensive retinopathy, DR3, DR1, Preretinal hemorrhage, Fundus neoplasm, Silicon oil in eye, Blur fundus with suspected PDR, Myelinated nerve fiber, Normal, Disc swelling and elevation, Peripheral retinal degeneration and break, Chorioretinal atrophy-coloboma, Dragged Disc {} presented in image 32
HyperKvasir [82] Barrett’s, Barrett’s short segment, BBPS 0-1, BBPS 2-3, Cecum, Dyed lifted polyps, Dyed resection margins, Esophagitis A, Esophagitis B-D, Hemorrhoids, Terminal ileum, Impacted stool, Polyps, Pylorus, Retroflex rectum, Retroflex stomach, Ulcerative colitis 0-1, Ulcerative colitis 1, Ulcerative colitis 1-2, Ulcerative colitis 2, Ulcerative colitis 2-3, Ulcerative colitis 3, Z-line this is an image of {};
{} presented in image
32
Kvasir [79] dyed lifted polyps, dyed resection margins, esophagitis, normal cecum, normal pylorus, normal z line, polyps, ulcerative colitis an endoscopic finding suggestive of {};
a snapshot from an endoscopic procedure showing {}
32
RSNA [48] No Finding, Pneumonia the chest radiology image of {};
the lung radiology illustrates that of {}
8
VinDr-SpineXR [73] No finding, Osteophytes, Disc space narrowing, Surgical implant, Foraminal stenosis, Other lesions, Vertebral collapse, Spondylolysthesis a spine x-ray image of {} 2
NIH ChestX-ray14 [74] Atelectasis, Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumonia, Pneumothorax, Consolidation, Edema, Emphysema, Fibrosis, Pleural Thickening, Hernia, No Finding an image of {} 16
CheXpert [58] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pneumothorax, Pleural Effusion, Pleural Other, Fracture, Support Devices, No Finding an image of {} 4
ODIR [83] Normal, Diabetes, Glaucoma, Cataract, Age related Macular Degeneration, Hypertension, Pathological Myopia, Other diseases/abnormalities {} presented in retinography 32
BUID [84] normal, malignant, benign a close-up ultrasound scan showing a {} in the breast 32
PAD-UFES-20 [85] Basal Cell Carcinoma, Squamous Cell Carcinoma, Actinic Keratosis, Seborrheic Keratosis, Melanoma, Nevus a dermatological image of {};
a diagnostic dermatological image illustrating a {}
16
OCTMNIST [86] choroidal neovascularization, diabetic macular edema, drusen, normal an image of {} 4
Breast Tumor MRI [87] glioma, meningioma, pituitary, no tumor a breast mri image of {} 4
Retinal OCT [91] Age-related macular degeneration, Choroidal neovascularization, Central serous retinopathy, Diabetic macular edema, Macular hole, Drusen, Diabetic retinopathy, Normal a retinal oct image of {} 4
Brain Tumor CT [65] Healthy, Tumor an image of {} 4
COVIDxCT [92] Normal, Pneumonia, COVID-19 an image of {} 4
Brain Tumor MRI [65] Healthy, Tumor an image of {} 4
DDSM [66] negative, positive the mammography image of {} 4
Breast Cancer [69] Normal, Tumor the mammography image of {} 4
Table 3: AUC scores for classification results across different modalities in the zero-shot setting. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
X-Ray
SIIM-ACR 55.13 (50.86,59.26) 67.87 (64.13,71.31) 64.19 (60.43,67.68) 77.08 (73.96,80.15) 83.05 (80.40,85.63)
Covid-CXR2 49.40 (47.53,51.26) 70.34 (68.70,72.16) 57.41 (55.26,59.32) 68.99 (66.88,70.98) 81.77 (80.19,83.35)
NLM-TB 65.64 (55.98,73.96) 82.92 (76.45,88.80) 74.22 (65.71,81.75) 88.60 (82.44,94.08) 92.92 (88.48,96.69)
MedFMC (Chest) 49.81 (47.37,50.00) 61.80 (57.86,64.99) 49.85 (47.37,50.00) 50.20 (47.62,50.70) 66.54 (62.12,70.67)
VinDr-PCXR 42.40 (35.27,48.47) 43.38 (35.73,48.83) 45.69 (36.56,51.25) 48.02 (38.63,55.35) 43.72 (34.59,49.95)
VinDr-CXR 48.92 (45.78,50.09) 58.58 (53.25,61.85) 49.06 (46.05,50.01) 48.84 (45.46,50.03) 61.24 (55.13,65.68)
RSNA 48.98 (45.26,52.37) 76.66 (73.73,79.40) 82.01 (79.43,84.33) 82.23 (79.76,84.52) 87.01 (84.81,89.13)
VinDr-SpineXR 50.00 (50.00,50.00) 63.23 (61.36,65.04) 50.01 (50.00,50.03) 50.01 (50.00,50.02) 67.76 (66.00,69.38)
NIH ChestXray14 49.83 (49.72,49.95) 57.69 (57.19,58.18) 50.67 (50.60,50.73) 51.93 (51.66,52.18) 62.92 (62.37,63.47)
CheXpert 50.66 (50.59,50.74) 55.26 (55.06,55.47) 50.15 (50.13,50.17) 50.43 (50.38,50.49) 57.07 (56.86,57.28)
Average 51.08 (49.76,52.37) 63.77 (62.42,65.02) 57.33 (56.05,58.48) 61.63 (60.43,62.70) 70.40 (69.18,71.50)
Fundus
RFMiD2 42.47 (37.20,47.42) 45.71 (39.56,51.92) 41.40 (35.43,46.38) 41.11 (35.44,46.24) 58.79 (49.66,66.21)
DRD 57.79 (57.28,58.35) 59.86 (59.35,60.41) 63.99 (63.55,64.45) 66.25 (65.80,66.72) 73.27 (72.88,73.65)
Fundus JSIEC 49.46 (43.41,55.20) 60.37 (52.99,66.60) 65.54 (58.07,72.05) 66.17 (58.56,73.11) 73.65 (65.78,80.56)
ODIR 52.29 (51.11,53.46) 67.44 (66.29,68.51) 67.20 (66.27,68.25) 69.61 (68.64,70.63) 73.31 (72.25,74.39)
Average 50.50 (48.37,52.38) 58.34 (56.02,60.60) 59.53 (57.27,61.68) 60.78 (58.54,62.83) 69.76 (66.90,72.50)
Pathology
LC25000 (COLON) 66.53 (64.24,68.81) 88.31 (86.79,89.74) 98.78 (98.36,99.15) 98.89 (98.56,99.18) 99.29 (99.06,99.49)
LC25000 (LUNG) 73.73 (72.73,74.71) 74.09 (73.22,75.03) 94.01 (93.48,94.55) 88.58 (87.67,89.42) 98.51 (98.22,98.78)
MedFMC (Colon) 38.34 (36.72,40.12) 80.54 (79.18,81.87) 72.59 (71.11,74.00) 82.00 (80.67,83.24) 94.66 (94.04,95.30)
PCam200 59.87 (59.10,60.72) 71.87 (71.07,72.60) 79.44 (78.81,80.14) 83.16 (82.54,83.76) 92.43 (92.04,92.79)
Average 59.62 (58.86,60.42) 78.70 (78.08,79.28) 86.20 (85.79,86.61) 88.16 (87.75,88.59) 96.22 (96.02,96.42)
Endoscopy
MedFMC (Endo) 62.49 (60.07,64.70) 60.88 (58.51,63.01) 66.04 (64.02,68.07) 70.84 (69.05,72.59) 75.32 (73.15,77.17)
WCE 53.24 (51.04,55.51) 73.29 (71.30,75.32) 93.85 (92.71,94.92) 96.75 (96.00,97.46) 97.63 (96.93,98.23)
HyperKvasir 58.77 (53.30,63.33) 71.48 (66.23,75.59) 68.49 (62.45,72.11) 79.76 (74.34,83.75) 81.71 (76.91,85.37)
Kvasir 57.23 (55.82,58.66) 76.12 (75.01,77.23) 84.40 (83.44,85.39) 91.74 (90.99,92.44) 96.72 (96.27,97.13)
Average 57.93 (56.50,59.45) 70.44 (68.93,71.80) 78.19 (76.65,79.33) 84.77 (83.38,85.95) 87.84 (86.42,89.06)
Mammography
VinDr-Mammo 49.32 (46.94,50.19) 47.71 (44.19,51.04) 49.62 (47.21,50.29) 49.72 (47.52,50.07) 51.78 (48.25,55.05)
Breast Cancer 46.45 (40.13,53.16) 53.26 (46.50,59.91) 50.77 (44.59,57.32) 54.84 (49.00,61.19) 55.12 (48.97,61.62)
DDSM 31.35 (30.22,32.54) 60.74 (59.42,61.95) 57.21 (55.81,58.52) 52.50 (51.25,53.71) 77.16 (75.99,78.30)
Average 42.37 (40.16,44.56) 53.90 (51.22,56.43) 52.53 (50.30,54.84) 52.35 (50.13,54.58) 61.35 (59.08,63.77)
Dermoscopy
HAM10000 62.40 (59.99,64.55) 72.14 (70.16,74.08) 65.87 (63.52,67.97) 67.54 (65.49,69.48) 82.40 (80.91,83.96)
PAD-UFES-20 78.33 (75.56,80.91) 80.99 (78.58,83.34) 66.69 (63.10,70.06) 76.06 (72.83,79.12) 86.75 (84.67,88.73)
Average 70.36 (68.42,72.10) 76.56 (74.88,78.08) 66.28 (64.22,68.24) 71.80 (69.83,73.64) 84.58 (83.27,85.82)
Ultrasound
BUSBRA 44.05 (36.45,52.09) 41.57 (34.52,47.85) 45.24 (37.05,52.21) 47.76 (37.52,56.64) 54.09 (44.10,62.95)
UBIBC 36.12 (32.36,39.97) 78.37 (75.30,81.27) 58.47 (54.90,62.43) 70.75 (66.91,74.28) 90.65 (88.38,92.65)
BUID 54.00 (47.64,60.75) 70.47 (64.56,76.00) 69.86 (64.18,75.20) 77.45 (71.80,82.63) 79.75 (73.76,85.67)
Average 44.72 (40.99,48.40) 63.47 (60.27,66.57) 57.86 (54.45,61.18) 65.32 (61.45,69.05) 74.83 (71.13,78.56)
OCT
OCTMNIST 51.80 (49.82,53.74) 52.36 (50.44,54.10) 93.05 (92.14,93.91) 83.88 (82.38,85.32) 93.90 (93.02,94.75)
Retinal OCT 35.07 (34.24,35.88) 39.12 (38.18,39.98) 72.77 (72.05,73.45) 48.10 (47.30,48.89) 74.84 (74.25,75.42)
Average 43.44 (42.36,44.47) 45.74 (44.71,46.70) 82.91 (82.31,83.43) 65.99 (65.14,66.74) 84.37 (83.84,84.87)
MRI
Breast Tumor MRI 48.68 (46.77,50.52) 76.57 (75.22,77.81) 79.43 (77.94,80.92) 74.65 (73.19,76.09) 85.79 (84.45,87.05)
Brain Tumor MRI 94.37 (92.91,95.73) 99.04 (98.47,99.49) 97.97 (97.24,98.61) 98.90 (98.42,99.33) 99.69 (99.40,99.90)
Average 71.53 (70.27,72.64) 87.80 (87.10,88.48) 88.70 (87.89,89.52) 86.78 (86.00,87.54) 92.74 (92.06,93.37)
CT
Brain Tumor CT 59.94 (56.29,63.63) 88.55 (86.45,90.51) 74.96 (71.77,78.01) 74.92 (71.69,78.05) 92.60 (90.76,94.20)
COVIDxCT 50.74 (50.34,51.15) 66.63 (66.26,67.00) 83.13 (82.76,83.49) 85.58 (85.26,85.90) 89.36 (89.03,89.67)
Average 55.34 (53.50,57.17) 77.59 (76.53,78.57) 79.04 (77.42,80.56) 80.25 (78.64,81.77) 90.98 (90.00,91.82)
Table 4: AUC scores for classification results across different modalities in the linear probes setting with 1% training data. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
X-Ray
SIIM-ACR 69.96 (66.47,73.58) 79.29 (76.32,82.24) 74.66 (71.40,77.90) 73.04 (69.62,76.06) 88.38 (86.15,90.42)
Covid-CXR2 96.29 (95.73,96.84) 93.39 (92.62,94.07) 90.56 (89.65,91.43) 88.27 (87.12,89.40) 87.52 (86.42,88.68)
NLM-TB 75.80 (67.51,83.60) 65.80 (56.69,74.00) 62.22 (53.17,71.13) 54.10 (44.33,62.72) 68.15 (59.59,76.48)
MedFMC (Chest) 53.49 (50.90,55.39) 56.52 (53.81,58.49) 60.81 (58.13,62.55) 61.28 (58.21,63.09) 62.94 (60.42,64.54)
VinDr-PCXR 46.80 (39.07,52.96) 47.32 (39.54,53.54) 47.24 (39.63,53.31) 48.01 (40.18,53.93) 47.74 (40.15,53.86)
VinDr-CXR 52.42 (49.12,54.31) 63.31 (59.90,65.25) 58.91 (55.17,60.90) 64.35 (60.68,66.07) 68.39 (65.06,70.13)
RSNA 73.93 (71.03,76.88) 82.35 (79.92,84.89) 82.80 (80.29,85.08) 84.86 (82.45,87.12) 85.52 (83.44,87.58)
VinDr-SpineXR 48.52 (46.76,50.31) 48.62 (46.66,50.51) 46.65 (44.67,48.62) 48.92 (47.04,50.55) 49.19 (47.47,51.04)
NIH ChestXray14 60.43 (59.93,60.96) 63.36 (62.73,64.01) 67.38 (66.96,67.84) 67.74 (67.20,68.29) 71.99 (71.50,72.46)
CheXpert 71.51 (71.29,71.73) 77.31 (77.07,77.54) 80.77 (80.53,81.08) 81.87 (81.40,82.10) 84.96 (84.75,85.19)
Average 64.91 (63.50,66.13) 67.73 (66.36,68.97) 67.20 (65.77,68.48) 67.24 (65.84,68.48) 71.48 (70.19,72.70)
Fundus
RFMiD2 43.01 (37.28,48.13) 42.83 (37.35,47.95) 43.00 (37.43,48.06) 43.11 (37.02,48.51) 42.79 (37.17,47.82)
DRD 66.50 (66.02,66.98) 68.00 (67.57,68.49) 69.43 (68.95,69.88) 67.67 (67.22,68.12) 72.08 (71.66,72.48)
Average 54.76 (25.96,28.67) 55.42 (26.34,28.92) 56.22 (26.73,29.37) 55.39 (26.20,29.03) 57.44 (27.32,29.96)
Pathology
LC25000 (COLON) 99.09 (98.81,99.34) 99.02 (98.71,99.30) 99.93 (99.89,99.97) 99.88 (99.81,99.94) 99.90 (99.84,99.94)
LC25000 (LUNG) 96.57 (96.16,96.97) 97.72 (97.32,98.05) 99.36 (99.20,99.50) 98.47 (98.20,98.73) 99.52 (99.40,99.63)
MedFMC (Colon) 83.69 (82.57,84.80) 94.73 (94.05,95.30) 96.06 (95.51,96.54) 96.72 (96.23,97.16) 98.25 (97.88,98.61)
PCam200 79.47 (78.84,80.14) 87.50 (87.00,88.02) 88.44 (87.98,88.92) 90.19 (89.77,90.63) 95.69 (95.40,95.96)
Average 89.71 (89.36,90.04) 94.74 (94.50,94.96) 95.95 (95.77,96.12) 96.32 (96.14,96.50) 98.34 (98.23,98.45)
Endoscopy
WCE 88.13 (87.09,89.20) 97.43 (96.86,97.96) 97.32 (96.77,97.80) 98.07 (97.63,98.51) 98.45 (97.94,98.85)
Kvasir 92.10 (91.62,92.58) 95.12 (94.81,95.43) 96.13 (95.83,96.41) 95.92 (95.66,96.18) 97.48 (97.24,97.71)
Average 90.12 (44.76,45.35) 96.28 (47.97,48.29) 96.72 (48.21,48.51) 97.00 (48.37,48.64) 97.96 (48.84,49.10)
Mammography
VinDr-Mammo 54.03 (51.10,56.14) 54.25 (51.25,56.29) 57.07 (54.09,58.87) 55.28 (52.45,57.22) 55.35 (52.13,57.56)
Breast Cancer 50.08 (43.77,56.59) 54.37 (48.06,61.13) 54.85 (48.57,61.26) 55.71 (49.71,62.06) 57.30 (51.00,63.43)
DDSM 90.53 (89.82,91.21) 86.17 (85.41,86.91) 93.32 (92.79,93.79) 88.44 (87.70,89.18) 90.64 (89.98,91.30)
Average 64.88 (62.67,67.21) 64.93 (62.58,67.49) 68.41 (66.11,70.72) 66.48 (64.30,68.77) 67.76 (65.45,70.09)
Dermoscopy
HAM10000 75.44 (73.35,77.50) 79.87 (78.05,81.52) 73.79 (71.76,75.79) 78.90 (77.30,80.52) 83.35 (82.13,84.63)
Average 75.44 (36.67,38.75) 79.87 (39.03,40.76) 73.79 (35.88,37.90) 78.90 (38.65,40.26) 83.35 (41.06,42.31)
Ultrasound
UBIBC 67.23 (63.00,71.19) 73.01 (69.52,76.57) 70.06 (66.24,73.68) 75.58 (72.45,78.82) 78.06 (74.79,81.20)
Average 67.23 (21.00,23.73) 73.01 (23.17,25.52) 70.06 (22.08,24.56) 75.58 (24.15,26.27) 78.06 (24.93,27.07)
OCT
OCTMNIST 87.50 (86.25,88.66) 90.03 (88.94,91.05) 97.69 (97.13,98.21) 91.55 (90.50,92.57) 97.64 (97.11,98.14)
Retinal OCT 81.27 (80.85,81.67) 81.28 (80.91,81.63) 84.73 (84.46,84.98) 77.71 (77.21,78.18) 84.24 (83.96,84.51)
Average 84.38 (83.74,85.02) 85.66 (85.09,86.21) 91.21 (90.92,91.49) 84.63 (84.05,85.21) 90.94 (90.65,91.21)
MRI
Breast Tumor MRI 85.05 (84.10,86.05) 92.65 (91.95,93.28) 94.51 (93.88,95.12) 93.53 (92.89,94.22) 96.25 (95.69,96.81)
Brain Tumor MRI 96.94 (96.06,97.79) 99.87 (99.76,99.95) 99.62 (99.39,99.79) 99.79 (99.60,99.92) 99.79 (99.55,99.95)
Average 91.00 (90.31,91.67) 96.26 (95.91,96.59) 97.06 (96.73,97.39) 96.66 (96.33,97.01) 98.02 (97.73,98.31)
CT
Brain Tumor CT 94.44 (92.53,96.06) 96.95 (95.91,97.86) 98.30 (97.34,99.05) 97.57 (96.68,98.42) 98.94 (98.28,99.49)
COVIDxCT 69.56 (69.15,69.97) 75.14 (74.75,75.52) 92.98 (92.76,93.20) 93.46 (93.24,93.66) 95.31 (95.13,95.50)
Average 82.00 (81.01,82.84) 86.04 (85.52,86.55) 95.64 (95.16,96.01) 95.51 (95.06,95.96) 97.12 (96.78,97.42)
Table 5: AUC scores for classification results across different modalities in the linear probes setting with 10% training data. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
X-Ray
SIIM-ACR 71.58 (68.30,74.92) 80.05 (76.99,82.91) 78.56 (75.52,81.38) 81.24 (78.42,83.90) 88.73 (86.56,90.78)
Covid-CXR2 97.29 (96.80,97.76) 96.44 (95.93,96.93) 95.07 (94.46,95.66) 92.61 (91.75,93.46) 92.61 (91.77,93.37)
NLM-TB 71.82 (63.16,79.77) 86.91 (81.10,91.73) 79.61 (72.06,86.45) 87.02 (81.32,92.16) 95.11 (91.00,98.03)
MedFMC (Chest) 59.73 (55.85,62.79) 67.63 (63.81,70.43) 71.55 (68.03,74.18) 72.82 (69.32,75.39) 73.87 (70.22,77.31)
VinDr-PCXR 51.10 (41.11,59.05) 47.69 (39.61,52.36) 50.72 (42.99,56.39) 50.30 (41.55,56.65) 47.97 (40.30,52.75)
VinDr-CXR 64.66 (59.72,67.22) 73.96 (68.81,76.30) 71.84 (66.61,74.25) 74.59 (69.22,76.80) 78.26 (73.81,81.06)
RSNA 75.40 (72.27,78.18) 83.39 (81.18,85.58) 84.76 (82.47,87.00) 85.91 (83.70,88.01) 86.50 (84.14,88.51)
VinDr-SpineXR 45.70 (44.02,47.53) 45.71 (43.86,47.59) 45.03 (43.04,46.92) 44.56 (42.74,46.29) 46.71 (44.86,48.55)
NIH ChestXray14 63.02 (62.48,63.57) 68.44 (67.92,68.92) 70.44 (69.95,70.86) 71.35 (70.87,71.82) 75.23 (74.78,75.69)
CheXpert 74.91 (74.66,75.18) 80.39 (80.20,80.70) 82.85 (82.60,83.06) 83.17 (82.99,83.36) 86.38 (86.12,86.58)
Average 67.52 (65.92,68.92) 73.06 (71.97,74.06) 73.04 (71.66,74.15) 74.36 (73.07,75.49) 77.14 (76.12,78.01)
Fundus
RFMiD2 52.00 (45.90,57.10) 54.24 (48.07,59.71) 55.35 (49.06,60.51) 54.38 (48.24,59.82) 55.33 (48.78,60.80)
DRD 71.71 (71.30,72.12) 72.71 (72.29,73.10) 73.21 (72.79,73.64) 69.81 (69.36,70.29) 76.44 (76.09,76.83)
ODIR 65.22 (64.28,66.22) 68.40 (67.44,69.43) 68.94 (67.98,69.81) 68.23 (67.27,69.23) 72.09 (71.16,73.06)
Average 62.98 (45.70,48.54) 65.12 (47.31,50.26) 65.83 (47.82,50.69) 64.14 (46.51,49.51) 67.95 (49.35,52.29)
Pathology
LC25000 (COLON) 99.53 (99.35,99.69) 99.32 (99.10,99.53) 99.97 (99.95,99.99) 99.92 (99.87,99.96) 99.94 (99.91,99.97)
LC25000 (LUNG) 97.62 (97.24,97.99) 98.68 (98.43,98.94) 99.45 (99.30,99.58) 98.70 (98.43,98.94) 99.76 (99.69,99.83)
MedFMC (Colon) 87.96 (86.97,89.00) 94.54 (93.88,95.14) 97.82 (97.47,98.15) 97.44 (97.03,97.83) 98.36 (98.03,98.68)
PCam200 83.33 (82.75,83.88) 89.41 (88.96,89.85) 91.31 (90.90,91.70) 91.80 (91.39,92.19) 96.01 (95.75,96.27)
Average 92.11 (91.80,92.42) 95.49 (95.27,95.69) 97.14 (97.00,97.27) 96.96 (96.81,97.11) 98.52 (98.42,98.64)
Endoscopy
MedFMC (Endo) 72.95 (70.73,75.02) 75.20 (73.07,77.38) 66.15 (63.98,68.49) 72.43 (70.17,74.68) 78.28 (76.49,80.17)
WCE 94.85 (94.04,95.68) 99.02 (98.63,99.35) 99.12 (98.85,99.35) 99.13 (98.83,99.38) 99.54 (99.30,99.75)
Kvasir 95.85 (95.53,96.17) 96.77 (96.49,97.05) 97.96 (97.72,98.17) 97.69 (97.47,97.93) 98.76 (98.59,98.93)
Average 87.88 (65.32,66.49) 90.33 (67.19,68.29) 87.74 (65.24,66.37) 89.75 (66.74,67.88) 92.19 (68.68,69.63)
Mammography
VinDr-Mammo 54.61 (50.35,56.95) 55.54 (52.68,58.10) 58.94 (55.46,61.03) 56.31 (53.33,58.62) 59.47 (55.55,61.82)
Breast Cancer 56.13 (49.88,62.00) 59.11 (53.13,65.52) 57.93 (51.95,64.30) 56.08 (49.37,62.08) 59.39 (52.80,65.89)
DDSM 92.96 (92.47,93.46) 92.64 (92.12,93.17) 95.28 (94.82,95.68) 91.93 (91.33,92.49) 93.23 (92.63,93.75)
Average 67.90 (65.36,70.04) 69.10 (66.98,71.27) 70.72 (68.26,73.00) 68.11 (65.87,70.43) 70.70 (68.07,72.99)
Dermoscopy
HAM10000 83.48 (81.85,85.13) 85.28 (83.84,86.68) 82.21 (80.28,83.92) 84.39 (83.04,85.64) 87.33 (86.22,88.42)
PAD-UFES-20 79.24 (77.06,81.44) 84.40 (82.51,86.22) 76.90 (74.21,79.36) 81.24 (79.08,83.34) 83.67 (81.74,85.58)
Average 81.36 (80.02,82.72) 84.84 (83.69,86.05) 79.56 (77.94,81.05) 82.82 (81.57,84.12) 85.50 (84.37,86.70)
Ultrasound
UBIBC 72.13 (68.34,75.67) 78.54 (75.34,81.76) 83.96 (81.09,86.89) 82.85 (80.04,85.71) 88.94 (86.47,91.35)
BUID 79.54 (73.49,85.26) 82.50 (76.92,87.17) 80.19 (74.46,85.49) 85.09 (79.95,89.63) 85.32 (79.54,90.47)
Average 75.84 (48.20,52.83) 80.52 (51.60,55.64) 82.07 (52.65,56.74) 83.97 (53.98,57.85) 87.13 (55.95,59.98)
OCT
OCTMNIST 94.28 (93.37,95.14) 94.95 (94.05,95.72) 98.66 (98.25,99.01) 95.81 (95.05,96.61) 99.19 (98.86,99.50)
Retinal OCT 83.66 (83.34,83.97) 83.92 (83.66,84.19) 86.17 (85.98,86.35) 82.63 (82.25,82.98) 86.06 (85.86,86.23)
Average 88.97 (88.51,89.46) 89.44 (88.97,89.88) 92.42 (92.20,92.62) 89.22 (88.81,89.65) 92.62 (92.44,92.81)
MRI
Breast Tumor MRI 89.81 (88.81,90.83) 97.06 (96.53,97.53) 98.12 (97.69,98.52) 96.36 (95.79,96.89) 98.47 (98.09,98.80)
Brain Tumor MRI 97.60 (96.70,98.34) 99.95 (99.89,99.99) 99.84 (99.68,99.95) 99.81 (99.63,99.95) 99.78 (99.46,99.96)
Average 93.70 (93.02,94.34) 98.50 (98.23,98.74) 98.98 (98.76,99.19) 98.08 (97.78,98.37) 99.12 (98.88,99.33)
CT
Brain Tumor CT 95.44 (93.82,96.88) 98.55 (97.86,99.13) 98.58 (97.68,99.32) 98.68 (98.04,99.20) 99.46 (99.11,99.76)
COVIDxCT 88.09 (87.81,88.39) 91.40 (91.16,91.65) 96.82 (96.67,96.96) 95.30 (95.11,95.49) 97.35 (97.22,97.48)
Average 91.76 (90.94,92.55) 94.98 (94.60,95.29) 97.70 (97.22,98.09) 96.99 (96.67,97.27) 98.40 (98.22,98.57)
Table 6: AUC scores for classification results across different modalities in the linear probes setting with 100% training data. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
X-Ray
SIIM-ACR 74.43 (70.89,77.61) 83.33 (80.66,86.03) 83.45 (80.62,86.06) 85.49 (83.02,87.95) 90.12 (88.08,92.20)
Covid-CXR2 98.24 (97.87,98.63) 98.74 (98.46,99.00) 98.11 (97.75,98.45) 96.07 (95.50,96.60) 96.94 (96.46,97.37)
NLM-TB 70.02 (61.77,78.29) 87.38 (81.15,92.84) 79.54 (71.57,86.50) 88.69 (83.13,93.07) 95.05 (91.50,98.18)
MedFMC (Chest) 66.28 (62.32,69.67) 73.40 (68.56,76.44) 76.33 (72.02,78.66) 76.31 (72.24,79.53) 79.40 (75.04,82.54)
VinDr-PCXR 61.39 (48.30,70.05) 63.77 (50.39,72.24) 67.28 (53.24,78.07) 63.59 (49.15,71.75) 63.89 (51.40,72.75)
VinDr-CXR 69.67 (63.36,73.04) 77.36 (71.45,81.09) 77.74 (72.06,81.40) 79.45 (73.26,82.37) 82.72 (75.55,86.44)
RSNA 78.38 (75.77,81.07) 83.96 (81.50,86.30) 85.56 (83.19,87.62) 86.46 (84.26,88.52) 87.52 (85.32,89.50)
VinDr-SpineXR 49.08 (47.20,50.86) 48.98 (47.28,50.63) 49.87 (48.04,51.62) 46.88 (45.33,48.47) 48.95 (47.09,50.52)
NIH ChestXray14 61.87 (61.35,62.42) 67.23 (66.61,67.79) 71.07 (70.59,71.56) 70.89 (70.39,71.41) 74.51 (73.96,75.08)
CheXpert 77.93 (77.70,78.20) 82.83 (82.62,83.01) 83.99 (83.84,84.16) 83.73 (83.54,83.92) 87.07 (86.90,87.22)
Average 70.73 (69.01,72.22) 76.70 (75.00,78.06) 77.29 (75.44,78.96) 77.76 (76.12,79.01) 80.62 (79.14,81.92)
Fundus
RFMiD2 65.63 (56.12,74.03) 68.54 (59.41,77.72) 70.89 (61.58,79.48) 67.71 (59.15,76.27) 70.25 (61.13,78.72)
DRD 77.04 (76.69,77.40) 77.93 (77.59,78.28) 76.72 (76.31,77.09) 73.17 (72.75,73.60) 79.42 (79.06,79.77)
Fundus JSIEC 83.61 (74.84,91.30) 84.69 (75.60,92.60) 87.31 (77.79,95.39) 84.34 (75.20,92.11) 87.29 (78.36,95.72)
ODIR 72.23 (71.29,73.14) 74.55 (73.49,75.54) 76.46 (75.42,77.52) 73.05 (71.92,74.13) 77.89 (77.05,78.88)
Average 74.63 (71.52,77.53) 76.43 (73.24,79.35) 77.84 (74.66,80.87) 74.57 (71.45,77.52) 78.71 (75.73,81.79)
Pathology
LC25000 (COLON) 99.89 (99.81,99.95) 99.87 (99.80,99.93) 100.00 (99.99,100.00) 99.98 (99.96,99.99) 99.99 (99.98,100.00)
LC25000 (LUNG) 98.76 (98.49,99.00) 99.31 (99.13,99.48) 99.73 (99.64,99.81) 99.38 (99.21,99.52) 99.88 (99.83,99.92)
MedFMC (Colon) 93.07 (92.34,93.80) 97.01 (96.55,97.44) 98.72 (98.47,98.96) 98.08 (97.74,98.39) 98.87 (98.59,99.11)
PCam200 87.42 (86.90,87.92) 92.02 (91.61,92.42) 93.00 (92.62,93.35) 92.57 (92.18,92.95) 96.54 (96.28,96.78)
Average 94.79 (94.56,95.01) 97.05 (96.89,97.21) 97.86 (97.75,97.97) 97.50 (97.36,97.64) 98.82 (98.73,98.90)
Endoscopy
MedFMC (Endo) 78.27 (76.30,80.01) 78.69 (76.55,80.70) 74.45 (72.52,76.34) 77.29 (75.22,79.17) 81.37 (79.77,83.03)
WCE 97.90 (97.33,98.47) 99.68 (99.44,99.85) 99.73 (99.59,99.86) 99.63 (99.41,99.80) 99.82 (99.69,99.93)
HyperKvasir 91.97 (87.27,95.64) 92.81 (85.20,96.53) 93.10 (87.25,97.45) 93.03 (85.40,97.24) 93.74 (88.64,97.30)
Kvasir 98.25 (98.03,98.47) 98.83 (98.65,99.01) 99.00 (98.81,99.17) 98.87 (98.67,99.05) 99.40 (99.25,99.52)
Average 91.60 (90.13,92.75) 92.50 (90.71,93.67) 91.57 (89.88,92.83) 92.21 (90.35,93.51) 93.58 (91.98,94.73)
Mammography
VinDr-Mammo 58.87 (54.64,61.87) 59.95 (56.46,62.99) 60.95 (56.99,64.00) 58.07 (54.17,61.78) 64.71 (59.69,67.13)
Breast Cancer 57.46 (50.82,63.71) 58.13 (51.52,64.32) 59.67 (52.95,66.36) 57.20 (50.79,63.63) 61.17 (54.44,67.31)
DDSM 95.95 (95.60,96.32) 96.26 (95.88,96.62) 96.91 (96.57,97.21) 94.04 (93.52,94.50) 96.32 (95.93,96.70)
Average 70.76 (68.20,73.15) 71.45 (69.01,73.81) 72.51 (69.98,74.95) 69.77 (67.32,72.11) 74.07 (71.50,76.37)
Dermoscopy
HAM10000 90.32 (89.11,91.48) 92.63 (91.70,93.47) 90.75 (89.66,91.82) 88.97 (87.78,90.12) 92.12 (91.22,92.98)
PAD-UFES-20 88.16 (86.33,90.01) 90.09 (88.45,91.80) 85.47 (83.30,87.61) 87.33 (85.21,89.41) 89.90 (88.21,91.53)
Average 89.24 (88.13,90.29) 91.36 (90.39,92.30) 88.11 (86.99,89.29) 88.15 (86.96,89.37) 91.01 (90.10,91.88)
Ultrasound
BUSBRA 57.73 (46.98,67.09) 59.93 (49.43,69.44) 63.07 (52.47,73.60) 59.69 (49.37,69.26) 64.79 (53.56,75.07)
UBIBC 83.70 (80.68,86.40) 88.27 (85.96,90.55) 91.06 (89.04,93.16) 87.13 (84.49,89.59) 93.44 (91.73,94.98)
BUID 84.76 (79.68,89.26) 86.60 (81.51,91.44) 88.95 (83.99,93.01) 87.22 (82.40,91.57) 89.18 (83.99,93.70)
Average 75.40 (71.38,79.02) 78.27 (74.47,82.00) 81.03 (77.08,84.80) 78.01 (73.92,81.59) 82.47 (78.48,86.15)
OCT
OCTMNIST 96.83 (96.14,97.49) 98.06 (97.56,98.50) 98.17 (97.69,98.60) 96.77 (95.99,97.53) 99.47 (99.23,99.70)
Retinal OCT 85.31 (85.07,85.55) 85.61 (85.39,85.82) 86.64 (86.49,86.77) 84.55 (84.26,84.82) 86.73 (86.57,86.87)
Average 91.07 (90.69,91.44) 91.84 (91.57,92.08) 92.40 (92.16,92.62) 90.66 (90.23,91.06) 93.10 (92.96,93.23)
MRI
Breast Tumor MRI 94.30 (93.48,95.11) 98.58 (98.21,98.94) 99.23 (98.99,99.45) 97.94 (97.50,98.35) 99.17 (98.90,99.40)
Brain Tumor MRI 98.34 (97.64,98.86) 99.96 (99.91,99.99) 99.91 (99.80,99.98) 99.87 (99.72,99.97) 99.87 (99.67,99.99)
Average 96.32 (95.77,96.83) 99.27 (99.08,99.45) 99.57 (99.43,99.69) 98.90 (98.67,99.12) 99.52 (99.36,99.66)
CT
Brain Tumor CT 97.05 (95.76,98.18) 99.13 (98.65,99.54) 98.85 (98.04,99.52) 99.13 (98.68,99.49) 99.61 (99.34,99.84)
COVIDxCT 93.87 (93.67,94.07) 96.46 (96.31,96.62) 97.59 (97.47,97.71) 96.14 (95.98,96.31) 98.11 (97.99,98.22)
Average 95.46 (94.79,96.03) 97.79 (97.56,98.01) 98.22 (97.82,98.56) 97.64 (97.40,97.84) 98.86 (98.71,98.99)
Table 7: AUC scores for classification results across different modalities in the fully fine-tuning setting. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
X-Ray
SIIM-ACR 88.29 (86.07,90.45) 89.46 (87.19,91.55) 90.05 (87.94,92.09) 90.59 (88.56,92.47) 91.72 (89.82,93.58)
Covid-CXR2 99.76 (99.53,99.91) 99.93 (99.85,99.98) 99.76 (99.57,99.91) 99.85 (99.66,99.97) 99.89 (99.77,99.98)
NLM-TB 71.37 (63.30,79.07) 92.83 (88.02,97.04) 87.07 (81.66,92.69) 92.64 (87.29,97.12) 94.32 (90.11,97.65)
MedFMC (Chest) 75.70 (71.89,78.92) 73.56 (69.91,76.59) 72.50 (68.61,75.99) 74.41 (70.19,78.33) 82.41 (78.08,84.86)
VinDr-PCXR 57.89 (46.23,66.38) 65.14 (50.66,74.33) 59.53 (47.55,68.18) 60.69 (49.10,70.10) 67.25 (53.99,76.13)
VinDr-CXR 62.09 (56.20,67.83) 70.30 (64.01,74.03) 58.00 (53.27,62.36) 71.43 (65.68,75.29) 83.47 (77.23,86.20)
RSNA 83.75 (81.42,86.01) 84.05 (81.65,86.36) 84.73 (82.36,86.92) 82.71 (80.14,85.15) 86.21 (83.86,88.27)
VinDr-SpineXR 51.28 (49.25,53.24) 51.16 (49.35,52.96) 53.55 (51.78,55.19) 52.06 (50.11,54.16) 54.15 (52.42,55.86)
NIH ChestXray14 71.21 (70.68,71.75) 78.91 (78.43,79.35) 73.17 (72.68,73.62) 70.67 (70.18,71.17) 78.99 (78.51,79.43)
CheXpert 89.29 (89.14,89.42) 89.34 (89.17,89.49) 87.85 (87.67,88.06) 87.56 (87.38,87.75) 89.16 (88.97,89.35)
Average 75.06 (73.43,76.46) 79.47 (77.85,80.81) 76.62 (75.21,77.89) 78.26 (76.77,79.69) 82.76 (81.29,83.99)
Fundus
RFMiD2 60.81 (51.87,68.51) 58.34 (50.65,64.88) 63.39 (54.47,71.84) 63.42 (54.52,71.52) 69.09 (60.30,77.22)
DRD 71.44 (70.84,72.00) 83.51 (83.15,83.89) 82.55 (82.15,82.94) 77.31 (76.84,77.74) 80.35 (79.95,80.72)
Fundus JSIEC 88.99 (79.43,97.36) 87.74 (78.07,95.87) 88.25 (78.95,96.58) 88.72 (79.15,97.00) 88.51 (79.13,96.62)
ODIR 76.77 (75.72,77.86) 81.24 (80.29,82.14) 81.07 (80.07,82.06) 79.08 (78.02,80.05) 83.37 (82.52,84.28)
Average 74.50 (71.18,77.17) 77.71 (74.71,80.47) 78.82 (75.75,81.74) 77.13 (73.99,80.13) 80.33 (77.28,83.03)
Pathology
LC25000 (COLON) 100.00 (100.00,100.00) 100.00 (100.00,100.00) 100.00 (100.00,100.00) 100.00 (100.00,100.00) 100.00 (100.00,100.00)
LC25000 (LUNG) 100.00 (100.00,100.00) 100.00 (100.00,100.00) 100.00 (100.00,100.00) 100.00 (100.00,100.00) 100.00 (100.00,100.00)
MedFMC (Colon) 98.27 (97.72,98.76) 99.20 (98.93,99.44) 99.05 (98.83,99.26) 99.49 (99.32,99.64) 99.50 (99.34,99.65)
PCam200 94.74 (94.39,95.05) 97.33 (97.09,97.56) 97.15 (96.91,97.38) 96.46 (96.17,96.73) 96.53 (96.25,96.80)
Average 98.25 (98.10,98.40) 99.13 (99.04,99.22) 99.05 (98.97,99.13) 98.99 (98.91,99.07) 99.01 (98.93,99.09)
Endoscopy
MedFMC (Endo) 70.87 (67.35,73.98) 80.04 (77.46,82.28) 73.15 (71.06,75.22) 77.02 (75.10,78.76) 79.20 (77.36,81.10)
WCE 99.39 (99.07,99.64) 99.80 (99.45,100.00) 99.42 (99.16,99.65) 99.91 (99.80,99.98) 99.85 (99.54,100.00)
HyperKvasir 93.69 (85.63,98.23) 94.52 (89.23,98.66) 93.89 (85.75,98.14) 93.34 (87.49,98.00) 94.72 (85.95,98.77)
Kvasir 99.52 (99.39,99.64) 99.08 (98.78,99.36) 99.49 (99.32,99.65) 99.33 (99.07,99.54) 99.34 (99.10,99.54)
Average 90.87 (88.94,92.49) 93.36 (91.42,94.77) 91.49 (89.58,92.89) 92.40 (90.59,93.73) 93.28 (91.40,94.55)
Mammography
VinDr-Mammo 61.20 (57.57,63.62) 68.56 (64.11,71.23) 62.58 (57.52,65.34) 62.56 (58.45,64.92) 66.38 (61.33,68.44)
Breast Cancer 59.89 (53.75,66.21) 75.68 (70.36,80.99) 76.40 (70.96,81.49) 72.09 (66.49,77.50) 80.46 (75.72,84.88)
DDSM 99.54 (99.42,99.64) 99.40 (99.16,99.59) 99.41 (99.18,99.60) 99.47 (99.31,99.60) 98.29 (97.81,98.69)
Average 73.54 (71.40,75.88) 81.21 (79.00,83.25) 79.46 (77.19,81.49) 78.04 (75.68,80.04) 81.71 (79.44,83.47)
Dermoscopy
HAM10000 93.60 (92.07,94.88) 96.97 (95.83,97.86) 95.39 (94.57,96.14) 94.79 (93.52,95.86) 96.47 (95.75,97.14)
PAD-UFES-20 91.11 (88.20,93.46) 93.61 (91.97,95.07) 91.59 (89.58,93.41) 90.90 (89.06,92.80) 92.92 (90.99,94.70)
Average 92.35 (90.67,93.75) 95.29 (94.26,96.18) 93.49 (92.42,94.49) 92.84 (91.73,93.93) 94.70 (93.66,95.66)
Ultrasound
BUSBRA 70.66 (58.76,81.79) 67.01 (55.81,77.43) 72.35 (59.85,83.70) 70.69 (57.44,81.85) 72.08 (60.57,83.76)
UBIBC 99.90 (99.82,99.96) 99.80 (99.43,99.99) 99.78 (99.61,99.90) 99.81 (99.58,99.96) 99.99 (99.96,100.00)
BUID 95.00 (91.95,97.58) 94.21 (90.90,96.68) 93.44 (89.84,96.68) 95.91 (93.60,97.73) 94.97 (92.37,97.19)
Average 88.52 (84.28,92.28) 87.01 (83.11,90.69) 88.52 (84.39,92.52) 88.80 (84.37,92.55) 89.01 (85.18,92.95)
OCT
OCTMNIST 98.87 (98.36,99.32) 99.13 (98.74,99.45) 98.69 (98.21,99.12) 98.44 (97.79,99.00) 99.17 (98.76,99.49)
Retinal OCT 86.56 (86.37,86.74) 86.19 (85.95,86.42) 86.53 (86.37,86.67) 84.05 (83.70,84.35) 86.70 (86.58,86.81)
Average 92.72 (92.43,92.95) 92.66 (92.45,92.86) 92.61 (92.36,92.83) 91.24 (90.89,91.58) 92.94 (92.73,93.11)
MRI
Breast Tumor MRI 99.99 (99.96,100.00) 99.96 (99.91,100.00) 100.00 (99.99,100.00) 99.98 (99.94,100.00) 100.00 (99.99,100.00)
Brain Tumor MRI 99.99 (99.96,100.00) 99.99 (99.96,100.00) 100.00 (100.00,100.00) 100.00 (99.99,100.00) 100.00 (99.99,100.00)
Average 99.99 (99.97,100.00) 99.98 (99.95,100.00) 100.00 (99.99,100.00) 99.99 (99.97,100.00) 100.00 (99.99,100.00)
CT
Brain Tumor CT 99.98 (99.94,100.00) 100.00 (99.99,100.00) 99.93 (99.81,100.00) 99.99 (99.98,100.00) 100.00 (99.99,100.00)
COVIDxCT 97.15 (97.01,97.29) 96.99 (96.84,97.13) 98.87 (98.78,98.95) 97.13 (97.00,97.25) 99.30 (99.24,99.36)
Average 98.56 (98.49,98.63) 98.50 (98.42,98.57) 99.40 (99.33,99.46) 98.56 (98.50,98.62) 99.65 (99.62,99.68)
Table 8: Performance of Recall@1,5,10 metrics of different models on the PMC-9K dataset. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Image-to-Text
Model Recall@1 Recall@5 Recall@10
CLIP 5.63 (5.18,6.10) 12.47 (11.80,13.15) 17.14 (16.38,17.92)
SigLIP-B 3.38 (3.04,3.77) 8.25 (7.69,8.79) 11.66 (10.98,12.29)
SigLIP-400M 2.86 (2.54,3.23) 7.04 (6.53,7.59) 9.68 (9.09,10.37)
PMC-CLIP 15.77 (15.05,16.49) 34.33 (33.34,35.37) 44.35 (43.35,45.35)
BiomedCLIP 73.41 (72.53,74.32) 91.93 (91.40,92.46) 95.30 (94.88,95.75)
ConceptCLIP 82.85 (82.05,83.60) 94.71 (94.23,95.20) 97.01 (96.61,97.38)
Text-to-Image
Model Recall@1 Recall@5 Recall@10
CLIP 6.31 (5.79,6.83) 13.11 (12.44,13.86) 18.39 (17.63,19.24)
SigLIP-B 3.29 (2.94,3.64) 8.06 (7.54,8.57) 11.68 (11.04,12.31)
SigLIP-400M 2.65 (2.32,2.97) 6.96 (6.47,7.47) 9.82 (9.22,10.40)
PMC-CLIP 15.69 (14.96,16.43) 32.93 (31.96,33.82) 42.71 (41.67,43.69)
BiomedCLIP 74.02 (73.15,74.93) 92.06 (91.49,92.65) 95.38 (94.96,95.80)
ConceptCLIP 83.24 (82.51,83.96) 94.50 (94.05,94.99) 96.86 (96.50,97.21)
Table 9: Performance of Recall@1,50,200 metrics of different models on the QUILT-1M dataset. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Image-to-Text
Model Recall@1 Recall@50 Recall@200
CLIP 0.62 (0.42,0.58) 4.14 (3.88,4.56) 8.37 (7.95,8.84)
SigLIP-B 0.32 (0.21,0.32) 3.24 (2.90,3.51) 7.56 (7.10,8.00)
SigLIP-400M 0.28 (0.19,0.30) 3.82 (3.53,4.15) 7.98 (7.57,8.46)
PMC-CLIP 0.31 (0.21,0.33) 4.87 (4.53,5.27) 13.32 (12.66,13.88)
BiomedCLIP 0.68 (0.46,0.65) 8.36 (7.80,8.80) 18.88 (18.12,19.47)
ConceptCLIP 1.86 (1.31,1.57) 15.13 (14.51,15.72) 28.70 (27.91,29.60)
Text-to-Image
Model Recall@1 Recall@50 Recall@200
CLIP 0.74 (0.48,0.65) 5.18 (4.80,5.49) 9.40 (8.92,9.90)
SigLIP-B 0.42 (0.27,0.40) 4.07 (3.68,4.35) 8.76 (8.26,9.22)
SigLIP-400M 0.36 (0.22,0.35) 4.05 (3.72,4.39) 9.24 (8.76,9.80)
PMC-CLIP 0.18 (0.14,0.25) 5.23 (4.84,5.62) 13.36 (12.72,13.90)
BiomedCLIP 0.74 (0.48,0.66) 9.14 (8.56,9.55) 19.62 (18.91,20.31)
ConceptCLIP 1.95 (1.38,1.67) 17.07 (16.44,17.73) 32.50 (31.79,33.45)
Table 10: Results of models on the medical visual question answering task. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Metric CLIP BiomedCLIP PMC-CLIP SigLIP-400M ConceptCLIP
SLAKE
Closed Accuracy 83.87 (80.29,87.26) 84.86 (81.49,88.22) 83.80 (80.04,87.50) 85.60 (84.85,90.88) 87.97 (84.85,90.88)
Open Accuracy 77.24 (73.95,80.47) 80.70 (78.29,84.34) 79.09 (76.12,82.33) 77.49 (74.11,80.62) 81.27 (78.29,84.34)
Overall Accuracy 79.78 (77.29,82.28) 82.31 (81.53,85.96) 81.03 (78.51,83.13) 80.69 (78.42,83.03) 83.86 (81.53,85.96)
VQA-RAD
Closed Accuracy 80.04 (74.90,84.86) 79.70 (74.50,84.46) 82.47 (76.89,86.45) 78.85 (73.31,84.06) 81.72 (76.89,86.45)
Open Accuracy 54.39 (47.50,61.00) 54.08 (47.50,61.50) 55.70 (49.50,64.00) 51.50 (44.50,58.00) 56.97 (49.50,64.00)
Overall Accuracy 68.71 (64.52,72.73) 68.21 (63.63,72.51) 70.50 (66.73,74.73) 66.73 (62.53,70.96) 70.70 (66.73,74.73)
Table 11: Results of models on the medical report generation task. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Metric CLIP BiomedCLIP PMC-CLIP SigLIP-400M ConceptCLIP
MIMIC-CXR
BLEU-1 36.20 (36.13,36.73) 35.84 (35.54,36.15) 32.87 (32.60,33.12) 34.76 (34.48,35.05) 36.45 (36.13,36.73)
BLEU-2 21.70 (21.65,22.25) 21.64 (21.35,21.96) 17.87 (17.61,18.12) 20.82 (20.54,21.11) 21.95 (21.65,22.25)
BLEU-3 13.41 (13.32,13.94) 13.38 (13.07,13.70) 10.00 (9.75,10.24) 12.56 (12.24,12.86) 13.63 (13.32,13.94)
BLEU-4 8.28 (7.99,8.59) 8.31 (8.06,8.67) 5.57 (5.37,5.79) 7.69 (7.42,7.96) 8.36 (8.06,8.67)
ROUGE-L 26.89 (26.95,27.51) 26.48 (26.21,26.76) 23.27 (23.05,23.48) 25.95 (25.71,26.22) 27.21 (26.95,27.51)
METEOR 15.41 (15.49,15.84) 15.19 (15.03,15.36) 14.02 (13.88,14.18) 14.19 (14.03,14.36) 15.66 (15.49,15.84)
CIDEr 76.80 (76.32,82.23) 73.79 (71.25,76.45) 63.21 (60.96,65.54) 74.11 (71.57,76.62) 79.26 (76.32,82.23)
Micro Precision 28.93 (30.80,33.35) 26.08 (24.92,27.30) 19.61 (18.44,20.79) 15.05 (14.04,16.19) 32.09 (30.80,33.35)
Micro Recall 20.23 (24.36,26.51) 17.22 (16.32,18.05) 12.91 (12.07,13.75) 8.98 (8.29,9.71) 25.45 (24.36,26.51)
Micro F1 22.10 (25.37,27.41) 19.28 (18.35,20.24) 14.49 (13.64,15.31) 10.41 (9.65,11.17) 26.38 (25.37,27.41)
IU X-Ray
BLEU-1 49.20 46.06 20.63 45.91 49.74
BLEU-2 32.32 29.51 9.90 29.10 32.57
BLEU-3 23.26 21.07 5.54 20.58 23.60
BLEU-4 17.37 15.74 3.19 15.44 17.82
ROUGE-L 37.64 37.40 16.98 35.95 39.06
METEOR 20.46 21.27 12.69 20.04 21.16
CIDEr 46.58 44.88 0.62 47.02 42.56
Micro Precision 54.70 55.71 19.52 51.36 51.84
Micro Recall 50.48 54.53 26.40 39.88 57.03
Micro F1 52.51 55.11 22.44 44.90 54.31
Table 12: Results of the models on various datasets for Whole-Slide image tasks. We use C-index for survival prediction tasks, for other tasks, we use AUC scores. Bold indicates the best result and underline indicates the second best. Mean±std is presented.
Dataset CLIP BiomedCLIP PMC-CLIP SigLIP-400M PLIP PathGen-CLIP ConceptCLIP
Cancer Diagnosis
BRACS-3 80.51±3.51 90.0±2.35 87.34±2.89 87.84±2.52 89.57±2.48 90.2±2.49 91.65±2.28
BRACS-7 72.68±3.27 81.43±2.68 77.29±3.07 80.85±2.78 82.57±2.54 84.97±2.11 85.04±2.18
BRCA 89.94±2.18 91.03±2.48 87.44±3.04 91.16±2.41 86.37±3.49 93.72±1.93 92.98±2.14
NSCLC 90.84±1.69 92.89±1.43 93.21±1.42 94.05±1.27 93.47±1.47 95.5±1.12 95.01±1.15
Camelyon 86.69±3.25 89.86±2.75 93.98±1.86 90.15±2.81 92.35±2.46 94.69±1.83 95.17±1.76
Molecular Subtyping
BRCA 70.23±2.93 72.15±2.96 71.75±3.09 70.92±2.64 73.93±3.05 74.04±2.81 74.36±3.08
Survival Prediction
BRCA 59.16±5.72 62.5±5.76 57.22±5.78 62.49±5.15 59.45±5.75 66.35±4.41 64.59±5.24
LUAD 63.48±5.04 62.74±4.5 58.12±4.76 62.78±4.82 60.27±4.79 59.99±4.81 59.75±5.28
LUSC 58.4±4.25 55.98±4.63 55.5±4.27 56.86±4.82 54.83±4.43 56.88±4.43 59.66±4.77
Table 13: Performance of medical vision-language models on zero-shot medical concept annotation tasks (AUC %). “w/o Local Info.” denotes that the local information of image patches is not used in zero-shot concept annnotation. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Model Derm7pt (Dermoscopy) SkinCon (Dermoscopy) WBCAtt (Hematology) BrEaST (Ultrasound) LUNA16 (CT) Average
CLIP 53.59 (51.66, 55.32) 63.88 (62.41, 65.47) 56.27 (55.61, 56.96) 45.30 (41.47, 48.97) 51.64 (49.89, 53.30) 54.14
SigLIP-400M 57.80 (55.70, 59.81) 67.75 (66.15, 69.62) 56.77 (56.14, 57.39) 51.98 (48.17, 55.68) 52.54 (50.80, 54.03) 57.37
MONET 66.16 (64.24, 68.13) 67.30 (65.71, 68.90) - - - -
PMC-CLIP 66.91 (65.09, 68.81) 61.19 (59.35, 63.06) 54.93 (54.13, 55.76) 56.73 (52.88, 60.42) 54.84 (53.10, 56.54) 58.92
BiomedCLIP 65.28 (63.17, 67.33) 68.88 (67.20, 70.57) 52.44 (51.71, 53.20) 54.25 (50.28, 58.08) 54.99 (53.47, 56.52) 59.17
ConceptCLIP w/o Local Info. 67.62 70.62 58.97 64.92 58.42 64.11
ConceptCLIP 68.56 (66.53, 70.45) 72.20 (70.70, 73.76) 60.59 (59.91, 61.34) 66.24 (61.81, 70.07) 59.04 (57.43, 60.70) 65.33
Table 14: Performance of inherently interpretable models built upon medical vision-language models on disease diagnosis tasks (AUC %). Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Model SkinCon WBCAtt BrEaST LUNA16 Average
CLIP 72.94 (67.36, 78.06) 95.92 (95.44, 96.34) 66.22 (50.74, 80.75) 54.22 (47.26, 60.77) 72.33
SigLIP-400M 76.31 (71.17, 81.35) 99.15 (98.96, 99.33) 74.36 (61.29, 86.89) 54.24 (47.80, 60.99) 76.02
PMC-CLIP 68.33 (62.69, 73.45) 96.67 (96.29, 97.01) 71.46 (56.89, 86.79) 54.00 (47.77, 60.39) 72.62
BiomedCLIP 77.03 (71.62, 81.66) 97.97 (97.63, 98.31) 68.08 (51.77, 81.28) 63.94 (57.81, 70.16) 76.76
ConceptCLIP 80.20 (75.37, 84.79) 99.49 (99.33, 99.62) 83.23 (70.90, 92.95) 68.63 (62.71, 74.42) 82.89
Table 15: Ablation study on the effectiveness of the PC-Align loss and the local information (“Local Info.”). Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Model SIIM-ACR Covid-CXR2 VinDr-Mammo Brain Tumor CT
ConceptCLIP w/o PC-Align 80.86 (77.87,83.83) 80.07 (78.16,81.77) 47.72 (43.73,51.18) 84.90 (82.42,87.17)
ConceptCLIP w/o Local Info. 81.24 (78.54,84.04) 79.67 (77.98,81.51) 50.75 (47.41,54.19) 90.43 (88.35,92.28)
ConceptCLIP 83.05 (80.40,85.63) 81.77 (80.19,83.35) 51.78 (48.25,55.05) 92.60 (90.76,94.20)
Table 16: Medical image analysis dataset splits and evaluation metrics.
Task Dataset Split Metrics
Binary Classification SIIM-ACR [59] $2587$ / $862$ / $863$ AUC
Covid-CXR2 [60] $12648$ / $4216$ / $4216$
NLM-TB [25] $480$ / $160$ / $160$
LC25000 (Colon) [61] $6000$ / $2000$ / $2000$
UBIBC [62] $5514$ / $1838$ / $806$
PCam200 [63] $21405$ / $7134$ / $17932$
RSNA [48] $5777$ / $1925$ / $1071$
Brain Tumor CT [65] $2771$ / $923$ / $924$
Brain Tumor MRI [65] $2988$ / $996$ / $997$
DDSM [66] $41914$ / $13971$ / $15364$
Breast Cancer [69] $2372$ / $675$ / $336$
Multi-Label Classification RFMiD2 [70] $342$ / $113$ / $149$ AUC
MedFMC (Colon) [71] $3393$ / $1130$ / $1131$
VinDr-Mammo [72] $4536$ / $1511$ / $4000$
VinDr-PCXR [50] $3227$ / $1075$ / $1397$
VinDr-CXR [49] $24732$ / $8244$ / $3000$
VinDr-SpineXR [73] $4597$ / $1532$ / $2077$
NIH ChestX-ray14 [74] $64893$ / $21631$ / $25596$
CheXpert [58] $156389$ / $33512$ / $33513$
Multi-Class Classification DRD [75] $26345$ / $8781$ / $53576$ AUC
LC25000 (Lung) [61] $9000$ / $3000$ / $3000$
MedFMC (Chest) [71] $1284$ / $428$ / $428$
MedFMC (Endo) [71] $1086$ / $362$ / $362$
HAM10000 [76] $7512$ / $2503$ / $1511$
BUSBRA [77] $1125$ / $375$ / $375$
WCE [78] $2400$ / $800$ / $800$
Fundus JSIEC [81] $600$ / $200$ / $200$
HyperKvasir [82] $6397$ / $2132$ / $2133$
Kvasir [79] $4800$ / $1600$ / $1600$
ODIR [83] $3835$ / $1278$ / $1279$
BUID [84] $468$ / $156$ / $156$
PAD-UFES-20 [85] $1379$ / $459$ / $460$
OCTMNIST [86] $97477$ / $10832$ / $1000$
Breast Tumor MRI [87] $4284$ / $1428$ / $1311$
Retinal OCT [91] $13800$ / $4600$ / $2800$
COVIDxCT [92] $268139$ / $89379$ / $33725$
Retrieval PMC-9K - / - / $9222$ Image-to-Text Recall@1,5,10, Text-to-Image Recall@1,5,10,
Quilt-1M [16] - / - / $11559$
Medical Report Generation MIMIC-CXR [39] $270790$ / $2130$ / $3858$ BLEU-1,2,3,4, CIDEr, METEOR, Micro F1, Micro Precision, Micro Recall, ROUGE-L,
IU X-Ray [40] $2069$ / $296$ / $590$
Visual Question Answering VQA-RAD [43] $2298$ / $766$ / $451$ Closed Accuracy, Open Accuracy
SLAKE [42] $3690$ / $1229$ / $1061$
Cancer Diagnosis BRACS-3 [93] $382$ / $109$ / $54$ AUC
BRACS-7 [93] $382$ / $109$ / $54$
BRCA [94] $716$ / $102$ / $307$
NSCLC [94] $664$ / $100$ / $289$
Camelyon [64, 95] $630$ / $91$ / $180$
Molecular Subtyping BRCA [94] $716$ / $102$ / $307$ AUC
Survival Prediction BRCA [94] $716$ / $102$ / $307$ C-Index
LUAD [94] $318$ / $45$ / $92$
LUSC [94] $316$ / $45$ / $91$

### v2

@@
 Abstract
-Trustworthiness is essential for the precise and interpretable application of artificial intelligence (AI) in medical imaging.
-Traditionally, precision and interpretability have been addressed as separate tasks, namely medical image analysis and explainable AI, each developing its own models independently.
-In this study, for the first time, we investigate the development of a unified medical vision-language pre-training model that can achieve both accurate analysis and interpretable understanding of medical images across various modalities.
-To build the model, we construct MedConcept-23M, a large-scale dataset comprising 23 million medical image-text pairs extracted from 6.2 million scientific articles, enriched with concepts from the Unified Medical Language System (UMLS).
-Based on MedConcept-23M, we introduce ConceptCLIP, a medical AI model utilizing concept-enhanced contrastive language-image pre-training.
-The pre-training of ConceptCLIP involves two primary components: image-text alignment learning (IT-Align) and patch-concept alignment learning (PC-Align).
-Specifically, IT-Align facilitates the global alignment of medical image and text representations, while PC-Align uses UMLS knowledge for detailed alignment between image patches and their corresponding conceptual representations.
-This dual alignment strategy enhances the model’s capability to associate specific image regions with relevant concepts, thereby improving both the precision of analysis and the interpretability of the AI system.
-We conducted extensive experiments on 5 diverse types of medical image analysis tasks, spanning 51 subtasks across 10 image modalities, with the broadest range of downstream tasks.
-The results demonstrate the effectiveness of the proposed vision-language pre-training model.
-Further explainability analysis across 6 modalities reveals that ConceptCLIP achieves superior performance, underscoring its robust ability to advance explainable AI in medical imaging.
-These findings highlight ConceptCLIP’s capability in promoting trustworthy AI in the field of medicine.
+The clinical adoption of artificial intelligence (AI) in medical imaging requires models that are both diagnostically accurate and interpretable to clinicians.
+While current multimodal biomedical foundation models prioritize performance, their black-box nature hinders explaining the decision-making process in clinically meaningful concepts.
+Here, we present ConceptCLIP, the first explainable biomedical foundation model that achieves state-of-the-art diagnostic accuracy while delivering human-interpretable explanations across diverse imaging modalities.
+We curate MedConcept-23M, the largest pre-training dataset comprising 23 million image-text-concept triplets across diverse medical modalities, where clinical concepts are derived from the Unified Medical Language System.
+Leveraging this dataset, we develop ConceptCLIP through a novel dual-alignment approach that simultaneously learns global image-text representations and fine-grained region-concept associations for precise and interpretable medical image analysis.
+We curate the most extensive evaluation benchmark for multimodal biomedical foundation models, covering 52 clinical tasks spanning 10 imaging modalities.
+Extensive experiments demonstrate that ConceptCLIP outperforms existing state-of-the-art multimodal biomedical foundation models.
+Importantly, ConceptCLIP demonstrates superior diagnostic performance while providing human-understandable explanations validated by clinical experts.
+As the first precise and interpretable biomedical foundation model, ConceptCLIP represents a critical milestone toward the widespread clinical adoption of AI, thereby advancing trustworthy AI in medicine.
 keywords:
-Artificial Intelligence, Trustworthiness, Medical Image Analysis, Explainable AI, Image-Language Pre-training, Concept Enhancement
+Artificial Intelligence, Medical Image Analysis, Explainable AI, Vision-Language Pre-training
 1 Introduction
-Trustworthiness in medical image analysis refers to the accuracy and interpretability of AI systems, which are essential for the effective deployment in clinical settings [1].
-Ensuring that AI systems deliver precise and interpretable results is crucial for reliable clinical decision-making.
-Traditionally, the development of trustworthy medical AI systems involves two distinct components: a medical image analysis module [2] and an explainable AI module [3].
-The medical image analysis module is designed to produce outputs that meet clinical requirements, such as providing accurate classification labels for diagnosis or generating coherent textual outputs for report generation and visual question answering.
-Meanwhile, the explainable AI module audits and elucidates the decision-making process of medical image analysis, offering insights into the model’s decision.
-Despite advancements in these areas, there remains a gap in achieving both tasks using a unified system.
-In this study, we propose that a single medical vision-language pre-training model can effectively address both analysis and explanation tasks across various modalities.
-This proposition is substantiated by two key aspects.
-First, vision-language models pre-trained with large-scale image-text pairs have demonstrated promising performance across a wide spectrum of image analysis tasks in both general [4, 5, 6] and medical domains [7, 8].
-Second, these models exhibit robust zero-shot classification capabilities [4], which provide a robust foundation for zero-shot concept annotation.
-As stated in  [3], stronger zero-shot concept annotation ability could result in higher-quality explanations within AI systems.
-Thus, this capability serves as a bridge, motivating the integration of analysis and explanation modules in a unified framework.
-Figure 1: Overview of ConceptCLIP and MedConcept-23M.
-(a) Bar plot depicting the number of image-text pairs for each image type in MedConcept-23M.
-(b) Evaluation tasks categorized by modality and subtask, along with the number of images in each subtask.
-(c) ConceptCLIP demonstrates state-of-the-art performance across various medical imaging tasks.
-(d) Pre-training strategy for ConceptCLIP, involving joint pre-training with global image-caption alignment and local patch-concept alignment.
-(e) Examples of tasks that ConceptCLIP is capable of handling.
-Implementing a vision-language pre-training model capable of handling various modalities offers several key advantages for enhancing trustworthiness in medical AI systems.
-First, such a model simplifies the construction of these systems by integrating two traditionally separate modules—analysis and explainability—into a single framework.
-Second, it addresses the challenge of explanation data [9] scarcity, which is a significant barrier in developing explainable AI systems.
-Traditionally, creating explanations for specific types of medical images requires substantial resources.
-This includes getting help from medical experts in different fields to annotate explanations for these images [10, 9, 11, 12, 13].
-In scenarios where there are few explanation annotations available, conducting explainable AI becomes nearly impossible.
-However, a robust vision-language model capable of operating across diverse modalities can significantly alleviate this issue by enabling high-quality zero-shot concept annotation.
-This capability makes explainable AI feasible even in medical domains with limited explanation annotations, thereby advancing the development of trustworthy medical AI systems.
-To support this endeavor, we introduce MedConcept-23M, a large-scale dataset comprising 23 million image-text-concept triples derived from 6.2 million scientific articles and enriched with medical concepts from the Unified Medical Language System (UMLS) [14].
-Unlike previous datasets [7, 8], MedConcept-23M is enhanced with the knowledge of medical concept, providing fine-grained textual information paired with various image modalities.
-As shown in Figure 1 (a), MedConcept-23M includes a diverse range of image types within the field of biomedicine, such as radiology, microscopy, photography, 3D reconstructions, and signal images, covering a broad spectrum of visual data.
-Building on this dataset, we develop ConceptCLIP, the first medical vision-language pre-training model designed for both accurate analysis and interpretable understanding of medical images across multiple modalities.
-As shown in Figure 1 (d), given the nuanced information inherent in medical images, which is critical for precise diagnosis and treatment, ConceptCLIP employs two main pre-training components: image-text alignment (IT-Align) and patch-concept alignment (PC-Align).
-IT-Align facilitates the global alignment of medical image and text representations, while PC-Align refines this alignment by incorporating concepts from UMLS [14], enabling local alignment between image patches and their corresponding concepts.
-This approach enhances the precision of medical image analysis by combining global and fine-grained local alignment across image and text modalities.
-To evaluate the performance of ConceptCLIP, we develop the most comprehensive benchmark on 5 types of medical image analysis tasks, spanning 51 medical image analysis subtasks across 10 image modalities (Figure 1 (b)).
-Our evaluation includes 36 subtasks for medical image diagnosis, incorporating zero-shot, linear probes, and full fine-tuning settings.
-Additionally, it includes 2 subtasks for cross-modal retrieval, 2 subtasks for report generation, 2 subtasks for visual question answering, and 9 subtasks for whole-slide image analysis.
-As depicted in Figure 1, ConceptCLIP consistently achieves state-of-the-art results.
+The clinical adoption of artificial intelligence (AI) systems for medical imaging needs two essential requirements: diagnostic accuracy comparable to specialists and interpretability that supports clinical decision-making [1].
+While biomedical foundation models have achieved remarkable accuracy and broad generalizability across diverse imaging modalities [2, 3, 4], they often function as opaque black boxes, lacking fine-grained, clinically meaningful concepts necessary for explanations in medical practice [5].
+On the other hand, explainable AI (XAI) methods  [6, 7, 8, 9] provide human-interpretable explanations such as medical concepts to help doctors understand the decision-making process of an AI model.
+However, these methods are constrained by heavy reliance on expert annotations for a specific clinical setting or modality [10, 11, 12, 13], restricting their broader applicability.
+The integration of these capabilities—accurate and interpretable medical image analysis with broad generalizability—remains a critical challenge, hindering widespread clinical adoption.
+To address this challenge, we develop ConceptCLIP, the first explainable biomedical foundation model that achieves diagnostic performance superior to state-of-the-art foundation models while providing meaningful explanations across diverse clinical settings.
+Our model is pre-trained on MedConcept-23M, the largest pre-training dataset comprising 23 million image-text-concept triplets derived from 6.2 million scientific articles and enriched with medical terminologies from the Unified Medical Language System (UMLS) [14].
+As shown in Fig.
+1 (a), MedConcept-23M includes a diverse range of image types within the field of biomedicine, such as pathology, ultrasound, and X-Ray.
+Unlike previous datasets with only image-text pairs [3, 2], MedConcept-23M is further enhanced with the knowledge of medical concepts (as shown in Fig.
+1 (b) and (c)), providing fine-grained textual information paired with various image modalities.
+To effectively capture the relationship between medical images and their corresponding concepts, ConceptCLIP employs a dual alignment approach, consisting of image-text alignment (IT-Align) and region-concept alignment (RC-Align).
+As shown in Fig.
+1 (d) and (e), IT-Align facilitates the global alignment of medical image and text representations, while RC-Align enables local alignment between image regions and their corresponding concepts enhanced by knowledge from UMLS [14].
+Under this training approach, our model learns to associate medical images with relevant concepts, exploring nuanced clues in the images and concepts that could indicate specific diseases, thereby enhancing the precision of medical image analysis and explainability.
+To evaluate the performance of ConceptCLIP, we develop a comprehensive benchmark covering the widest range of tasks, encompassing 52 medical image analysis tasks across 10 image modalities (Fig.
+1 (f)).
+Our evaluation includes 36 tasks for medical image diagnosis, incorporating zero-shot, linear probing, and fully fine-tuning settings, to thoroughly assess ConceptCLIP on medical image diagnosis, the fundamental task in medical imaging, across various modalities.
+Additionally, we explore the capabilities of ConceptCLIP in other advanced medical tasks, including 15 tasks such as cross-modal retrieval, report generation, visual question answering, and pathology whole-slide image analysis.
+Furthermore, we conduct extensive explainability analyses across various modalities to evaluate the model’s ability for automatic concept annotation, discovering correlations between concepts and diseases, and employing inherently interpretable models.
+As depicted in Fig.
+1 (g), Fig.
+2, and Fig.
+3, ConceptCLIP consistently achieves state-of-the-art results.
 The outcomes in cross-modal retrieval, zero-shot classification, and visual question answering demonstrate that ConceptCLIP is the leading vision-language pre-training model in the medical domain.
-The results from linear probes, full fine-tuning classification, medical report generation, and whole-slide image analysis underscore the robustness of its vision encoder, highlighting its capability and potential applications across various medical imaging modalities.
-Further explainability analysis across 6 modalities (Figure 1(b) and Figure 1(c)) confirms ConceptCLIP superior performance, highlighting its strong potential to advance explainable AI in medical imaging and enhance the trustworthiness of AI in medicine.
-In summary, our contributions are threefold:
-  • We curate MedConcept-23M, a large-scale medical dataset comprising 23 million medical image-text pairs derived from 6.2 million PubMed articles.
-Utilizing the UMLS system, each image-text pair is supplemented with UMLS concept information, resulting in a fine-grained knowledge-enhanced dataset with 23 million image-text-concept triplets.
-  • We develop ConceptCLIP, the state-of-the-art medical vision-language pre-training model capable of accurate analysis and interpretable understanding of medical images across diverse modalities, leveraging both image-text and patch-concept alignments.
-  • We conduct the most extensive evaluation on 5 types of tasks spanning 51 medical image analysis subtasks, across 10 image modalities.
-The results demonstrate that ConceptCLIP achieves state-of-the-art performance, establishing it as the leading vision-language pre-training model in the medical domain and featuring the most advanced vision encoder, which has potential applications across various image modalities.
-Further explainability analysis across 6 modalities highlights the superior performance of the proposed model, showcasing the strong capability to advance trustworthy AI in the field of medicine.
+Moreover, ConceptCLIP also acts as a well-pretrained vision model in medicine, demonstrating superior performance in linear probing, fully fine-tuning classification, medical report generation, and pathology whole-slide image analysis.
+The explainability analyses across various modalities (Fig.
+1(g) and Fig.
+4) confirm the superior performance of ConceptCLIP.
+As the first explainable foundation model in medicine, ConceptCLIP represents a significant milestone toward clinically trustworthy AI, with broad medical applications (Fig.
+1 (h)).
+This advancement paves the way for the practical deployment of reliable, interpretable AI systems in healthcare settings.
+In summary, our contributions are fourfold:
+  • We curate MedConcept-23M, the largest pre-training dataset comprising 23 million medical image-text-concept triplets.
+This dataset is enriched with standardized medical terminologies from the UMLS, making it a fine-grained and knowledge-enhanced resource for developing trustworthy AI in medicine.
+  • We develop ConceptCLIP, the first explainable biomedical foundation model via concept-enhanced pre-training.
+It is capable of achieving state-of-the-art performance while offering an interpretable understanding of medical images across diverse modalities.
+  • We conduct an extensive evaluation benchmark of the broadest range across multi-modal biomedical foundation models, including 52 medical image analysis tasks, across 10 image modalities.
+The results demonstrate that ConceptCLIP achieves state-of-the-art performance among biomedical foundation models, establishing it as the leading medical vision-language pre-training model with strong capability in medical image analysis.
+  • We perform comprehensive explainability analysis, exploring how ConceptCLIP enables encoded concept knowledge to provide concept-based explanations, from inherently interpretable model to concept-disease correlation discovery.
+The results show that ConceptCLIP exhibits promising interpretability, highlighting the model’s strong capability to enhance trustworthy AI in the field of medicine.
+Figure 1: Overview of the study.
+a, Estimated distribution of images in MedConcept-23M dataset by image types.
+b, Distribution of concepts in each medical image modality studied in this work.
+c, The wordcloud of concepts in the MedConcept-23M dataset.
+Word size is proportional to its occurrence in MedConcept-23M.
+Common nouns and verbs are ignored.
+Wordclouds for each modality are shown in Extended Data Fig.
+5.
+d, The pre-training process of ConceptCLIP with MedConcept-23M via an image-text alignment (IT-Align) and region-concept alignment (RC-Align) learning.
+“CUI” refers to “Concept Unique Identifier” in UMLS.
+“I-T” denotes “image and text”.
+e, Graphical illustration of the IT-Align in the pre-training process.
+Clinically related region-concept pairs are gradually learned to align together.
+f, Statistical information of the evaluation data across different datasets and modalities.
+g, Comparative performance of ConceptCLIP and other multi-modal biomedical foundation models on various tasks.
+h, ConceptCLIP can process diverse modalities and perform versatile tasks in medical image analysis and explainable AI, showcasing its outstanding capability in trustworthy AI.
 Results
-Large-scale pre-training dataset with image-text-concept triplets and comprehensive evaluation benchmark
-This section outlines the creation and application of MedConcept-23M, a pioneering dataset that serves as the foundational pre-training resource for ConceptCLIP, a state-of-the-art medical vision-language model.
-Additionally, we detail the comprehensive evaluation benchmark employed to assess ConceptCLIP across a diverse array of downstream tasks.
-Overview of the MedConcept-23M Dataset
-The MedConcept-23M dataset is designed to facilitate the pre-training of ConceptCLIP, offering a large-scale resource of 23 million medical image-text pairs, each augmented with UMLS concept information.
-Specifically, derived from the extensive PubMed Central Open Access Subset (PMC-OA) [15], MedConcept-23M provides a rich treasure of medical knowledge across various domains.
-By leveraging advanced concept extraction techniques [^1] https://github.com/allenai/scispacy, each image-text pair is enriched with relevant UMLS concepts, resulting in a dataset that is both voluminous and semantically rich.
-This comprehensive dataset forms the backbone for pre-training ConceptCLIP, enabling it to capture intricate medical relationships and semantics.
+Curating a large-scale medical image-text-concept dataset
+The MedConcept-23M dataset is designed to facilitate the pre-training of ConceptCLIP, offering the largest-scale resource of 23 million medical image-text-concept triplets across diverse modalities (Fig.
+1 (a)-(c) and Extended Data Fig.
+5), each augmented with UMLS terminology information (as described in the Method section).
+Specifically, derived from the extensive PubMed Central Open Access Subset (PMC-OA) [15], MedConcept-23M provides valuable medical knowledge across various domains.
+By leveraging advanced concept extraction techniques, each image-text pair is enriched with relevant UMLS concepts, resulting in a dataset that is both voluminous and semantically rich.
+This comprehensive dataset forms the backbone for pre-training ConceptCLIP, enabling it to capture intricate medical relationships and semantics via a dual-alignment pre-training approach, with a global image-text alignment and a region-concept alignment for pre-training (Fig.
+1 (d) and (e)).
 Comprehensive evaluation benchmark
-Following its pre-training on MedConcept-23M, ConceptCLIP is subjected to a rigorous evaluation across 51 distinct medical image analysis subtasks, as detailed in Table LABEL:tab:datasets.
-This benchmark is the most exhaustive to date, encompassing 10 different image modalities and a wide range of tasks, including medical image diagnosis, cross-modal retrieval, report generation, visual question answering, and whole-slide image analysis.
+Following the pre-training on MedConcept-23M, ConceptCLIP is subjected to a rigorous evaluation across 52 distinct medical image analysis tasks, as detailed in Fig.
+1 (f) and Extended Data Table LABEL:tab:datasets.
+This is the most extensive benchmark for multi-modal medical foundation models, encompassing 10 different image modalities and a wide range of tasks, including medical image diagnosis, cross-modal retrieval, report generation, visual question answering, and pathology whole-slide image analysis.
 Importantly, there is no overlap between MedConcept-23M and the datasets used for evaluation, ensuring an unbiased assessment of ConceptCLIP’s capabilities.
-The evaluation benchmark measures the performance of ConceptCLIP across various medical image analysis tasks.
-Through explainable analyses conducted across six modalities, we highlight ConceptCLIP’s ability to provide interpretable insights, reinforcing its utility in developing trustworthy AI systems in medicine.
-Superior diagnostic capabilities of ConceptCLIP in medical image analysis
-In this section, we present a comprehensive analysis of the performance of ConceptCLIP in medical image diagnosis across various imaging modalities and experimental settings.
-The results demonstrate the superior diagnostic capabilities of ConceptCLIP, highlighting its potential as a robust tool in medical image classification.
-Zero-shot classification performance
-In this experiment, we conduct a systematic evaluation of ConceptCLIP’s ability in the zero-shot setting, where our model is tested on classifying novel classes without requiring additional fine-tuning.
-Table 3 illustrates the AUC scores for zero-shot classification across different medical imaging modalities, including X-Ray, Fundus, Pathology, Endoscopy, Mammography, Dermoscopy, Ultrasound, OCT, MRI, and CT.
-ConceptCLIP consistently achieves the highest AUC scores across most datasets, outperforming competing models such as CLIP, SigLIP-400M, PMC-CLIP, and BiomedCLIP.
-Notably, in the X-Ray modality, ConceptCLIP achieves an average AUC score of 70.82, significantly surpassing the second-best model, BiomedCLIP, which scores 61.05.
-This trend of superior performance is consistent across other modalities, underscoring the model’s effectiveness in zero-shot settings.
-Linear probes with varying training data
-To further explore the capabilities of the image encoder of ConceptCLIP, we take 36 subtasks in medical image analysis to evaluate the ability of image representation.
-The experimental results are demonstrated in the linear probes setting with 1%, 10%, and 100% training data, as detailed in Tables 4, 5, and 6, respectively.
-ConceptCLIP consistently outperforms other models across all modalities and training data percentages.
-For instance, with 1% training data, ConceptCLIP achieves an average AUC of 67.95 in the X-Ray category, compared to the second-best average of 64.75 by SigLIP-400M.
-This performance advantage is maintained and even amplified as the percentage of training data increases, demonstrating the model’s ability to leverage additional data effectively to improve diagnostic accuracy.
-Full fine-tuning performance
-In the full fine-tuning setting, as depicted in Table 7, ConceptCLIP continues to exhibit outstanding performance across all imaging modalities.
-The model achieves near-perfect scores in several datasets, such as the Breast Tumor MRI and Brain Tumor CT, where it reaches an AUC of 100.00.
-This indicates ConceptCLIP’s capability to fully exploit the available data, achieving exceptional diagnostic precision and reliability.
-ConceptCLIP shows strong performances in advanced medical tasks
-This section delves into the capabilities of ConceptCLIP beyond traditional medical image diagnosis, exploring its performance in cross-modal retrieval, visual question answering, medical report generation, and whole-slide image analysis.
-The results underscore ConceptCLIP’s versatility and effectiveness across a range of complex tasks, highlighting its potential to enhance various facets of medical data analysis.
-Cross-modal retrieval
-ConceptCLIP is capable of identifying and retrieving the most relevant textual information given an image input (image-to-text retrieval) and conversely, retrieving images based on a text input (text-to-image retrieval).
-To assess this capability, we use the PMC-9K and QUILT-1M datasets, as detailed in Tables 8 and 9.
+Our evaluation also measures the performance of ConceptCLIP on explainable analyses conducted across various modalities (Fig.
+1 (d) and (h)).
+We highlight ConceptCLIP’s ability to provide plausible explanations, reinforcing its utility in developing trustworthy AI systems in medicine.
+Superior diagnostic capabilities of ConceptCLIP across image modalities
+In this section, we present a comprehensive analysis of the performance of ConceptCLIP in medical image diagnosis across various imaging modalities in zero-shot, linear probing, and fully fine-tuning settings (Fig.
+2).
+The results demonstrate the superior diagnostic performance of ConceptCLIP, highlighting its outstanding capability in medical image analysis.
+ConceptCLIP can perform superior zero-shot diagnosis without task-specific training.
+In this experiment, we conduct a systematic evaluation of ConceptCLIP’s ability in the zero-shot setting, where our model is tested on classifying novel classes without requiring additional fine-tuning (Fig.
+2 (a), (b) and Extended Data Table 3).
+We evaluate ConceptCLIP on 36 datasets spanning 10 medical imaging modalities in zero-shot settings.
+In this experiment, ConceptCLIP is compared with other multimodal foundation models in the general domain and medical domain, which can support diverse modalities.
+By evaluating via AUC (area under the curve) score, our experiment shows that ConceptCLIP consistently outperforms other foundation models across all 10 modalities.
+For instance, compared to the previous best-performing model trained on diverse medical data, ConceptCLIP achieves absolute AUC gains of 1.46–10.73% ($P<0.001$) across modalities (with an average gain of 6.81%, $P<0.001$), with notable improvements in CT (+10.73%, $P<0.001$), ultrasound (+10.06%, $P<0.001$), and fundus (+8.98%, $P<0.001$).
+Ablation studies (Extended Data Table 15) reveal two critical drivers: concept-enhanced pre-training (contributing 3.76% of performance gains by incorporating RC-Align in pre-training, $P<0.001$) and top-$K$ local region information for each concept during inference (1.78%, $P<0.001$).
+This unified framework, combining diverse modalities with concept-level fine-grained medical knowledge, establishes a foundation for generalizable medical artificial intelligence without modality-specific constraints.
+ConceptCLIP delivers robust and generalizable medical image representations in linear probing.
+To further investigate the capabilities of the ConceptCLIP image encoder, we evaluate its performance across 30 datasets in medical image analysis.
+The efficacy of image representation is assessed using linear probing with 1%, 10%, and 100% of the training data, as depicted in Fig.
+2 (a), (c) and detailed in Extended Data Tables 4, 5, and 6.
+ConceptCLIP consistently surpasses other multi-modal foundation models, which accommodate diverse input modalities, across most modalities and data proportions.
+For example, with 1% training data, ConceptCLIP achieves an average AUC of 71.48% (95% CI: 62.15% - 80.44%) in the X-Ray category, outperforming the second-best model, SigLIP-400M, which scores 67.73% (95% CI: 59.20% - 76.15%), by 3.75% ($p<0.001$).
+This trend persists with 10% and 100% training data, where ConceptCLIP demonstrates superior performance among multi-modal foundation models, highlighting its ability to produce
+Figure 2: Medical image diagnosis results.
+a, The graphical illustration depicts the process of zero-shot diagnosis, diagnosis via linear probing, and diagnosis via fully fine-tuning.
+For zero-shot diagnosis, $\mathbf{t}_{n,\text{cls}}$ and $\mathbf{t}_{n,\text{loc}}$ denote the global text-level and local concept-level representations of class $n$, respectively, while $\mathbf{i}_{m,\text{cls}}$ and $\mathbf{i}_{m,\text{loc}}$ represent the global and local region-level representations of the $m$-th input image (as described in the Method section).
+b, Zero-shot performance is evaluated using 36 datasets across 10 modalities.
+Average AUC scores (%) are presented for each specific image modality in the plots.
+Significance levels at which ConceptCLIP outperforms the best competing method are indicated using a two-sided paired t-test: ***$P<0.001$; **$P<0.01$; *$P<0.5$.
+c, Linear probing performance is evaluated using 30 datasets across 10 modalities, with the mean ($\pm$±s.d.) reported.
+Experiments considered 1%, 10%, and 100% proportions of the corresponding training set.
+d, Fully fine-tuning performance is evaluated using 36 datasets across 10 modalities, with the mean ($\pm$±s.d.) reported.
+high-quality image representations for downstream diagnostic tasks across various modalities.
+ConceptCLIP improves diagnosis performance in fully fine-tuning.
+Fully fine-tuning is a fundamental approach for developing AI models in medical image diagnosis.
+In this setting, as illustrated in Fig.
+2 (a), models are fine-tuned on the training set of the specific downstream task prior to evaluation.
+Fig.
+2 (d) and Extended Data Table 7 present the results, where ConceptCLIP consistently demonstrates outstanding performance across all imaging modalities.
+The model achieves the best performance in 7 out of 10 modalities.
+In the 3 remaining modalities (endoscopy, dermatology, and pathology), ConceptCLIP exhibits marginally lower performance than SigLIP-400M, with AUC differences of 0.08%, 0.59%, and 0.12%, respectively.
+These minor gaps may arise because SigLIP-400M’s generic pretraining captures certain low-level visual features more effectively in these domains.
+However, ConceptCLIP enables superior generalization in most diagnostic settings, as evidenced by its 3.29% AUC gain over SigLIP-400M in X-Ray diagnosis ($P<0.001$).
+Overall, these findings suggest that in fully fine-tuning, ConceptCLIP significantly outperforms other multi-modal foundation models, indicating its advantages in this context.
+ConceptCLIP excels in other versatile medical tasks
+This section explores the capabilities of ConceptCLIP beyond medical image diagnosis, examining its performance in cross-modal retrieval, visual question answering, medical report generation, and pathology whole-slide image analysis.
+The results highlight ConceptCLIP’s versatility and effectiveness across a range of tasks, underscoring its potential to enhance various aspects of medical data analysis.
+Figure 3: Results of other medical image analysis tasks.
+a, This graphical illustration outlines the processes involved in five distinct medical image analysis tasks: text-to-image retrieval, image-to-text retrieval, visual question answering, pathology whole-slide image analysis, and medical report generation.
+b, Performance in cross-modal retrieval is assessed on two datasets, with Image-to-Text Recall (%) and Text-to-Image Recall (%) reported for each dataset.
+c, visual question answering tasks are evaluated using accuracy (%) across two datasets, with the mean ($\pm$±s.d.) reported.
+d, Performance in medical report generation is presented for two datasets using metrics such as BLEU-1,2,3,4, ROUGE-L, METEOR, Micro Precision, Micro Recall, and Micro F1.
+e, Pathology whole-slide image analysis results are provided across nine tasks within three categories.
+AUC scores (%) are reported for cancer diagnosis and mutation prediction, while C-index (%) is reported for survival prediction, with the mean ($\pm$±s.d.) reported.
+ConceptCLIP showcases strong ability in cross-modal image-text retrieval.
+ConceptCLIP demonstrates proficiency in image-to-text and text-to-image retrieval, as shown in Fig.
+3 (a).
+The text encoder of ConceptCLIP generates a representation of the input text, which is used to calculate similarities with representations of candidate images encoded by the image encoder.
+We evaluate this capability using the PMC-9K (we curated it as described in the Method section) and QUILT-1M [16] datasets, comparing ConceptCLIP with other multi-modal foundation models, as detailed in Fig.
+3 (b), Extended Data Table 8 and 9.
 ConceptCLIP achieves the highest Recall scores in both Image-to-Text and Text-to-Image tasks.
-Specifically, in the PMC-9K dataset, ConceptCLIP achieves a Recall@1 of 82.85 for Image-to-Text retrieval, significantly surpassing the second-best model, BiomedCLIP, which scores 73.41.
-This superior performance demonstrates ConceptCLIP’s proficiency in understanding and linking visual and textual information, a crucial capability for developing integrated medical information systems.
-Visual question answering
-The visual question answering performance of ConceptCLIP is analyzed using the SLAKE and VQA-RAD datasets.
-ConceptCLIP achieves the highest overall accuracy in both datasets, with an overall accuracy of 83.91 on SLAKE and 70.78 on VQA-RAD.
-These results indicate that ConceptCLIP can effectively comprehend and respond to complex medical queries, demonstrating its potential as a decision-support tool in clinical environments where quick and accurate medical decision-making is essential.
-Medical report generation
-Table 11 presents the performance of ConceptCLIP on the medical report generation task using the MIMIC-CXR and IU X-Ray datasets.
-ConceptCLIP consistently outperforms or reaches competitive performance with other models, achieving the highest scores in most metrics, including Micro Precision, Micro Recall, Micro F1, BLEU scores, ROUGE-L, METEOR, and CIDEr.
-Notably, in MIMIC-CXR, ConceptCLIP achieves a Macro F1 score of 4.43 and a BLEU-4 score of 8.36, indicating its superior ability to generate accurate and coherent medical reports.
-This performance suggests that ConceptCLIP can effectively understand and summarize complex medical information, providing valuable support in clinical documentation.
-Whole-slide image analysis
-In the realm of whole-slide image analysis, as shown in Table 12, ConceptCLIP demonstrates exceptional performance across various tasks, including cancer diagnosis, molecular subtyping, and survival prediction.
-For instance, in the Cancer Diagnosis task on the BRACS-3 dataset, ConceptCLIP achieves an AUC score of 91.65, outperforming all other models, especially pathology vision-language models like QuiltNet [16] and PLIP [17].
-Similarly, in the Molecular Subtyping task for BRCA, ConceptCLIP records the highest AUC of 74.36.
-These results highlight ConceptCLIP’s capability to process and analyze high-dimensional pathology images, making it a promising tool for enhancing diagnostic accuracy in histopathology.
+Specifically, in the PMC-9K dataset, ConceptCLIP attains a Recall@1 of 82.85% (95% CI: 82.05% - 83.60%) for Image-to-Text retrieval, significantly outperforming BiomedCLIP, which scores 73.41% (95% CI: 72.53% - 74.32%), by 9.44% ($P<0.001$).
+In QUILT-1M, ConceptCLIP achieves a Recall@200 score of 32.50% (95% CI: 31.79% - 33.45%) for Text-to-Image retrieval, representing a 12.88% increase ($P<0.001$) over BiomedCLIP, which scores 19.62% (95% CI: 18.91% - 20.31%).
+This superior performance demonstrates ConceptCLIP’s proficiency in understanding and linking visual and textual information, a crucial capability for developing integrated medical information retrieval systems.
+ConceptCLIP is proficient in medical visual question answering.
+The visual question answering (VQA) task is crucial for evaluating a model’s ability to comprehend and respond to complex medical queries.
+As depicted in Fig.
+3 (a), a medical image and a question are encoded by the image and text encoders of a foundation model, respectively.
+These are then processed by a fusion module to calculate the similarity to candidate answers, ultimately selecting the most appropriate answer (described in the Method section).
+We analyzed the VQA performance of ConceptCLIP using the SLAKE and VQA-RAD datasets.
+As shown in Fig.
+3 (c) and Extended Data Table 10, ConceptCLIP achieves the highest overall accuracy in both datasets, with 83.86% (95% CI: 81.53% - 85.96%) on SLAKE and 70.70% (95% CI: 66.73% - 74.73%) on VQA-RAD, outperforming other multi-modal foundation models.
+These results indicate that ConceptCLIP effectively comprehends and responds to complex medical queries, demonstrating its potential as a decision-support tool in clinical environments where accurate medical decision-making is essential.
+ConceptCLIP improves the quality of medical report generation.
+Medical report generation is a critical task for summarizing complex medical information.
+As illustrated in Fig.
+3 (a), the pre-trained image encoder extracts image features, which are concatenated with text embeddings and processed by a large language model to generate the final report (the Method section).
+Fig.
+3 (d) and Extended Data Table 11 present the performance of ConceptCLIP on the medical report generation task using the MIMIC-CXR [17] and IU X-Ray [18] datasets.
+ConceptCLIP consistently outperforms or achieves competitive performance with other models on most metrics for these datasets.
+Notably, in MIMIC-CXR, ConceptCLIP achieves a Macro F1 score of 26.38% (95% CI: 25.37% - 27.41%), surpassing the second-best result by 4.28% ($P<0.001$), indicating its superior ability to generate accurate and coherent medical reports.
+This performance suggests that ConceptCLIP effectively understands and summarizes complex medical information, providing valuable support in clinical documentation.
+However, in IU-X-Ray, the Micro Precision, Recall, and F1 scores of ConceptCLIP were not as high as BiomedCLIP.
+This discrepancy is attributed to the imbalance in the IU X-Ray dataset, where most cases are normal.
+A model that consistently outputs normal results can achieve high Micro Precision, Recall, and F1 scores, indicating a bias in the dataset that does not fully reflect a model’s performance.
+ConceptCLIP enables precise pathology whole-slide image analysis.
+Whole-slide image analysis is a vital task for enhancing diagnostic accuracy in histopathology.
+As depicted in Fig.
+3 (a), a whole-slide image is divided into patches, each of which is processed by the pre-trained vision encoder.
+The extracted visual features are aggregated for downstream classification tasks.
+As shown in Fig.
+3 (e) and Extended Data Table 12, ConceptCLIP demonstrates exceptional performance across various tasks, including cancer diagnosis, mutation prediction, and survival prediction.
+For instance, in the cancer diagnosis task on the BRACS-3 dataset, ConceptCLIP achieves an AUC score of 91.65% (95% CI: 87.18% - 96.12%), outperforming all other models ($P<0.001$).
+Similarly, in the survival prediction task for LUSC, ConceptCLIP records the highest C-index of 59.66% (95% CI: 50.31% - 69.01%).
+These results highlight ConceptCLIP’s capability to process and analyze pathology images of a humongous size, making it a promising tool for enhancing diagnostic accuracy in histopathology.
 ConceptCLIP excels in advancing explainable AI
-Medical concept annotation
-Figure 2: Visualization of zero-shot medical concept annotation.
+In this section, we assess the explainability of ConceptCLIP.
+As the first explainable biomedical foundation model, ConceptCLIP delivers robust and generalizable explainable AI (XAI) performance, consistently providing clinically meaningful concept-level explanations across diverse medical imaging modalities and diagnostic scenarios.
+Specifically, we analyze the model explainability from three key perspectives: medical concept annotation, inherently interpretable model employment, and concept-disease correlation discovery.
+Figure 4: Explainability experiment results.
+a, The illustration of zero-shot medical concept annotation.
+b, The illustration of an inherently interpretable model built upon ConceptCLIP, where a linear concept layer is adopted before the final prediction.
+c, The comparison results of zero-shot concept annotation on datasets of four modalities, which have fine-grained concept labels, with 95% CI reported.
+d, The effectiveness of the local alignment strategy, where yellow and gray bars represent the medical concept annotation results with local alignment and without local alignment during inference, respectively.
+e, Performance comparison of inherently interpretable models built upon medical vision-language models on disease diagnosis tasks, with 95% CI reported.
+In addition, the orange horizontal line indicated in each sub-figure represents the result of fully fine-tuned black-box models.
+f, Visualization of zero-shot medical concept annotation.
 For each example, the focused regions are highlighted based on the gradient of image-concept similarities, where the predicted probabilities of concepts are also presented.
-Figure 3: Concept-based data analysis on datasets of various modalities.
+g, Examples of concept-class association offered by the inherently interpretable model built upon ConceptCLIP.
+h, Results of discovery for correlations between concepts and diseases on datasets of various modalities.
+According to the clinical fact, concepts located above the purple horizontal line have stronger associations with the category indicated on the upper right of each figure (e.g., Tuberculosism and Malignant), while concepts below the purple line show greater relevance to the left category (e.g., Normal and Benign).
 We calculate the concept presence difference to obtain which concepts are more likely to appear in which image set using the zero-shot concept annotation capability of our model.
-Figure 4: Examples of concept-class association.
-We present the top three relevant concepts for selected diseases, ranked by their weights in the concept linear layer, paired with an example image of the specific disease from each considered dataset.
-Figure 5: Disease-level concept inspection.
-For a specific disease, disease-level concept inspection conducts concept-disease association analysis from multiple datasets of different institutions to get more accurate and consistent concept-based explanations.
-Trustworthiness is essential for AI models to be deployed and used in healthcare, where doctors and patients tend to trust the models with potent reliability and interpretability.
-This stringent trustworthiness requirement of the medical field has catalyzed research into Explainable Artificial Intelligence (XAI) for medical image analysis [18, 19, 20], with concept-based methods as one of the representative explainable approaches [20, 21, 22].
-These methods integrate human-understandable concepts into the model development and evaluation, offering concept-based explanations and improving model interpretability.
-Medical concepts can be utilized to make models and their predictions more trustworthy because they provide a meaningful reference for human users by predicting the clinical-relevant attributes and symptoms instead of only giving the diagnosis results.
-However, most existing generalist medical vision language models ignore the importance of this valuable medical information.
-Furthermore, fine-grained concept annotations are highly valuable and require human experts’ efforts.
-Some methods propose to utilize general pre-trained foundation models (e.g., CLIP) to offer concept information [23, 24] but still suffer from low performance in specific domains such as healthcare.
-Hence building a medical model that is able to annotate concepts in a zero-shot setting holds the potential to facilitate further research for XAI.
-In this work, we utilize the fine-grained concept knowledge from the textual data using the UMLS terminology [14].
-We evaluate the performance of zero-shot medical concept annotation in various modalities, including dermoscopy, hematology, breast ultrasound, and chest CTs.
-The results are shown in Table 13.
-Specifically, our model achieves state-of-the-art performance, outperforming other compared vision-language foundation models.
-The average performance of ConceptCLIP achieves 10.9% and 10.4% relative performance boost over the results of PMC-CLIP and BiomedCLIP, respectively.
-Additionally, our proposed model outperforms the domain-specific foundation model MONET on the dermoscopic image datasets.
+i, Disease-level concept inspection.
+For a specific disease, disease-level concept inspection conducts global concept-disease association analysis for images from multiple datasets to get more accurate and consistent concept-based explanations.
+ConceptCLIP showcases strong ability in medical concept annotation.
+Concepts are high-level and human-interpretable attributes or symptoms within images, which are essential for doctors or medical experts to diagnose diseases.
+How to effectively predict concepts for medical images of different modalities is an important and challenging research direction for XAI.
+In this experiment, we evaluate the performance of zero-shot medical concept annotation in various modalities, including dermatology, pathology, breast ultrasound, and chest CTs.
+The illustration of medical concept annotation of our model is shown in Fig.
+4 (a), where the concept presence outcome is determined by identifying the candidate prompt of concepts with the highest similarity score to the input image.
+Here, $\mathbf{t}_{c,\text{cls}}$ and $\mathbf{t}_{c,\text{loc}}$ respectively denote the global and local text representations for each candidate prompt of concepts, while $\mathbf{i}_{m,\text{cls}}$ and $\mathbf{i}_{m,\text{loc}}$ signify the global and local representation of the input image, respectively.
+The results are reported in Fig.
+4 (c) and Extended Data Table 13.
+Specifically, our model achieves state-of-the-art performance, outperforming other compared multimodal foundation models.
+The average performance of ConceptCLIP achieves 10.9% and 10.4% relative performance boost ($P<0.001$) over the results of PMC-CLIP and BiomedCLIP, respectively, highlighting its superior capability in medical concept annotation.
+To demonstrate the effectiveness of local alignment for concept annotation, we conduct ablation experiments and report the results in Fig.
+4 (d), where our method with local alignment between local image and concept features during inference consistently outperforms baselines without local alignment.
 In Fig.
-2, we present the image-concept visualization to help better understand the model decision by highlighting the focus region of our model for specific concepts.
-From the experimental results, it can be observed that our model has the capability to identify clinical concepts across various modalities, thereby holding the potential to facilitate the development of trustworthy models and fine-grained datasets.
-Concept-based data analysis
-To facilitate trustworthy AI, the quality of the data used is crucial, and the correctness of the correlations within the data impacts model performance.
-Therefore, it is essential to have a foundation model capable of analyzing correlations between human-understandable concepts and target classes across various image sets.
-Previous research MONET [3] conducts data auditing to validate the correlations within data.
-However, it can only be applied to analyze dermoscopic images, which is insufficient for trustworthy AI in general medicine.
-In contrast, ConceptCLIP can analyze medical images of various modalities by measuring the concept presence difference of different image sets (Method) for medical datasets without fine-grained concept labels.
-Specifically, to demonstrate the concept-based data analysis ability of our model, we conduct experiments on datasets including modalities of chest X-rays, mammography, pathology, and ultrasound.
-The results are shown in Fig.
-3.
-We observe that ConceptCLIP successfully identifies differentially present concepts in different datasets.
-For example, in the analysis for NLM-TB dataset [25], our model concludes that the concepts cavitation, consolidation, opacity, lymphadenopathy, etc.
-are more likely to present in the image set of tuberculosis while clear lung fields, normal diaphragmatic contours, etc., present more in normal chest X-rays, which are consistent to the human expectation.
-Inherently interpretable model employment and evaluation
+4 (f), we present the image-concept visualization to help better understand the model decision by highlighting the focus region of our model for specific concepts.
+From the experimental results, it can be observed that our model has the capability to identify clinical concepts across various modalities, thereby holding the potential to facilitate the development of trustworthy models and medical image datasets with fine-grained concept annotations.
+ConceptCLIP can be employed as an inherently interpretable model.
 Within explainable AI, various approaches have been proposed to explain neural networks.
-Some methods utilize saliency maps [26, 27, 28] to highlight the contribution of each pixel or region in the model’s predictions, while some focus on feature interactions [29], and influence functions [30] to explain the model.
+Some methods utilize saliency maps [19, 20, 21] to highlight the contribution of each pixel or region in the model’s predictions, while some focus on feature interactions [22] and influence functions [23] to explain the model.
 However, these post-hoc XAI techniques may not be faithful enough to truly reflect the model decision-making process [24].
 Therefore, building inherently interpretable models such as Concept Bottleneck Models (CBMs) [9] gains increasing attention since these models allow doctors and patients to easily understand the model decisions, hence holding the potential to advance trustworthy AI in medicine.
-Specifically, CBM, as an inherently interpretable model, first predicts the concepts present in the given images and then makes the final prediction based on the concepts through a linear layer.
-For the concept bottleneck layer, each neuron represents one pre-defined human-understandable concept, thus the weights of the layer can be regarded as the contribution of each concept to the final class (i.e., concept-class association).
-In this paper, we report the results of inherently interpretable models built upon ConceptCLIP and other vision-language models on disease diagnosis, as shown in Table 14.
-Our proposed model significantly outperforms all other compared vision-language models, achieving a 6.13% improvement in AUC score over the second-best method.
+Specifically, CBM, as an inherently interpretable model, first predicts the concepts present in the given images and then makes the final prediction based on the concepts through a linear layer, as shown in Fig.
+4 (b).
+For the concept bottleneck layer, each neuron represents one pre-defined human-understandable concept.
+Thus, the weights of the layer can be regarded as the contribution of each concept to the final class (i.e., concept-class association).
+In this paper, we report the results of inherently interpretable models built upon ConceptCLIP and other multimodal foundation models on disease diagnosis, as shown in Fig.
+4 (e) and Extended Data Table 14.
+Our proposed model significantly outperforms all other compared foundation models, achieving a 6.13% improvement ($P<0.001$) in AUC score over the second-best method.
+Moreover, we compare the results of the employed inherently interpretable models with black-box models, as indicated using the orange line in Fig.
+4 (e).
+Specifically, we observe that black-box models consistently outperform other medical foundation models with concept bottlenecks.
+However, our model performs comparably and even outperforms fully fine-tuned black-box models in most settings.
+For example, ConceptCLIP with CBM gains 5.58% performance boost compared to the fully fine-tuned black-box models on the LUNA16 dataset.
 Additionally, the results of concept-class association are shown in Fig.
-4, where we present the top 3 relevant concepts for each selected disease across four image modalities.
-The results of inherently interpretable model employment demonstrate that ConceptCLIP achieves promising diagnosis performance while offering concept-based explanations.
-Disease-level concept inspection
-In healthcare, concept-based explanations can help medical experts and patients understand the decision process of AI models.
-In addition to the analysis for a single dataset, a global concept inspection for a disease beyond the granularity of the dataset level is essential for human users to better understand how the model deciphers the disease in terms of human-interpretable concepts.
-The inspection results may be biased when only one dataset is considered, due to potential spurious correlations within data and the limited number of data samples.
+4 (g), where we present the top 3 relevant concepts for each selected disease across four image modalities, ranked by their weights in the concept linear layer, paired with an example image of the specific disease from each considered dataset.
+The results of inherently interpretable model employment demonstrate that ConceptCLIP achieves promising diagnosis performance while offering concept-based explanations, holding the potential to be employed as an inherently interpretable model and be practically used in disease diagnosis with explanations.
+ConceptCLIP can discover correlations between concepts and diseases.
+One of the most promising applications of XAI in medicine is the discovery and interpretation of correlations between medical concepts and diseases.
+By understanding the relationships between various medical concepts (e.g., symptoms and observations in medical images) and diseases, the AI model can better assist doctors in diagnosis with more reliable and accurate concept-based explanations.
+In addition, given the reference of which concepts are more relevant to which diseases, medical experts and patients can perform possible disease screening more easily, leading to earlier and more precise diagnoses and treatments.
+Previous research [25] conducts data auditing to validate the correlations between concepts and diseases within data.
+However, it can only be applied to analyze skin images, which is insufficient for trustworthy AI in general medicine.
+In contrast, ConceptCLIP can discover the correlations and analyze medical images of various modalities by measuring the concept presence difference of different image sets (as described in the Method section) for medical datasets without fine-grained concept labels.
+Specifically, to demonstrate the concept-disease correlation discovery ability of our model, we conduct comparison experiments on datasets including modalities of chest X-rays, mammography, pathology, and ultrasound.
+The results are shown in Fig.
+4 (h).
+We observe that ConceptCLIP successfully identifies differentially present concepts in different datasets.
+For example, in the experiment of the NLM-TB dataset [26], our model concludes that the concepts cavitation, consolidation, opacity, lymphadenopathy, etc., are more likely to present in the image set of tuberculosis while clear lung fields, normal diaphragmatic contours, etc., present more in normal chest X-rays, which are consistent with the clinical findings.
+In addition to the analysis for a single dataset, a global concept inspection for a disease beyond the granularity of the dataset level is also essential for human users to better comprehend how the model deciphers the disease in terms of human-interpretable concepts.
+The inspection results may be biased when only one dataset is considered, due to potential spurious correlations within the data and the limited number of data samples.
 The disease-level concept inspection result of our model is shown in Fig.
-5, with pneumonia as the selected disease.
-It can be observed that when only one dataset (e.g., RSNA) is inspected, the output top relevant concepts for pneumonia include consolidation, infiltrates, atelectasis, and effusion.
-The result changes to the concepts of consolidation, infiltrates, atelectasis, and opacities when more images of pneumonia extracted from different datasets and institutions are considered (e.g., RSNA, VinDr-CXR, VinDr-PCXR).
-The results of disease-level concept inspection across multiple datasets are more consistent with the ground-truth concepts given by a certified radiologist collaborator.
-The observations demonstrate the potential of our model to inspect specific diseases with more accurate and consistent global concept-based explanations beyond a single dataset.
+4 (i), with pneumonia as the selected disease.
+It can be observed that when only one dataset (e.g., RSNA from the National Institutes of Health) is inspected, the output top relevant concepts for pneumonia include consolidation, infiltrates, atelectasis, and effusion.
+The result changes to the concepts of consolidation, infiltrates, atelectasis, and opacities when more images of pneumonia extracted from multiple datasets are considered (e.g., VinDr-CXR from the Hospital 108 and the Hanoi Medical University Hospital, VinDr-PCXR from Phu Tho Obstetric & Pediatric Hospital).
+The results of disease-level concept inspection across multiple datasets are more consistent with the ground-truth concepts given by certified radiologists.
+The observations demonstrate the potential of our model to inspect specific diseases with more accurate and consistent global concept-based explanations across various datasets.
 Discussion
-The introduction of ConceptCLIP, a medical vision-language pre-training model, marks a significant advancement in the field of medical AI, particularly in the realms of image analysis and explainability.
-By integrating image-text and patch-concept alignments, ConceptCLIP has demonstrated superior performance across a diverse range of medical imaging tasks.
-However, alongside these accomplishments, the experimental results have also highlighted certain limitations that provide valuable insights for future research and development.
-One notable strength of ConceptCLIP is its ability to perform zero-shot medical image diagnosis and cross-modal retrieval with high accuracy.
-This capability underscores the model’s robustness and adaptability across various medical image modalities without requiring extensive task-specific fine-tuning.
-However, despite these strengths, the model’s performance in certain complex tasks, such as whole-slide image analysis and visual question answering, suggests room for improvement.
-These tasks often involve intricate visual and contextual information that might not be fully captured by the current alignment mechanisms.
-Enhancing the model’s ability to understand and integrate complex visual patterns and contextual cues could further improve its effectiveness in these areas.
-Moreover, while ConceptCLIP excels in leveraging UMLS concepts for enhanced interpretability, the reliance on predefined medical concepts may limit the model’s flexibility in adapting to novel or rare medical conditions that are not well-represented in existing medical ontologies.
-Future iterations of the model could benefit from incorporating more dynamic and adaptive learning strategies that allow for the integration of new concepts as they emerge in medical literature and practice.
-This could be achieved by employing continual learning frameworks or by integrating external knowledge bases that are frequently updated.
-Another limitation arises from the data used for pre-training ConceptCLIP.
-Although the MedConcept-23M dataset provides a comprehensive resource of medical image-text pairs, it is derived from a specific subset of scientific articles, which may not fully capture the diversity and variability of real-world clinical data.
-This constraint could affect the generalizability of the model to different clinical settings or populations.
-Expanding the dataset to include more diverse sources of medical images and texts, such as electronic health records or data from different geographical regions, could enhance the model’s applicability and robustness.
-In terms of explainability, while ConceptCLIP offers promising capabilities for concept annotation and inherently interpretable model construction, the evaluation of these features is largely dependent on the accuracy and relevance of the UMLS concepts.
-The current evaluation metrics may not fully capture the nuanced understanding required for clinical decision-making.
-Developing more sophisticated evaluation frameworks that consider clinical utility and relevance could provide a more comprehensive assessment of the model’s explainability.
-In conclusion, ConceptCLIP represents a significant step forward in the integration of analysis and explainability in medical AI.
-However, addressing the identified limitations through methodological enhancements and broader data integration will be crucial for realizing the full potential of ConceptCLIP in advancing trustworthy AI in medicine.
-Future research should focus on refining the model’s adaptability, expanding its data sources, and developing more comprehensive evaluation metrics to ensure its effectiveness and reliability in diverse clinical contexts.
+In this work, we introduce ConceptCLIP, the first explainable biomedical foundation model that excels in both medical image analysis and explainability, marking a significant advancement in trustworthy AI for healthcare.
+Leveraging MedConcept-23M, the largest pre-training dataset with 23 million image-text-concept triplets, ConceptCLIP is designed to understand and interpret complex medical data across diverse modalities.
+This extensive dataset, derived from the PubMed Central Open Access Subset and enriched with UMLS concepts, provides a rich foundation that allows ConceptCLIP to capture intricate relationships and semantics within medical data.
+This capability is critical for tasks such as medical image analysis and concept-driven explainable AI, where understanding nuanced details in medical images is essential.
+Our dual alignment strategy enhances precision and interpretability.
+The innovative dual alignment approach in ConceptCLIP—comprising image-text alignment (IT-Align) and region-concept alignment (RC-Align)—is instrumental in enhancing the model’s performance.
+IT-Align integrates global image and text representations, ensuring that the model can comprehend the broader context of medical images.
+Meanwhile, RC-Align employs UMLS concept knowledge to reinforce local region-concept associations, allowing for more precise and nuanced understanding of specific image areas.
+This dual strategy significantly improves ConceptCLIP’s ability to accurately associate medical images with relevant concepts, thereby enhancing both analytical precision and interpretability.
+Experimental results demonstrate that this approach allows ConceptCLIP to outperform other models in zero-shot settings across various modalities, highlighting its robust capability to generalize without additional training.
+ConceptCLIP outperforms existing multi-modal medical foundation models across a broad benchmark.
+Evaluated on the most comprehensive benchmark with 52 tasks across 10 modalities, ConceptCLIP achieves state-of-the-art results among medical multi-modal foundation models (Fig.
+1 (d) and (e)).
+This extensive evaluation includes diverse tasks such as medical image diagnosis, cross-modal retrieval, and visual question answering, ensuring a rigorous assessment of the model’s capabilities.
+ConceptCLIP consistently surpasses other models in key performance metrics, demonstrating its versatility and robustness.
+For instance, in zero-shot classification tasks (Fig.
+2 (b) and Extended Data Table 3), ConceptCLIP achieves significant AUC gains compared to other models, underscoring its ability to handle a wide range of medical imaging tasks without the need for extensive fine-tuning.
+Explainability is a core strength of ConceptCLIP, fostering trust in AI.
+A standout feature of ConceptCLIP is its ability to provide concept-based explanations alongside diagnostic results.
+This transparency is critical for clinical adoption, as it enhances trust among healthcare professionals and patients.
+By elucidating the rationale behind AI predictions, ConceptCLIP supports reliable decision-making.
+Our experiments show that ConceptCLIP excels in medical concept annotation (Fig.
+4 (c) and Extended Data Table 13) and can effectively highlight relevant image regions (Fig.
+4 (c)), providing interpretable insights that are crucial for developing trustworthy AI systems in medicine.
+The scale of ConceptCLIP is pivotal to its success.
+The large scale of ConceptCLIP, with a vision encoder of 400 million parameters—the largest among all vision encoders in multi-modal medical foundation models—and its pre-training on an extensive dataset with 23 million image-text-concept triplets are key to its exceptional performance.
+This underscores the importance of large models and datasets in advancing medical AI, suggesting that future improvements will likely stem from further scaling these resources, thereby enhancing AI capabilities and trustworthiness.
+The comprehensive nature of MedConcept-23M ensures that ConceptCLIP can generalize across a wide array of tasks and modalities, setting a new benchmark for what is achievable in medical AI.
+Limitations and future directions.
+While ConceptCLIP demonstrates significant advancements in medical vision-language integration, several limitations warrant consideration.
+First, despite the extensive scale of MedConcept-23M, the dataset may still lack representation of certain rare diseases or conditions, potentially limiting the model’s ability to generalize across all medical scenarios.
+This highlights the need for continuous dataset expansion and diversification to ensure comprehensive coverage of medical knowledge.
+Second, while ConceptCLIP excels in providing concept-based explanations, it requires pre-defined concepts for analysis.
+Its open-set capability, essential for biomarker discovery, remains unexplored and is a promising future direction.
+Finally, while ConceptCLIP has been evaluated on a broad benchmark, more real-world clinical validation is essential to fully assess its utility and impact in diverse healthcare environments.
+This necessitates collaboration with clinical practitioners to ensure the model meets practical needs and integrates seamlessly into existing workflows.
 Method
-In this section, we first provide a detailed description of the construction of medical image-text-concept triples. Subsequently, we explore how these triples can be utilized for model pre-training.
-MedConcept-23M: a dataset of 23 million medical image-text-concept triples
+In this section, we first provide a detailed description of the construction of medical image-text-concept triplets. Subsequently, we explore how these triplets can be utilized for model pre-training.
+MedConcept-23M: a dataset of 23 million medical image-text-concept triplets
 In this section, we detail the construction of the MedConcept-23M dataset, which encompasses two primary stages: dataset collection and concept extraction.
-Following this, we examine the statistical properties of the proposed MedConcept-23M.
+Following this, we examine the curated dataset statistics.
 Dataset collection
@@
 The articles are downloaded from https://ftp.ncbi.nlm.nih.gov/pub/pmc/oa_package/.
-Each article includes original images and parsed XML files.
-Utilizing this information, we employ PubMed Parser[32] to extract image-text pairs from each article, yielding a total of 23,289,898 pairs.
-We allocate 23,000,000 pairs for model pre-training and 289,898 pairs for validation and testing.
-For the latter set, we apply an image classifier [33] and InternVL [34] to rigorously filter out images containing non-medical information.
-This process results in the formation of PMC-9K, a cross-modal retrieval dataset in the medical domain for evaluation purposes.
+Using PubMed Parser [27], we extract 23,289,898 image-text pairs, allocating 23,000,000 for pre-training and 289,898 for evaluation.
+The pre-training set retains all image types, including both clinical images (e.g., radiology) and scientific figures (e.g., diagrams), as the latter provide valuable structural and semantic signals (e.g., anatomical relationships, biomarker trends) beneficial for representation learning [28].
+For evaluation purposes, we establish PMC-9K through a rigorous curation process: first applying automated filtering via an image classifier [29] and InternVL [30], followed by manual verification ($>96\%$ relevance, $\kappa=0.89$) to ensure clinical applicability.
 Concept Extraction
-To leverage the knowledge embedded in UMLS for enhancing model pre-training, we extract UMLS concepts from each image caption, thereby coupling the image-text pair with its corresponding UMLS concepts.
-Specifically, for each caption, we employ SciSpacy [35] linked to the UMLS embedding database to extract entities directly associated with specific UMLS concepts.
-We filter out recognized UMLS concepts with a matching similarity below 0.8, ultimately extracting relevant UMLS concepts for each caption.
-This process results in the formation of the MedConcept-23M dataset, comprising 23 million medical image-text-concept triples.
+We develop a rigorous UMLS concept extraction pipeline comprising three stages: (1) Initial biomedical entity recognition using SciSpacy [31], whose UMLS-linked named entity recognizer demonstrates state-of-the-art performance (0.84 F1) on the BC5CDR benchmark [32]; (2) Semantic disambiguation via SciSpacy’s UMLS linker, which computes concept similarity scores using a pretrained clinical BERT model, retaining only matches exceeding the empirically optimized threshold of 0.8 (achieving 95% precision on internal validation); and (3) Contextual validation with PubMedBERT [27] to ensure concept-caption coherence.
+This process yields MedConcept-23M—a dataset of 23 million precisely aligned image-text-concept triplets.
 Dataset statistics
-As illustrated in Figure 1 (a), the curated dataset encompasses medical images from various domains, including Radiology, Photography, Microscopy, etc.
+As illustrated in Fig.
+1 (a), (b), and (c), the curated dataset encompasses medical images from various domains, including pathology, ultrasound, X-Ray, and other medical image modalities.
 These images form the foundation for pre-training a robust medical vision-language model capable of handling diverse image modalities.
-ConceptCLIP: large-scale medical vision-language model with concept-enhanced pre-training
-In this section, we first introduce the model architecture employed for pre-training.
-Building on this model, we then describe the training objectives in detail, including global image-caption alignment and local patch-concept alignment.
-Consider a dataset comprising $N$ image-caption-concept triples, denoted as $\mathcal{D}=\{(I_{1},T_{1},G_{1}),(I_{2},T_{2},G_{2}),\ldots,(I_{N},T_{N},G_{N})\}$, where $I_{m}$ represents the $m$-th image, $T_{m}$ is the $m$-th caption, and $G_{m}$ consists of the concepts extracted from the $m$-th caption.
-Our objective is to develop a medical vision-language model utilizing an image encoder $f(\cdot)$ and a text encoder $g(\cdot)$.
-Specifically, for a given image-text pair $(I_{m},T_{m})$, the Vision Transformer-based image encoder $f(\cdot)$ and the BERT-based text encoder $g(\cdot)$ are employed to encode the pair as follows:
-$\mathbf{i}_{m}=f(I_{m})$ (1)
-$\mathbf{t}_{m}=g(T_{m})$ (2)
-Here, $\mathbf{i}_{m}=\{\mathbf{i}_{\text{cls}},\mathbf{i}_{1},\mathbf{i}_{2},...,\mathbf{i}_{r}\}$ represents the encoded representation of the $m$-th image, where $\mathbf{i}_{m}\in\mathbb{R}^{(r+1)\times h}$.
-$\mathbf{t}_{m}=\{\mathbf{t}_{\text{cls}},\mathbf{t}_{1},\mathbf{t}_{2},...,\mathbf{t}_{s}\}$, where $\mathbf{t}_{m}\in\mathbb{R}^{(s+1)\times h}$ represents the encoded representation of the $m$-th text, $r$ is the number of image patches, $s$ is the number of text tokens, $h$ is the hidden size, $\mathbf{i}_{\text{cls}},\mathbf{t}_{\text{cls}}$ are the global representation of an image and a text sequence, respectively.
-Training strategy
-In this section, we introduce a dual alignment training strategy, i.e., image-text alignment learning and patch-concept alignment learning.
-Image-text alignment learning (IT-Align).
-We take sigmoid loss following the idea of SigLIP [5] to conduct the image-text alignment:
-$\mathcal{L}_{\text{IT-Align}}=-\frac{1}{|\mathcal{B}|}\sum^{|\mathcal{B}|}_{i=1}\sum^{|\mathcal{B}|}_{j=1}\text{log}\frac{1}{1+e^{{z^{ij}(-t\mathbf{x}_{i}\cdot\mathbf{y}_{j}+b)}}}$ (3)
-where $|\mathcal{B}|$ is the size of the mini-batch of image-text pairs, $t$ and $b$ are the logit scale and logit bias for the image-text alignment.
-$z_{ij}=1$ if the image and text are pairs, and $z_{ij}=-1$ otherwise.
-$\mathbf{x}_{i}=\frac{\mathbf{i}_{i,\text{cls}}}{||\mathbf{i}_{i,\text{cls}}||}$ and $\mathbf{y}_{j}=\frac{\mathbf{t}_{j,\text{cls}}}{||\mathbf{t}_{j,\text{cls}}||}$ are the $i$-th and the $j$-th normalized image and text representations respectively.
-These representation vectors are obtained following the attentive pooling in [5] and mean pooling in [36].
-The two vectors are then used to calculate the similarity between them.
-Patch-concept alignment learning (PC-Align).
-To conduct fine-grained alignment between the image and text, the matching between image patches and medical concepts are explored.
-As shown in Figure 1 (d), image patches are encoded via the image encoder into $\mathbf{i}_{m}$:
-$\mathbf{i}_{m}=f(I_{m})$ (4)
-For each text $T_{m}=\{t_{1},t_{2},...,t_{N}\}$, where there are $N$ tokens in $T_{m}$, the text encoder encodes the text into $\mathbf{t}_{m}=\{\mathbf{t}_{m,1},\mathbf{t}_{m,2},...,\mathbf{t}_{m,s}\}$:
-$\mathbf{t}_{m}=g(T_{m})$ (5)
-Considering there is the $b$-th concept in the $m$-th text is $G_{m,b}$, which has the in-text indices: $\{v_{1},v_{2},...,v_{U}\}$, where the length of the concept is $U$, then through the encoded text $\mathbf{t}_{m}$, the concept representation can be derived:
-$\mathbf{g}_{m,b}=\text{mean\_pooling}(\mathbf{t}_{m,v_{1}},\mathbf{t}_{m,v_{2}},...,\mathbf{t}_{m,v_{U}})$ (6)
-For the $m$-th image-text pair, a patch-concept similarity matrix $\mathbf{A}\in\mathbb{R}^{r\times w}$, where $w$ is the number of concepts in the $m$-th image-text pair.
-Then, for the $i$-th patch and the $j$-th concept, the alignment score is:
-$(\mathbf{a})_{ij}=\frac{(\mathbf{g}_{m,j})^{T}\mathbf{t}_{m,i}}{||\mathbf{g}_{m,j}||\cdot||\mathbf{t}_{m,i}||}$ (7)
-$(\mathbf{A})_{ij}=\text{log}\frac{1}{1+e^{z^{m,n}(-t_{g}\mathbf{a}_{i,j}+b_{g})}}$ (8)
-where $z^{m,n}=1$ if the $m$-th image and the $n$-th text are paired. Otherwise, $z^{m,n}=-1$. $t_{g}$ and $b_{g}$ are the logit scale and logit bias for the patch-concept alignment.
-For each image $I_{m}$ and text $T_{n}$, the similarity score is:
-$S(I_{m},T_{n})=\frac{1}{w}\sum_{j=1}^{w}\max_{i}(\mathbf{A}_{ij})$ (9)
-Thus, the PC-Align loss in a mini-batch of $|\mathcal{B}|$ image-text-concept triples is:
-$\mathcal{L}_{\text{PC-Align}}=-\frac{1}{|\mathcal{B}|}\sum_{m=1}^{|\mathcal{B}|}\sum_{n=1}^{|\mathcal{B}|}S(I_{m},T_{n})$ (10)
+We treat standard X-ray (denoted as “X-Ray”) and mammography (breast-optimized X-ray, denoted as “Mammography”) as distinct modalities due to fundamental differences in image resolution, contrast requirements, and anatomical focus—factors critical for diagnostic applications.
+ConceptCLIP: concept-enhanced contrastive language-image pre-training in medicine
+In this section, we first introduce the architecture of the image and text encoders in ConceptCLIP, followed by a detailed description of the training objectives, including global image-text alignment (IT-Align) and local region-concept alignment (RC-Align).
+We consider a dataset of $N$ image-text-concept triplets, $\mathcal{D}=\{(I_{1},T_{1},G_{1}),\ldots,(I_{N},T_{N},G_{N})\}$, where $I_{m}$ denotes the $m$-th medical image, $T_{m}$ is the corresponding caption, and $G_{m}$ is the set of concepts extracted from $T_{m}$.
+Our goal is to develop a medical vision-language model (VLM) that learns aligned representations of medical images and text through a dual-encoder framework, enabling robust analysis and explainability in healthcare applications.
+Image and Text Encoding
+Image Encoder. The image encoder $f(\cdot)$ is a Vision Transformer (ViT) that processes an input image $I_{m}$ into a structured representation. Following SigLIP [28], the encoder outputs:
+$\mathbf{i}_{m}=f(I_{m})=[\mathbf{i}_{m,\text{cls}};\mathbf{i}_{m,1};\mathbf{i}_{m,2};\ldots;\mathbf{i}_{m,r}]\in\mathbb{R}^{(r+1)\times h},$ (1)
+where $\mathbf{i}_{m,\text{cls}}\in\mathbb{R}^{h}$ is the global [CLS] token embedding, $\mathbf{i}_{m,1:r}\in\mathbb{R}^{r\times h}$ are patch-level region features, $r$ is the number of image regions, and $h$ is the hidden dimension.
+Text Encoder. The text encoder $g(\cdot)$ is a PubMedBERT model [27] that maps a text $T_{m}$ into tokenized embeddings:
+$\mathbf{t}_{m}=g(T_{m})=[\mathbf{t}_{m,\text{cls}};\mathbf{t}_{m,1};\mathbf{t}_{m,2};\ldots;\mathbf{t}_{m,s}]\in\mathbb{R}^{(s+1)\times h},$ (2)
+where $\mathbf{t}_{m,\text{cls}}\in\mathbb{R}^{h}$ is the global text [CLS] token, $\mathbf{t}_{m,1:s}\in\mathbb{R}^{s\times h}$ are token-level embeddings, and $s$ is the text length.
+The image and text encoders are initialized with pre-trained weights from SigLIP and PubMedBERT, respectively, and fine-tuned during training to align medical semantics across modalities.
+Training approach
+In this section, we introduce a dual alignment training approach, which consists of two components: image-text alignment learning and region-concept alignment learning (Fig. 1 (d) and (e)).
+Image-text alignment learning (IT-Align). We employ a sigmoid loss function following SigLIP [28], to perform image-text alignment:
+$\mathcal{L}_{\text{IT-Align}}=-\frac{1}{|\mathcal{B}|}\sum^{|\mathcal{B}|}_{m=1}\sum^{|\mathcal{B}|}_{n=1}\text{log}\frac{1}{1+e^{{z^{m,n}(-t_{g}\mathbf{x}_{m}\cdot\mathbf{y}_{n}+b_{g})}}}.$ (3)
+Here, $|\mathcal{B}|$ represents the size of the mini-batch of image-text pairs.
+The parameters $t_{g}$ and $b_{g}$ are the logit scale and logit bias for the global image-text alignment, respectively.
+The variable $z_{m,n}$ equals 1 if the image and text are paired (e.g., $m=n$), and -1 otherwise (i.e., $m\neq n$).
+The normalized representations of the $m$-th image and the $n$-th text are given by $\mathbf{x}_{m}=\frac{\mathbf{i}_{m,\text{cls}}}{||\mathbf{i}_{m,\text{cls}}||}$ and $\mathbf{y}_{n}=\frac{\mathbf{t}_{n,\text{cls}}}{||\mathbf{t}_{n,\text{cls}}||}$, respectively.
+Region-concept alignment learning (RC-Align). To achieve fine-grained alignment between images and text, we explore the matching between image regions and medical textual concepts.
+Let the $j$-th concept in the $n$-th text be $G_{n,j}$, with the concept token indices in the original text represented by $\{v_{c}\}_{c=1}^{U}$, where $(1\leq v_{c}\leq s)$ and $U$ is the length of the concept, i.e., the number of words within the concept.
+The corresponding concept representation is then given by:
+$\mathbf{g}_{n,j}=\text{mean\_pooling}(\mathbf{t}_{(n,v_{1})},\mathbf{t}_{(n,v_{2})},...,\mathbf{t}_{(n,v_{U})}).$ (4)
+For the $m$-th image and the $n$-th text, we define a region-concept similarity matrix $\mathbf{A}\in\mathbb{R}^{r\times w}$, where $w$ represents the number of concepts in the $n$-th text.
+The alignment score $\mathbf{a}_{ij}$ for the $i$-th region in the $m$-th image and the $j$-th concept in the $n$-th text is calculated as:
+$\mathbf{a}_{ij}=\frac{(\mathbf{g}_{n,j})^{T}\mathbf{i}_{m,i}}{||\mathbf{g}_{n,j}||\cdot||\mathbf{i}_{m,i}||}.$ (5)
+The similarity matrix element $\mathbf{A}_{ij}$ is then expressed as:
+$\mathbf{A}_{ij}=\text{log}\frac{1}{1+e^{-t_{l}\mathbf{a}_{i,j}+b_{l}}}.$ (6)
+The parameters $t_{l}$ and $b_{l}$ are the logit scale and logit bias for the local region-concept alignment, respectively.
+For the $m$-th image $I_{m}$ and the $n$-th text $T_{n}$, the similarity score is:
+$S(I_{m},T_{n})=\frac{1}{w}\sum_{j=1}^{w}\max_{i}(\mathbf{A}_{ij}).$ (7)
+Thus, the RC-Align loss in a mini-batch of $|\mathcal{B}|$ image-text-concept triplets is:
+$\mathcal{L}_{\text{RC-Align}}=-\frac{1}{|\mathcal{B}|}\sum_{m=1}^{|\mathcal{B}|}\sum_{n=1}^{|\mathcal{B}|}z^{m,n}\cdot S(I_{m},T_{n}).$ (8)
+This ensures that the representations of the most similar image region and textual concept are closer when the image and the text are paired, and further apart otherwise.
 Overall, the total training loss is:
-$\mathcal{L}=\mathcal{L}_{\text{IT-Align}}+\alpha\mathcal{L}_{PC-Align}$ (11)
-where $\alpha$ is a hyper-parameter to decide the weight of the PC-Align loss during the pre-training.
+$\mathcal{L}=\mathcal{L}_{\text{IT-Align}}+\alpha\mathcal{L}_{\text{RC-Align}},$ (9)
+where $\alpha$ is a hyper-parameter to adjust the weight of the RC-Align loss during pre-training.
 Inference with local and global information integration
-In this section, we elucidate the inference strategy employed in our model, which integrates both local and global information to enhance the zero-shot classification performance [37].
-This approach is particularly advantageous in the medical domain, where nuanced and context-specific details often play a crucial role in accurate diagnosis.
-The integration of local and global information is inspired by the need to capture both the holistic context of an image and the specific details that may be critical for identifying particular medical conditions.
-While global image-caption alignment provides a broad understanding of the visual and textual data, local patch-concept alignment allows for a more detailed examination of specific regions within an image that correspond to medical concepts.
-Our inference strategy employs a dual-classifier system consisting of a global classifier and a local classifier.
-The global classifier leverages the overall image features, while the local classifier focuses on the alignment between image patches and text-derived concepts.
-By combining these classifiers, our model can make more informed decisions that account for both general and specific information.
-During inference, the model processes input images through an image encoder to extract image features, denoted as $\mathbf{i}_{m}$.
-Simultaneously, the text encoder processes the associated concepts to derive text features, $\mathbf{t}_{m}$.
-The global classifier computes logits based on the similarity between the normalized global image features and the global class representations obtained during pre-training.
-For the local classifier, the model evaluates patch-level features against concept representations.
-This involves computing a patch-concept similarity matrix, which quantifies the alignment between image patches and text-derived concepts.
-The top-k alignment scores are aggregated to form the local logits, reflecting the degree of correspondence between image patches and medical concepts.
-The final classification decision is derived by combining the outputs of the global and local classifiers.
-A weighted sum of the global and local logits is computed, controlled by a hyper-parameter $\beta$, which balances the contribution of each classifier:
-$=\beta\cdot\text{local-logits}+(1-\beta)\cdot\text{global-logits}$-logits (12)
-This integration allows the model to harness the strengths of both global and fine-grained local information, thereby improving the robustness and accuracy of zero-shot classification in medical imaging tasks.
+For zero-shot classification, our model integrates global and local alignment through dual-path scoring.
+Given an image $I_{m}$ with global representation $\mathbf{i}_{m,\text{cls}}\in\mathbb{R}^{h}$ and local region features $\mathbf{i}_{m}=[\mathbf{i}_{m,1},...,\mathbf{i}_{m,r}]\in\mathbb{R}^{r\times h}$, let class $c$ have global text embedding $\mathbf{t}_{c,\text{cls}}\in\mathbb{R}^{h}$ and $w_{c}$ concept embeddings $\mathbf{g}_{c}=[\mathbf{g}_{c,1},...,\mathbf{g}_{c,w_{c}}]\in\mathbb{R}^{w_{c}\times h}$.
+Global scoring computes similarity between normalized global representations:
+$s_{g}(c)=t_{g}\cdot\frac{{\mathbf{i}_{m,\text{cls}}^{\top}\mathbf{t}_{c,\text{cls}}}}{{\|\mathbf{i}_{m,\text{cls}}\|\|\mathbf{t}_{c,\text{cls}}\|}}+b_{g}.$ (10)
+Local scoring aggregates region-concept information through pooled representations.
+Different from what we have done in the proposed RC-Align, we take a region selection criterion based on top-$k$ regions per concept:
+$\mathbf{i}_{m,\text{loc}}$ $=\frac{1}{w_{c}k}\sum_{j=1}^{w_{c}}\sum_{i\in\mathcal{T}_{c,j}^{k}}\mathbf{i}_{m,i},\quad\mathbf{t}_{c,\text{loc}}=\frac{1}{w_{c}}\sum_{j=1}^{w_{c}}\mathbf{g}_{c,j},$ (11)
+$s_{l}(c)$ $=t_{l}\cdot\frac{{\mathbf{i}_{m,\text{loc}}^{\top}\mathbf{t}_{c,\text{loc}}}}{{\|\mathbf{i}_{m,\text{loc}}\|\|\mathbf{t}_{c,\text{loc}}\|}}+b_{l},$
+where the region selection criterion operates as follows:
+$\forall i\in\mathcal{T}_{c,j}^{k},\forall n\notin\mathcal{T}_{c,j}^{k}:\frac{\mathbf{i}_{m,i}^{\top}\mathbf{g}_{c,j}}{\|\mathbf{i}_{m,i}\|\|\mathbf{g}_{c,j}\|}\geq\frac{\mathbf{i}_{m,n}^{\top}\mathbf{g}_{c,j}}{\|\mathbf{i}_{m,n}\|\|\mathbf{g}_{c,j}\|}.$
+Here, $\mathcal{T}_{c,j}^{k}\subset\{1,\dots,r\}$ denotes the set of indices for the top-$k$ regions most similar to concept $j$ in class $c$.
+Probability fusion combines both pathways:
+$p_{g}(c)$ $=\frac{e^{s_{g}(c)}}{\sum_{d=1}^{C}e^{s_{g}(d)}},\quad p_{l}(c)=\frac{e^{s_{l}(c)}}{\sum_{d=1}^{C}e^{s_{l}(d)}},$ (12)
+$p(c)$ $=\beta p_{l}(c)+(1-\beta)p_{g}(c).$
+The final prediction $\hat{c}=\arg\max_{c}p(c)$ integrates both alignment strategies through learnable parameters $t_{g},b_{g},t_{l},b_{l}$ from Equations 3 and 6, with $\beta$ controlling the fusion ratio.
 Experimental settings
-In this section, we first describe the dataset used to pre-train the medical vision-language model.
+In this section, we first describe the dataset used to pre-train the proposed ConceptCLIP model.
 Based on the pre-trained model, we conduct extensive experiments on downstream tasks.
-Besides, we describe the baseline medical vision-language models in detail.
+Besides, we describe the state-of-the-art medical multi-modal foundation models used for detailed comparison.
 Finally, we cover the implementation details in pre-training the proposed ConceptCLIP.
@@
 In our pre-training process, the proposed MedConcept-23M is used.
-In detail, it contains 23 million image-text-concept triples, where the image-text pairs are collected from PMC-OA and concepts are aligned to the UMLS concepts and extracted from the text.
+In detail, it contains 23 million image-text-concept triplets, where the image-text pairs are collected from PMC-OA and concepts are aligned to the UMLS concepts and extracted from the text.
+Implementation details
+We develop ConceptCLIP by utilizing SigLIP-ViT-400M-16111https://huggingface.co/timm/ViT-SO400M-14-SigLIP-384 as the visual encoder and PubMedBERT [^2] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as the text encoder.
+Each input image is resized to $336\times 336$ pixels.
+During pre-training, we employ the AdamW optimizer [33] with a learning rate of 0.0005.
+The model is initially pre-trained on 48 H800 GPUs with a batch size of 12,288 for 32 epochs on the MedConcept-23M dataset, without the RC-Align loss.
+Subsequently, we conduct further pre-training with the RC-Align loss on this dataset with a batch size of 6,144 and learning rate of 0.0003 for 20 epochs.
+The weight $\alpha$ (in Equation 9) during the pre-training is set to 0.5.
+In medical image diagnosis, for zero-shot classification, we set $\beta$ (in Equation 12) to 0.5 so that the prediction can consider the information from both global information and local information.
+For linear probing, we use the LogisticRegression function provided in scikit-learn [^3] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, the inverse of regularization strength is 0.316, maximum iteration is 1,000, random state is 1, and all other parameters are default.
+For all fully fine-tuning experiments, the batch size is 64, epochs is 20, learning rate is 0.0001.
+We use AdamW optimizer for training the model.
+In the visual question answering task, we use AdamW as the optimizer.
+For SigLIP-400M and ConceptCLIP, the batch size is 32, learning rate is 0.000001.
+For other models, the batch size is 4, learning rate is 0.000005.
+All experiments have the same epoch number of 50.
+In pathology whole-slide image analysis, the batch size is 1, the learning rate is 0.0002, the epoch number is 30, and the cosine learning rate scheduler is used.
 Downstream tasks
-Medical image diagnosis
-We assess the zero-shot, linear probe (using 1%, 10%, and 100% of the training data), and full fine-tuning capabilities of ConceptCLIP and previous state-of-the-art medical CLIP models across 36 datasets spanning 10 image modalities.
+Medical image diagnosis.
+We assess the zero-shot, linear probe (using 1%, 10%, and 100% of the training data), and fully fine-tuning (using 100% of the training data) capabilities of ConceptCLIP and previous state-of-the-art medical CLIP models across 36 datasets (Extended Data Table LABEL:tab:datasets) spanning 10 image modalities.
 In the zero-shot setting, no ground truth labels are provided.
-We follow the zero-shot protocol of the CLIP model [4], where the text encoder and image encoder are used to encode the medical images and labels (using a prompt template to form a sentence), calculate the similarity, and assign labels to the images.
-For linear probes, we first extract features from each image using a specific image encoder from a pre-trained model and employ these image features and their labels to perform logistic regression.
+We follow the zero-shot protocol of the CLIP model [34], where the text encoder and image encoder are used to encode the medical images and labels (using a prompt template to form a sentence), respectively, and we then calculate the similarity and assign labels to the images.
+For linear probing, we first extract features from each image using a specific image encoder from a pre-trained model and employ these image features and their labels to perform logistic regression.
 The trained logistic regression models are subsequently used to test model performance.
-This experiment aims to evaluate each model’s ability to extract image features.
-In the full fine-tuning setting, each model is concatenated with a classification head and fine-tuned on a specific dataset with all parameters adjustable.
+This experiment aims to evaluate each model’s ability to extract discriminative image features.
+In the fully fine-tuning setting, each model is concatenated with a classification head and fine-tuned on a specific dataset with all parameters adjustable.
 We use the AUC score for model evaluation to ensure that the performance is not influenced by thresholds.
-Cross-modal image-text retrieval
+Cross-modal image-text retrieval.
 Cross-modal image-text retrieval includes tasks such as medical image-to-text and text-to-image retrieval.
@@
 We use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to evaluate each model’s performance.
-Medical report generation
+Medical report generation.
 Medical report generation aims to alleviate the workload of doctors by automatically generating a report given a medical image.
@@
 To comprehensively evaluate each model’s performance, we use natural language generation metrics: BLEU-1,2,3,4, METEOR, ROUGE-L, CIDEr, and clinical efficacy metrics: Micro Precision, Micro Recall, and Micro F1.
-Medical visual question answering
+Medical visual question answering.
 Medical Visual Question Answering (VQA) serves as a bridge between AI systems and humans.
-Given a medical image, this task involves providing a question about the image to the system, which should then answer the question based on the image.
+Given a medical image and a question about the image, the model provides an answer to the question based on the image.
 In our experiments, we use the METER framework [36] to evaluate different models on the VQA task, employing the SLAKE [37] and VQA-RAD [38] datasets.
 Following METER’s evaluation metrics, we use Closed Accuracy, Open Accuracy, and Overall Accuracy to assess each model’s performance.
-Whole-slide image analysis
-To explore our model’s ability to understand whole-slide images, we conduct experiments on 6 related datasets for 9 subtasks in Cancer Diagnosis, Molecular Subtyping, and Survival Prediction, each of which can be regarded as a classification task, with a whole-slide image as input to the vision encoder and a predicted label as output.
-Given the high resolution of whole-slide images, we follow the conventional approach by segmenting the whole-slide image into smaller patches and using each pre-trained vision encoder to extract features from each image patch.
+Pathology whole-slide image analysis.
+To explore our model’s ability to understand pathology whole-slide images, we conduct experiments on 6 related datasets for 9 subtasks in cancer diagnosis, mutation prediction, and survival prediction, each of which can be regarded as a classification task, with a whole-slide image as input to the vision encoder and a predicted label as output.
+Given the high resolution of whole-slide images, we follow the conventional approach by splitting the whole-slide image into smaller regions and using each pre-trained vision encoder to extract features from each image region.
 These features are input into the ABMIL model [39] for multiple instance learning (MIL).
 The model outputs a predicted label.
-Similar to the Medical Image Diagnosis task, we use the AUC score to evaluate different models in Cancer Diagnosis and Molecular Subtyping, and concordance index (C-index) for survival analysis.
-Medical concept annotation
-We evaluate to what extent can the proposed ConceptCLIP annotate fine-grained concepts for medical images, hence facilitating the development of concept-based explainable artificial intelligence.
+Similar to the medical image diagnosis task, we use the AUC score to evaluate different models in cancer diagnosis and molecular subtyping, and the concordance index (C-index) for survival prediction.
+Medical concept annotation.
+We evaluate to what extent the proposed ConceptCLIP can annotate fine-grained concepts for medical images, hence facilitating the development of concept-based explainable artificial intelligence.
 The performance of medical concept annotation is evaluated using the concept labels of the datasets.
 For each concept, two positive and negative prompts are designed to calculate the similarity with input images to get the predicted output.
-Specifically, we adopt five datasets with clinical concept annotations to evaluate the zero-shot concept annotation performance, including dermoscopic images, blood cell images (hematology), breast ultrasound, and chest CT.
-Derm7pt [10]: A dermoscopic image dataset containing 1,011 images with clinical concepts for melanoma skin lesions in dermatology.
-The test set is adopted for evaluation.
-Only the dermoscopic images are considered in this paper for concept annotations.
-The seven clinical concepts include “Pigment network”, “Dots and globules”, “Pigmentation”, “Streaks”, “Regression structures”, “Blue-whitish veil”, and “Vascular structures”.
-SkinCon [9]: A skin disease dataset densely annotated by experts for fine-grained model debugging and analysis.
-We choose 22 concepts that have at least 50 images representing the concept from the Fitzpatrick 17k (F17k) [45] subset.
-350 images are sampled for evaluation.
-The clinical concepts include “Papule”, “Plaque”, “Pustule”, “Bulla”, “Patch”, “Nodule”, “Ulcer”, “Crust”, “Erosion”, “Atrophy”, “Exudate”, “Telangiectasia”, “Scale”, “Scar”, “Friable”, “Dome-shaped”, “Brown (Hyperpigmentation)”, “White (Hypopigmentation)”, “Purple”, “Yellow”, “Black”, and “Erythema”.
-WBCAtt [11]: This dataset is a densely annotated dataset for white blood cell (WBC) recognition, containing 11 morphological attributes for 10,298 cell images.
-The test set which contains 3,099 images is used for evaluation.
-There are 11 concepts, including “Cell size”,“Cell shape”,“Nucleus shape”,“Nuclear cytoplasmic ratio”,“Chromatin density”, “Cytoplasm vacuole”, “Cytoplasm texture”, “Cytoplasm colour”, “Granule type”, “Granule colour”, and “Granularity”.
-BrEaST [12]: A breast ultrasound dataset containing 256 images with fine-grained concepts for breast lesion analysis.
-Seven BI-RADS descriptors are adopted as the evaluated concepts, including “Shape”, “Magrin”, “Echogenicity”, “Posterior features”, “Halo”, “Calcifications”, and “Skin thickening”.
-LUNA16 [13]: A curated version of LIDC-IDRI [46] dataset, which contains lung images with clinical concepts.
-1,185 slices of lesions from the chest CTs with six clinical concepts are used for evaluation.
-The concepts include “Calcification”, “Lobulation”, “Margin”, “Sphericity”, “Spiculation”, and “Texture”.
-Concept-based data analysis
-To conduct concept-based data analysis, concept presence difference is calculated for datasets of four modalities, including NLM-TB (chest X-ray), CBIS-DDSM (mammography), UBIBC (breast ultrasound), and MedFMC-Colon (pathology).
+Specifically, we adopt five datasets with clinical concept annotations to evaluate the zero-shot concept annotation performance, including dermatology images, blood cell images (pathology), breast ultrasound, and chest CT.
+Inherently interpretable model employment and evaluation.
+Building inherently interpretable models is essential for deploying AI in medicine since it is more trustworthy for doctors and patients during clinical diagnosis.
+In this paper, we employ the Concept Bottleneck Model (CBM) [9] to build the inherently interpretable model based on our method.
+Inspired by previous works [40, 41], we use ConceptCLIP to calculate the cosine similarity between input images and medical concepts, then the similarity is used as the input to the concept bottleneck layer, which maps the concept similarity to the final prediction.
+The concept bottleneck layer is trained using the class labels.
+For concept-class association, the weights of each neuron within the linear layer can be regarded as the contribution to the final prediction.
+To evaluate the effectiveness of our method, we compare ConceptCLIP with black-box models and other vision-language models of both the natural and medical domains on four datasets of different modalities, including dermatology, pathology, ultrasound, and CT.
+Discovery for correlations between concepts and diseases.
+To conduct correlation discovery between concepts and diseases, concept presence difference is calculated for datasets of four modalities, including NLM-TB (chest X-ray), CBIS-DDSM (mammography), UBIBC (breast ultrasound), and MedFMC-Colon (pathology).
 Specifically, for each dataset, assume there are two image sets of different classes (e.g., tuberculosis and normal in NLM-TB dataset), denoted as $I_{+}=\{I_{1+},I_{2+},...,I_{M+}\}$ and $I_{-}=\{I_{1-},I_{2-},...,I_{N-}\}$, and the pre-defined concept list $C$ (e.g., cavitation, consolidation, etc.), where $M$ and $N$ are the numbers of images in the positive and negative image sets, respectively, and $C$ are generated by a large language model (i.e., GPT-4 [42]) since the dataset does not have concept labels.
@@
 A positive value of $D_{c_{i}}$ means concept $c_{i}$ is more likely to appear in the positive image set $I_{+}$, and vice versa.
-Inherently interpretable model employment and evaluation
-Building inherently interpretable models is essential for deploying AI in medicine since it is more trustworthy for doctors and patients during clinical diagnosis.
-In this paper, we employ the Concept Bottleneck Model (CBM) [22] to build the inherently interpretable model based on our method.
-Inspired by previous works [23, 24], we use ConceptCLIP to calculate the cosine similarity between input images and medical concepts, then the similarity is used as the input to the concept bottleneck layer, which maps the concept similarity to the final prediction.
-The concept bottleneck layer is trained using the class labels.
-For concept-class association, the weights of each neuron within the linear layer can be regarded as the contribution to the final prediction.
-To evaluate the effectiveness of our method, we compare ConceptCLIP with other vision-language models of both the natural and medical domains on four datasets of different modalities, including dermoscopy, hematology, ultrasound, and CT.
-Disease-level concept inspection
-As shown in Fig.
-5, pneumonia is selected as the inspected disease.
-For concept inspection on a single dataset, RSNA [48] dataset is adopted, which only includes classes of No finding and Pneumonia.
-For disease-level concept inspection, we collect datasets that also include the two classes, i.e., VinDr-CXR [49] and VinDr-PCXR [50], where VinDr-PCXR focus on the chest X-rays of children.
+For disease-level concept inspection, as shown in Fig.
+4 (i), pneumonia is selected as the inspected disease.
+For concept inspection on a dataset from one institution, the RSNA [43] dataset is adopted, which only includes classes of No finding and Pneumonia.
+For disease-level concept inspection, we collect datasets from various sources that also include the two classes, i.e., VinDr-CXR from the Hospital 108 and the Hanoi Medical University Hospital [44] and VinDr-PCXR from the Phu Tho Obstetric & Pediatric Hospital [45], where VinDr-PCXR focuses on the chest X-rays of children.
 It is noteworthy that VinDr-CXR and VinDr-PCXR contain other classes such as Lung tumor and Tuberculosis, but we only consider the images of No Finding and Pneumonia to analyze the pneumonia disease.
 For each dataset, we use ConceptCLIP to obtain the concept-disease association.
-Then we average the weights of the concept bottleneck layer across all datasets to derive the final top relevant concepts for pneumonia.
-General domain models and the state-of-the-art medical models
-CLIP [4] is a vision-language model pre-trained on large-scale image-text pairs via contrastive learning in the general domain, which showcases strong ability in zero-shot classification.
-We take the version of CLIP-ViT-B-16 pre-trained on LAION-2B [51] under the framework of OpenCLIP [52].
-SigLIP [5] is a variant of the CLIP model, which is pre-trained via contrastive learning using the sigmoid loss. In our experiments, we use the version of ViT-SO400M-14-SigLIP (SigLIP-400M).
-PMC-CLIP [7] is a vision-language model tailored for medicine.
-The authors collected PubMed open access articles and conducted a series of data processing operations, producing a dataset with 1.6 million medical image-text pairs.
-The dataset is used to pre-train a CLIP model, thus resulting in the PMC-CLIP model.
-We use the origin checkpoint [^2] https://huggingface.co/datasets/axiong/pmc_oa_beta/blob/main/checkpoint.pt that the authors provided.
-BiomedCLIP [8] is a medical vision-language model pre-trained on PMC-15M, a dataset with 15 million medical image-text pairs sourced from PMC-OA.
-We take the released checkpoint [^3] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224 for evaluation.
-Implementation details
-We develop ConceptCLIP by utilizing SigLIP-ViT-400M-16444https://huggingface.co/timm/ViT-SO400M-14-SigLIP-384 as the visual encoder and PubMedBERT [^5] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as the text encoder.
-Each input image is resized to $336\times 336$ pixels.
-During pre-training, we employ the AdamW optimizer [53] with a learning rate of $5\times 10^{-4}$.
-The model is initially pre-trained on H800 GPUs with a batch size of 12,288 for 32 epochs on the MedConcept-23M dataset, without the PC-Align loss.
-Subsequently, we conduct further pre-training with the PC-Align loss on this dataset with a batch size of 6,144 and learning rate of $3\times 10^{-4}$ for 20 epochs.
-The weight $\alpha$ during the pre-training is set to 0.5.
-In medical image diagnosis, for zero-shot classification, we set $\beta$ to 0.5 so that the prediction can consider the information from both global information and local information.
-For different datasets, we choose different “Top-K” values and prompts as described in LABEL:tab:zero_shot_settings.
-For linear probes, for all experiments, we use the LogisticRegression function provided in scikit-learn [^6] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, with inverse of regularization strength is 0.316, maximum iteration is 1,000, random state is 1, and all other parameters are default.
-For full fine-tuning experiments, for all experiments, the batch size is 64, epochs is 20, learning rate is 1e-4.
-We use AdamW optimizer for training the model.
-In the visual question answering task, we use AdamW as the optimized.
-For SigLIP-400M and ConceptCLIP, the batch size is 32, learning rate is 1e-6.
-For other models, the batch size is 4, learning rate is 5e-6.
-All experiments have the same epoch number 50.
-In whole-slide image analysis, the batch size is 1, learning rate is 2e-4, epochs are 30, and the cosine learning rate scheduler is used.
+Then we average the weights of the concept bottleneck layer across all considered datasets to derive the final top relevant concepts for pneumonia.
 Metrics
-AUC
+AUC.
 The Area Under the Receiver Operating Characteristic Curve (AUC) is a performance measurement for classification problems at various threshold settings.
 It is widely used in medical image diagnosis and explainable AI tasks to evaluate the ability of a model to distinguish between classes.
-Accurary
-Accuracy (ACC) is the proportion of true results among the total number of cases examined.
-In the context of visual question answering tasks, we use closed accuracy, open-accuracy, and overall accuracy to assess the model’s ability to select the correct answer from a candidate answer set.
-The formula is:
-$\text{ACC}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$
-C-Index
+Accuracy. Accuracy (ACC) quantifies the proportion of correct predictions relative to the total evaluated cases.
+ACC=Number of Correct PredictionsTotal Number of $\text{ACC}=\frac{\text{Number of Correct Predictions}}{\text{Total Number of Predictions}}.$
+For visual question answering (VQA) tasks, we evaluate three variants: closed accuracy, which measures performance on predefined answer candidates (e.g., multiple-choice questions); open accuracy, assessing free-form answer generation without predefined options; and overall accuracy, providing a combined assessment of both closed and open settings.
+C-Index.
 The concordance index (C-Index) measures the predictive accuracy of a survival model.
@@
 It is calculated as the proportion of all usable patient pairs whose predictions and outcomes are concordant.
-Recall
-Recall, or sensitivity, measures the ability of a model to correctly identify all relevant instances.
-In cross-modal retrieval tasks, we use Image-to-Text Recall@1,5,10 and Text-to-Image Recall@1,5,10 to assess the model’s retrieval performance at different levels.
-The formula is:
-$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$
-BLEU
-The Bilingual Evaluation Understudy (BLEU) score [54] is a metric for evaluating the quality of text generated by a machine, such as translations or reports.
-It is used for medical report generation to assess the accuracy and fluency of the generated text.
-The formula for BLEU is:
-$\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$
-where $p_{n}$ is the precision for n-grams, $w_{n}$ are weights, and BP is the brevity penalty.
-METEOR
+Recall.
+Recall (sensitivity) quantifies the proportion of relevant instances successfully retrieved from the total relevant population.
+In classification tasks, this metric evaluates a model’s ability to minimize false negatives (FN), defined as:
+$\text{Recall}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}},$
+[…]
+BLEU.
+The Bilingual Evaluation Understudy (BLEU) score [46] quantifies the similarity between machine-generated text and reference human-authored texts by measuring lexical and syntactic overlap.
+Originally developed for machine translation, BLEU is widely adopted in medical report generation to evaluate both semantic fidelity (alignment with clinical content) and fluency (linguistic coherence).
+The score is computed as:
+$\text{BLEU}=\underbrace{\text{BP}}_{\text{Brevity Penalty}}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$
+where $p_{n}$ represents the modified n-gram precision (the fraction of machine-generated n-grams that match the reference texts), $w_{n}$ are weighting factors (typically $w_{n}=\frac{1}{N}$ for n-grams up to length $N=4$), and BP is the brevity penalty $\min\left(1,e^{1-\frac{\text{reference length}}{\text{generated length}}}\right)$ that prevents artificially high scores for shorter outputs.
+Higher BLEU scores (range: 0–1) indicate stronger alignment with reference texts.
+In medical applications, this metric ensures generated reports preserve critical clinical terminology (via $n$-gram precision) while maintaining natural narrative structure (via brevity penalty).
+METEOR.
 The Metric for Evaluation of Translation with Explicit ORdering (METEOR) [47] is another metric for evaluating text generation quality, focusing on precision, recall, and alignment of phrases.
 It is used for medical report generation to provide a more nuanced assessment of generated text quality.
-ROUGE-L
-The Recall-Oriented Understudy for Gisting Evaluation (ROUGE-L) [56] measures the longest common subsequence (LCS) between the generated text and reference texts.
+ROUGE-L.
+The Recall-Oriented Understudy for Gisting Evaluation (ROUGE-L) [48] measures the longest common subsequence (LCS) between the generated and reference text sequences.
 It is used for evaluating medical report generation to capture the overlap in content.
@@
 […]
-CIDEr
-Consensus-based Image Description Evaluation (CIDEr) [57] is a metric designed to assess the similarity of a generated text to multiple reference texts, emphasizing consensus.
+CIDEr.
+Consensus-based Image Description Evaluation (CIDEr) [49] is a metric designed to assess the similarity of a generated text to multiple reference text sequences, emphasizing consensus.
 It is used for medical report generation to evaluate the relevance and quality of the generated content.
-Clinical efficacy
-Clinical Efficacy measures the practical effectiveness of a generated medical report in a clinical setting, focusing on its impact on clinical outcomes and decision-making.
-It includes precision, recall, and F1, by annotating the generated report and comparing it with the ground truth report via 14 disease classification labels in CheXbert [58].
-Micro precision calculates the precision of the model by considering all true positives and false positives across all classes, providing a detailed view of the model’s performance:
-Micro $\text{Micro precision}=\frac{\sum\text{TP}}{\sum\text{TP}+\sum\text{FP}}$
-Micro recall aggregates the true positives and false negatives across all classes to compute recall, offering a comprehensive assessment of the model’s ability to identify relevant instances:
-Micro $\text{Micro recall}=\frac{\sum\text{TP}}{\sum\text{TP}+\sum\text{FN}}$
-Micro F1 is the harmonic mean of Micro Precision and Micro Recall, providing a balanced measure of the model’s accuracy and completeness across all classes:
+Clinical Efficacy.
+Clinical efficacy assesses the practical utility of AI-generated medical reports by measuring their concordance with clinical standards and potential impact on diagnostic decisions.
+We employ micro-averaged precision, recall, and F1 scores computed using the CheXbert classifier [50], which identifies 14 standardized radiological observations (including pneumothorax and edema) in both generated and reference reports.
+The micro-averaged approach aggregates performance across all disease classes: Micro Precision measures the fraction of correctly identified findings among all system-predicted abnormalities (true positives/[true positives + false positives]), while Micro Recall calculates the proportion of reference report findings correctly detected by the system (true positives/[true positives + false negatives]).
+The Micro F1 score, defined as:
 […]
+provides a balanced measure of diagnostic performance by combining precision and recall.
+This comprehensive evaluation framework ensures clinically meaningful assessment of report quality, where high precision minimizes overdiagnosis and high recall reduces missed findings.
+These metrics reflect system-wide reliability: high precision minimizes overdiagnosis (false positives), while high recall mitigates missed diagnoses (false negatives).
+Micro-averaging prioritizes per-case accuracy over per-class balancing, mirroring clinical workflows where all observations in a report hold equal urgency.
 Data availability
-The dataset used for pre-training is described in Section Method. The datasets used for evaluation are listed in Extended Data Table LABEL:tab:datasets.
+The pre-training dataset, described in Section Method, is derived from publicly available sources and will be released to the research community upon publication to ensure reproducibility.
+All evaluation datasets referenced in Extended Data Table LABEL:tab:datasets are publicly accessible benchmarks, with complete citation information provided to enable direct access by other researchers.
 Code availability
-The implementation of ConceptCLIP will be available at https://github.com/JerrryNie/ConceptCLIP. The weights of ConceptCLIP will be released at https://huggingface.co/JerrryNie/ConceptCLIP.
+The implementation of ConceptCLIP is available at https://github.com/JerrryNie/ConceptCLIP. The weights of ConceptCLIP is released at https://huggingface.co/JerrryNie/ConceptCLIP.
 Author contribution
@@
 conceived and designed the work.
-Y.N., S.H., Y.B., Y.W., and Z.C.
+Y.N., S.H., Y.B., Y.W., Zhixuan C., Zhiyuan C., and H.W.
 collected the data for pre-training and downstream task evaluation.
@@
 and S.H.
-contributed to the pre-training technical implementation.
+contributed to the technical implementation of the pre-training process.
 Y.B.
@@
 evaluated whole-slide image analysis tasks.
-Z.C.
+Zhixuan C.
 evaluated medical report generation tasks.
@@
 wrote the manuscript with inputs from all authors.
+Xi.W., L.L., M.W., Xian.W., R.C.K.C., Y.M.L., Y.Z., and P.R.
+provided suggestions on the framework and experiments.
 All authors reviewed and approved the final paper.
@@
 The authors have no conflicts of interest to declare.
+Ethics declarations
+This study has been reviewed and approved by the Human and Artefacts Research Ethics Committee (HAREC). The protocol number is HREP-2024-0212.
+Acknowledgements
+This work was supported by the Hong Kong Innovation and Technology Commission (Project No.
+MHP/002/22, GHP/006/22GD and ITCPD/17-9), HKUST (Project No.
+FS111), and the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference Number: T45-401/22-N),
 Appendix A Extended Data
-Table 1: Medical image analysis datasets.
-Dataset Website Description
+Table 1: Medical image analysis datasets grouped by task. Click the hyperlink in the “link” column to access the corresponding dataset.
+Dataset Description Link
 Binary Classification
-SIIM-ACR [59] https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation A binary classification dataset containing chest radiographic images.
-Covid-CXR2 [60] https://www.kaggle.com/datasets/andyczhao/covidx-cxr2 A dataset of over 16,000 chest X-ray images from more than 15,100 patients across 51 countries, including 2,300 positive COVID-19 images, designed to differentiate between no pneumonia, non-COVID-19 pneumonia, and COVID-19 pneumonia in the COVIDx V8A dataset, while the COVIDx V8B dataset focuses on detecting COVID-19 positive and negative cases.
-NLM-TB [25] https://www.ncbi.nlm.nih.gov/pmc/articles/PMC4256233/ A radiology image dataset with 800 total images with the label of normal or tuberculosis.
-LC25000 (Colon) [61] https://github.com/tampapath/lung_colon_image_set A dataset comprises 10,000 pathology images from colon tissue, featuring benign tissues and adenocarcinomas.
-UBIBC [62] https://www.kaggle.com/datasets/vuppalaadithyasairam/ultrasound-breast-images-for-breast-cancer This dataset consists of ultrasound images related to breast cancer, both benign and malignant.
-PCam200 [63] https://github.com/enigmanx20/patchtcga A dataset of pathological H&E images created from the Camelyon2016 challenge dataset [64].
-RSNA [48] https://www.rsna.org/rsnai/ai-image-challenge/rsna-pneumonia-detection-challenge-2018 A dataset of about 30,000 frontal view chest X-ray images, each labeled for binary classification to indicate the presence of pneumonia.
-Brain Tumor CT [65] https://www.kaggle.com/datasets/murtozalikhon/brain-tumor-multimodal-image-ct-and-mri A dataset of CT scans for brain tumor detection and analysis.
+SIIM-ACR [51] A binary classification dataset containing chest radiographic images.
+Covid-CXR2 [52] A dataset of over 16,000 chest X-ray images from more than 15,100 patients across 51 countries, including 2,300 positive COVID-19 images, designed to differentiate between no pneumonia, non-COVID-19 pneumonia, and COVID-19 pneumonia in the COVIDx V8A dataset, while the COVIDx V8B dataset focuses on detecting COVID-19 positive and negative cases.
+NLM-TB [26] A radiology image dataset with 800 total images with the label of normal or tuberculosis.
+LC25000 (Colon) [53] A dataset comprises 10,000 pathology images from colon tissues, featuring benign tissues and adenocarcinomas.
+UBIBC [54] A dataset consists of ultrasound images related to breast cancer, including benign and malignant.
+PCam200 [55] A dataset of pathological H&E images created from the Camelyon2016 challenge dataset [56].
+RSNA [43] A dataset of about 30,000 frontal view chest X-ray images, each labeled for binary classification to indicate the presence of pneumonia.
+Brain Tumor CT [57] A dataset of CT scans for brain tumor detection and analysis.
 It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
-Brain Tumor MRI [65] https://www.kaggle.com/datasets/murtozalikhon/brain-tumor-multimodal-image-ct-and-mri A dataset of MRI scans, for brain tumor detection and analysis.
-It features high-resolution images from multiple patients, labeled with tumor types (e.g., glioma, meningioma) and their locations in the brain, designed to aid in developing AI models for automatic detection, classification, and segmentation of brain tumors.
-DDSM [66] https://www.kaggle.com/datasets/skooch/ddsm-mammography A dataset containing 55,890 pre-processed images from the DDSM [67] and CBIS-DDSM [68] datasets, formatted as 299x299 pixel tfrecords for TensorFlow, with an imbalance of 14% positive and 86% negative examples, incorrectly split into test and validation sets.
-Breast Cancer [69] https://www.kaggle.com/datasets/hayder17/breast-cancer-detection/ A dataset of 3,383 annotated mammogram images focused on breast tumors, exported from Roboflow, designed for building and testing deep learning models for tumor detection.
+Brain Tumor MRI [57] This dataset is similar to the Brain Tumor CT dataset [57] but with the MRI modality.
+DDSM [58] A dataset containing 55,890 pre-processed images from the DDSM [59] and CBIS-DDSM [60] datasets, formatted as 299$\times$299 pixel TFRecords for TensorFlow, with 14% positive and 86% negative examples, split into test and validation sets.
+Breast Cancer [61] A dataset of 3,383 annotated mammogram images focused on breast tumors, exported from Roboflow, designed for building and testing deep learning models for tumor detection.
+HMC-QU [62] A dataset created by Hamad Medical Corporation, Tampere University, and Qatar University includes 2D echocardiography recordings from 2018 and 2019 for heart attack detection and left ventricle wall segmentation, with recordings at 25 fps and resolutions from 422$\times$636 to 768$\times$1024 pixels.
 Multi-Label Classification
-RFMiD2 [70] https://www.mdpi.com/2306-5729/8/2/29 A multi-label dataset of fundus images annotated by three eye specialists.
-MedFMC (Colon) [71] https://github.com/openmedlab/MedFM A dataset focused on facilitating the detection of early-stage cancer cells in tissue slides, enabling pathologists to classify and quantify cancerous regions for lesion and non-lesion classes.
-VinDr-Mammo [72] https://vindr.ai/datasets/mammo A large-scale benchmark dataset for computer-aided detection and diagnosis in full-field digital mammography.
-VinDr-PCXR [50] https://vindr.ai/datasets/pediatric-chest-x-ray A dataset focuses on pediatric chest X-rays for the interpretation of common thoracic diseases.
-VinDr-CXR [49] https://vindr.ai/datasets/cxr A dataset of chest x-ray images containing 18,000 high-quality postero-anterior scans with detailed localization of 22 critical findings and classification of 6 thoracic diseases, annotated by experienced radiologists from major hospitals in Vietnam.
-To make the official training and testing label consistent, we filter out part of inconsistent training data, thus resulting in 32,976 training images and 3,000 testing images.
-VinDr-SpineXR [73] https://vindr.ai/datasets/spinexr A dataset of annotated medical images for detecting and classifying spinal lesions from radiographs.
-NIH ChestX-ray14 [74] https://nihcc.app.box.com/v/ChestXray-NIHCC A dataset comprising 112,120 frontal-view X-ray images from 30,805 unique patients, featuring fourteen disease labels identified through NLP techniques, collected between 1992 and 2015.
-CheXpert [58] https://stanfordmlgroup.github.io/competitions/chexpert/ A dataset of chest radiographs from 65,240 patients.
+RFMiD2 [63] A multi-label dataset of fundus images annotated by three eye specialists.
+MedFMC (Colon) [64] A dataset focused on facilitating the detection of early-stage cancer cells in tissue slides, enabling pathologists to classify and quantify cancerous regions.
+VinDr-Mammo [65] A large-scale benchmark dataset for computer-aided detection and diagnosis in full-field digital mammography.
+VinDr-PCXR [45] A dataset focuses on pediatric chest X-rays for the interpretation of common thoracic diseases.
+VinDr-CXR [44] A dataset of chest X-Ray images containing 18,000 high-quality postero-anterior scans with detailed localization of 22 critical findings and classification of 6 thoracic diseases, annotated by experienced radiologists from major hospitals in Vietnam.
+VinDr-SpineXR [66] A dataset of annotated medical images for detecting and classifying spinal lesions from radiographs.
+NIH ChestX-ray14 [67] A dataset comprising 112,120 frontal-view X-ray images from 30,805 unique patients, featuring fourteen disease labels identified through natural language processing (NLP) techniques, collected between 1992 and 2015.
+CheXpert [50] A dataset of chest radiographs from 65,240 patients.
 Multi-Class Classification
-DRD [75] https://www.kaggle.com/competitions/diabetic-retinopathy-detection A dataset consists of high-resolution retina images labeled by subject ID and eye side, with each image rated for diabetic retinopathy severity on a scale of 0 (no DR) to 4 (proliferative DR) by a clinician.
-LC25000 (Lung) [61] https://github.com/tampapath/lung_colon_image_set A dataset consists of 15,000 pathology images from lung tissue, including benign tissues and various carcinomas.
-MedFMC (Chest) [71] https://github.com/openmedlab/MedFM A dataset for screening thoracic diseases encompasses 19 common thoracic abnormalities.
-MedFMC (Endo) [71] https://github.com/openmedlab/MedFM A dataset aimed at enhancing the automatic detection and classification of four different lesion types in colonoscopy images, facilitating early diagnosis of colorectal cancer.
-HAM10000 [76] https://www.kaggle.com/datasets/kmader/skin-cancer-mnist-ham10000 This dataset is composed of dermatoscopic images of pigmented lesions.
-BUSBRA [77] https://zenodo.org/records/8231412 A dataset of anonymized breast ultrasound images from 1,064 patients, including biopsy-proven tumors, BI-RADS annotations, and ground truth delineations of tumoral and normal regions.
-WCE [78] https://www.kaggle.com/datasets/francismon/curated-colon-dataset-for-deep-learning A dataset of colon disease images containing curated samples for training and testing, derived from the Kvasir [79] and ETIS-Larib-Polyp DB [80] datasets.
-Fundus JSIEC [81] https://www.kaggle.com/datasets/linchundan/fundusimage1000 A dataset of 1,000 fundus images belonging to 39 classes, sourced from the Joint Shantou International Eye Centre in Shantou city, Guangdong province, China.
-HyperKvasir [82] https://datasets.simula.no/hyper-kvasir/ A dataset consisting of 10,662 labeled images in JPEG format, categorized into 23 classes of medical findings, with each class represented by a specific folder, highlighting the imbalance in the number of images per class.
-Kvasir [79] https://datasets.simula.no/kvasir/ A dataset of annotated and verified gastrointestinal tract images, featuring various classes of anatomical landmarks and pathological findings, organized in separate folders with resolutions ranging from 720x576 to 1920x1072 pixels, suitable for tasks such as image retrieval and machine learning.
-ODIR [83] https://www.kaggle.com/datasets/andrewmvd/ocular-disease-recognition-odir5k A dataset of 5,000 patients featuring age, color fundus photographs from both left and right eyes, along with doctors’ diagnostic keywords, with 6,392 samples used for training.
-BUID [84] https://www.kaggle.com/datasets/aryashah2k/breast-ultrasound-images-dataset A dataset of 780 breast ultrasound images from 600 female patients aged 25 to 75, collected in 2018, categorized into three classes: normal, benign, and malignant, with ground truth images included.
-PAD-UFES-20 [85] https://www.notion.so/c1e0ff8f36814b94b205ee919e2e3ff8?pvs=21 A dataset of skin lesion images labeled into six categories: Basal Cell Carcinoma, Squamous Cell Carcinoma, Actinic Keratosis, Seborrheic Keratosis, Melanoma, and Nevus.
-OCTMNIST [86] https://medmnist.com/ A dataset for multi-class classification of retinal OCT images.
-Breast Tumor MRI [87] https://www.kaggle.com/datasets/masoudnickparvar/brain-tumor-mri-dataset A combined dataset (including three datasets: figshare [88], SARTAJ dataset [89].
+DRD [68] A dataset consists of high-resolution retina images labeled by subject ID and eye side, with each image rated for diabetic retinopathy severity on a scale of 0 (no DR) to 4 (proliferative DR) by ophthalmologists.
+LC25000 (Lung) [53] A dataset consists of 15,000 pathology images from lung tissue, including benign tissues and various carcinomas.
+MedFMC (Chest) [64] A dataset for screening thoracic diseases encompasses 19 common thoracic abnormalities.
+MedFMC (Endo) [64] A dataset aimed at enhancing the automatic detection and classification of four different lesion types in colonoscopy images, facilitating early diagnosis of colorectal cancer.
+HAM10000 [69] A dataset consists of dermatoscopic images of pigmented lesions.
+BUSBRA [70] A dataset of anonymized breast ultrasound images from 1,064 patients, including biopsy-proven tumors, BI-RADS annotations, and ground truth delineations of tumoral and normal regions.
+WCE [71] A dataset of colon disease images containing curated samples for training and testing, derived from the Kvasir [72] and ETIS-Larib-Polyp DB [73] datasets.
+Fundus JSIEC [74] A dataset of 1,000 fundus images belonging to 39 classes, sourced from the Joint Shantou International Eye Centre in Shantou city, Guangdong province, China.
+HyperKvasir [75] A dataset consisting of 10,662 labeled images in the JPEG format, categorized into 23 classes of medical findings, highlighting the imbalance in the number of images per class.
+Kvasir [72] A dataset of annotated and verified gastrointestinal tract images, featuring various classes of anatomical landmarks and pathological findings, with resolutions ranging from 720$\times$576 to 1920$\times$1072 pixels.
+ODIR [76] A dataset of 5,000 patients featuring color fundus photographs from both left and right eyes, along with ophthalmologists’ diagnostic keywords, with 6,392 samples used for training.
+BUID [77] A dataset of 780 breast ultrasound images from 600 female patients aged 25 to 75, collected in 2018, categorized into three classes: normal, benign, and malignant.
+PAD-UFES-20 [78] A dataset of skin lesion images labeled into six categories: basal cell carcinoma, squamous cell carcinoma, actinic keratosis, seborrheic keratosis, melanoma, and nevus.
+OCTMNIST [79] A dataset for multi-class classification of retinal OCT images.
+Breast Tumor MRI [80] A combined dataset (including three datasets: figshare [81], SARTAJ dataset [82].
 and Br35H dataset [83]) of 7,023 human brain MRI images classified into four classes: glioma, meningioma, no tumor, and pituitary.
-The no tumor class images are from the Br35H dataset.
-Retinal OCT [91] https://www.kaggle.com/datasets/obulisainaren/retinal-oct-c8 A dataset of high-quality retinal OCT images categorized into 8 classes of retinal diseases, designed for research and model training in classification using machine learning and deep learning.
-COVIDxCT [92] https://www.kaggle.com/datasets/hgunraj/covidxct/data A dataset consisting of two variants, ”A” with confirmed COVID-19 cases and ”B” which includes weakly verified cases, designed for training, validation, and testing of CT scans.
-We choose the “A” set in our experiment.
+The “no tumor” class images are from the Br35H dataset.
+Retinal OCT [84] A dataset of high-quality retinal OCT images categorized into 8 classes of retinal diseases.
+COVIDxCT [85] A dataset of CT scans consisting of two variants, “A” with confirmed COVID-19 cases and “B” which includes weakly verified cases. We choose the “A” set in our experiment.
 Retrieval
-PMC-9K - This newly curated dataset consists of a cleaned, held-out collection of 9,222 image-text pairs in MedConcept-23M.
+PMC-9K newly curated dataset consisting of a cleaned, held-out collection of 9,222 image-text pairs in MedConcept-23M.
 It is designed to comprehensively evaluate the cross-modal retrieval capabilities of ConceptCLIP across various modalities.
-Quilt-1M [16] https://quilt1m.github.io/ This dataset is a large-scale histopathology collection containing 768,826 image-text pairs.
-For evaluation purposes, we filter out the data sourced from PubMed Central to avoid data leakage and utilize a held-out subset comprising pathology image-text pairs to assess the retrieval performance.
+-
+Quilt-1M [16] A large-scale histopathology collection containing 768,826 image-text pairs.
+For evaluation purposes, we filter out the data sourced from PubMed Central and utilize a held-out subset comprising pathology image-text pairs to assess the retrieval performance.
 Medical Report Generation
-MIMIC-CXR [39] https://physionet.org/content/mimic-cxr/2.1.0/ A dataset contains 371,920 chest X-rays linked to 227,943 imaging studies from 65,079 patients.
-IU X-Ray [40] https://www.kaggle.com/datasets/raddar/chest-xrays-indiana-university/data A dataset consists of 7,470 pairs of chest X-ray images and their corresponding diagnostic reports.
+MIMIC-CXR [17] A dataset contains 371,920 chest X-rays linked to 227,943 imaging studies from 65,079 patients.
+IU X-Ray [18] A dataset consists of 7,470 pairs of chest X-ray images and their corresponding diagnostic reports.
 Visual Question Answering
-VQA-RAD [43] https://osf.io/89kps/ A dataset called VQA-RAD is manually constructed, featuring 3,064 question-answer pairs where clinicians ask natural questions about radiology images and provide reference answers.
-SLAKE [42] https://www.med-vqa.com/slake/ A bilingual radiology VQA dataset that includes 642 images and 14,000 questions. We only use the English part.
+VQA-RAD [38] A dataset manually constructed, featuring 3,064 question-answer pairs where clinicians ask questions about radiology images and provide reference answers.
+SLAKE [37] A bilingual radiology VQA dataset that includes 642 images and 14,000 questions. We only use the English part.
 Cancer Diagnosis
-BRACS-3 [93] https://www.bracs.icar.cnr.it/ A dataset contains 6 different subtypes of lesions including also images representing atypical lesions.
+BRACS-3 [86] A dataset contains 6 different subtypes of lesions, including images representing atypical lesions.
 Histological images representing normal tissue samples are also included.
 In this setting, a breast tumor according to a pathology image can be classified into “benign”, “atypical”, or “malignant”.
-BRACS-7 [93] https://www.bracs.icar.cnr.it/ A dataset contains 6 different subtypes of lesions including also images representing atypical lesions.
-Histological images representing normal tissue samples are also included.
+BRACS-7 [86] A dataset that contains the same images as BRACS-3 but with a different label set.
 In this setting, a breast tumor according to a pathology image can be classified into “normal”, “pathological benign”, “usual ductal hyperplasia”, “flat epithelial atypia”, “atypical ductal hyperplasia”, “ductal carcinoma in situ”, or “invasive carcinoma”.
-BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset features 985 histopathology whole slide images of Breast Invasive Carcinoma, including 787 Invasive Ductal Carcinoma and 198 Invasive Lobular Carcinoma cases.
-NSCLC [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset contains 1,053 histopathology slides of Non-Small Cell Lung Cancer, including 541 lung adenocarcinoma and 512 lung squamous cell carcinoma cases.
-Camelyon [64, 95] https://camelyon16.grand-challenge.org/https://camelyon17.grand-challenge.org/ A dataset based on CAMELYON16 [64] and CAMELYON17 [95], which evaluates new and existing algorithms for automated detection and classification of breast cancer metastases in whole-slide images of histological lymph node sections.
-Molecular Subtyping
-BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A dataset includes 985 histopathology whole slide images of breast invasive carcinoma, specifically 787 Invasive Ductal Carcinoma and 198 Invasive Lobular Carcinoma cases.
+NSCLC [87] A dataset contains 1,053 histopathology slides of non-small cell lung cancer, including 541 lung adenocarcinoma and 512 lung squamous cell carcinoma cases.
+Camelyon [56, 88] A dataset based on CAMELYON16 [56] and CAMELYON17 [88], which evaluates new and existing algorithms for automated detection and classification of breast cancer metastases in whole-slide images of histological lymph node sections.
+Mutation Prediction
+BRCA-TP53 [87] This dataset comprises molecular and clinical data from the Breast Invasive Carcinoma (BRCA) cohort of The Cancer Genome Atlas (TCGA), with specialized annotation for TP53-mutant tumors, the most frequently altered tumor suppressor in this malignancy.
+BRCA-TNN [87] This dataset comprises molecular and clinical profiles of triple-negative breast tumors (TNBC) from the TCGA BRCA cohort, defined by immunohistochemical absence of estrogen receptor (ER), progesterone receptor (PR), and HER2 amplification.
+LUAD-TP53 [87] This dataset provides comprehensive molecular profiling of TP53-mutant lung adenocarcinoma cases from the TCGA cohort.
 Survival Prediction
-BRCA [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis.
-We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
-LUAD [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis.
-We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
-LUSC [94] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A TCGA dataset for survival analysis.
-We employed case- and label-stratified splits with 7:1:2 training, validation, and testing sets over 400 cases.
+BRCA [87] A TCGA dataset of breast carcinoma for survival prediction.
+LUSC [87] A dataset of the Lung Squamous Cell Carcinoma LUSC cohort from the TCGA for survival prediction.
 Table 2: Zero-shot experimental settings for medical image analysis datasets.
-Dataset Classes Prompts Top-K
-SIIM-ACR [59] No Finding, Pneumothorax a chest radiology presents {} 256
-RFMiD2 [70] Neovascularization, Macular Edema, Myopia, Retinal Traction, Coloboma, Choroidal folds, Tortuous Vessels, Retinitis Pigmentosa, Retinal pigment epithelium changes, Optic Disc Pallor, Media Haze, Retinitis, Preretinal Hemorrhage, Asteroid Hyalosis, Drusens, Hemorrhagic Pigment Epithelial Detachment, Branch Retinal Vein Occlusion, Optic Disc Edema, Exudation, Haemorrhagic Retinopathy, Tilted Disc, Tessellation, Retinal Tears, Retinal Detachment, Optic Disc Cupping, Macular Hole, Silicone Oil-Filled Eye, Cotton Wool Spots, Vasculitis, Microaneurysm, Macular Scar, Age-Related Macular Degeneration, Optic neuritis, Anterior Ischemic Optic Neuropathy, Laser Scar, Chorioretinitis, Within Normal Limit, Epiretinal Membrane, Central retinal vein occlusion, Central Serous Retinopathy, Optociliary Shunt an image of {} 4
-DRD [75] No diabetic retinopathy, Mild diabetic retinopathy, Moderate diabetic retinopathy, Severe diabetic retinopathy, Proliferative diabetic retinopathy a detailed view of a retina indicating {};
+Dataset Classes Prompts
+SIIM-ACR [51] No Finding, Pneumothorax a chest radiology presents {}
+RFMiD2 [63] Neovascularization, Macular Edema, Myopia, Retinal Traction, Coloboma, Choroidal Folds, Tortuous Vessels, Retinitis Pigmentosa, Retinal Pigment Epithelium Changes, Optic Disc Pallor, Media Haze, Retinitis, Preretinal Hemorrhage, Asteroid Hyalosis, Drusens, Hemorrhagic Pigment Epithelial Detachment, Branch Retinal Vein Occlusion, Optic Disc Edema, Exudation, Haemorrhagic Retinopathy, Tilted Disc, Tessellation, Retinal Tears, Retinal Detachment, Optic Disc Cupping, Macular Hole, Silicone Oil-Filled Eye, Cotton Wool Spots, Vasculitis, Microaneurysm, Macular Scar, Age-Related Macular Degeneration, Optic Neuritis, Anterior Ischemic Optic Neuropathy, Laser Scar, Chorioretinitis, Within Normal Limit, Epiretinal Membrane, Central Retinal Vein Occlusion, Central Serous Retinopathy, Optociliary Shunt an image of {}
+DRD [68] No Diabetic Eetinopathy, Mild Diabetic Retinopathy, Moderate Diabetic Retinopathy, Severe Diabetic Retinopathy, Proliferative Diabetic Retinopathy a detailed view of a retina indicating {};
 a close-up of a retina highlighting {}
-32
-Covid-CXR2 [60] No Finding, Covid-19 a radiographic representation assessing for {} 256
+COVID-CXR2 [52] No Finding, COVID-19 a radiographic representation assessing for {}
 NLM-TB [26] Normal, Tuberculosis a close-up view of a chest x-ray presenting evidence of {};
@@
 an annotated image from a chest x-ray showing signs of {}
-32
 LC25000 (Colon) [53] Normal Colonic Tissue, Colon Adenocarcinomas the histopathological image illustrates that of {};
@@
 a pathological image highlighting a {}
-256
 LC25000 (Lung) [53] Lung Squamous Cell carcinomas, Lung Adenocarcinomas, Normal Lung Tissue histopathological image contains {};
@@
 an annotated histopathological image representing a {}
-256
-MedFMC (Chest) [71] pleural effusion, nodule, pneumonia, cardiomegaly, hilar enlargement, fracture old, fibrosis, aortic calcification, tortuous aorta, thickened pleura, TB, pneumothorax, emphysema, atelectasis, calcification, pulmonary edema, increased lung markings, elevated diaphragm, consolidation lung situation of {} 2
+MedFMC (Chest) [64] Pleural Effusion, Nodule, Pneumonia, Cardiomegaly, Hilar Enlargement, Fracture Old, Fibrosis, Aortic Calcification, Tortuous Aorta, Thickened Pleura, TB, Pneumothorax, Emphysema, Atelectasis, Calcification, Pulmonary Edema, Increased Lung Markings, Elevated Diaphragm, Consolidation lung situation of {}
 MedFMC (Colon) [64] Tumor, Normal this slide features an annotated section indicating a {};
 the image illustrates a {} in the context of colon tissue
-32
 MedFMC (Endo) [64] Ulcer, Erosion, Polyp, Tumor a diagnostic endoscopy showing features of {};
 an endoscopic finding suggestive of {}
-32
-VinDr-Mammo [72] birads negative, breast heterogeneously density, No Finding, breast scattered areas of fibroglandular, birads suspicious malignant, Mass, breast extremely density, birads highly suggestive of malignant, Suspicious Calcification, birads benign, breast almost entirely fatty, Suspicious Lymph Node, Focal Asymmetry, birads probably benign, Asymmetry, Architectural Distortion, Skin Thickening, Global Asymmetry, Nipple Retraction, Skin Retraction an image of {} 32
-HAM10000 [76] actinic keratoses and intraepithelial carcinoma, basal cell carcinoma, benign keratosis-like lesions, dermatofibroma, melanoma, melanocytic nevi, vascular lesions a dermatology has that of {} presented 32
-VinDr-PCXR [50] Bronchiolitis, Pneumonia, Other disease, Bronchitis, Brocho-pneumonia, Tuberculosis, Mediastinal tumor, nan, Situs inversus, Hyaline membrane disease, CPAM, No finding an image of {} 4
+VinDr-Mammo [65] Birads Negative, Breast Heterogeneously Density, No Finding, Breast Scattered Areas of Fibroglandular, Birads Suspicious Malignant, Mass, Breast Extremely Density, Birads Highly Suggestive of Malignant, Suspicious Calcification, Birads Benign, Breast Almost Entirely Fatty, Suspicious Lymph Node, Focal Asymmetry, Birads Probably Benign, Asymmetry, Architectural Distortion, Skin Thickening, Global Asymmetry, Nipple Retraction, Skin Retraction an image of {}
+HAM10000 [69] Actinic Keratoses and Intraepithelial Carcinoma, Basal Cell Carcinoma, Benign Keratosis-Like Lesions, Dermatofibroma, Melanoma, Melanocytic Nevi, Vascular Lesions a dermatology has that of {} presented
+VinDr-PCXR [45] Bronchiolitis, Pneumonia, Other Disease, Bronchitis, Brocho-Pneumonia, Tuberculosis, Mediastinal Tumor, Nan, Situs Inversus, Hyaline Membrane Disease, CPAM, No Finding an image of {}
 BUSBRA [70] Metaplasia Apocrina, Ductal Carcinoma in Situ, Lipoma, Papillary Carcinoma, Hyperplasia, Hamartoma, Ductal Hyperplasia, Cyst, Lymphoma, Invasive Ductal Carcinoma, Fibroadenoma, Lymph Node, Sclerosing Adenosis, Medullary Carcinoma, Adenocarcinoma, Intraductal Papilloma, Duct Ectasia, Mastitis, Fibrosis, Phyllodes Tumor, Lobular Atrophy, Lobular Carcinoma in Situ, Fibrocystic Changes, Mucinous Carcinoma, Galactocele, Invasive Lobular Carcinoma, Proliferative Lesions, Fat Necrosis this is an image of {};
 {} presented in image
-32
-VinDr-CXR [49] Atelectasis, Lung tumor, Lung cavity, Tuberculosis, Pulmonary fibrosis, Clavicle fracture, Lung cyst, Pneumonia, Calcification, No finding, Rib fracture, Pleural thickening, Other disease, Mediastinal shift, Enlarged PA, Nodule/Mass, ILD, COPD, Pneumothorax, Consolidation, Infiltration, Pleural effusion, Other lesion, Cardiomegaly, Emphysema, Lung Opacity an image of {} 2
+VinDr-CXR [44] Atelectasis, Lung Tumor, Lung Cavity, Tuberculosis, Pulmonary Fibrosis, Clavicle Fracture, Lung Cyst, Pneumonia, Calcification, No Finding, Rib Fracture, Pleural Thickening, Other Disease, Mediastinal Shift, Enlarged PA, Nodule/Mass, ILD, COPD, Pneumothorax, Consolidation, Infiltration, Pleural Effusion, Other Lesion, Cardiomegaly, Emphysema, Lung Opacity an image of {}
 UBIBC [54] Benign, Malignant {} show in an ultrasound image of breast cancer;
 an ultrasound image of breast cancer showing signs of {}
-32
-WCE [78] normal, ulcerative colitis, polyps, esophagitis an endoscopic finding suggestive of {} 32
+WCE [71] Normal, Ulcerative Colitis, Polyps, Esophagitis an endoscopic finding suggestive of {}
 PCam200 [55] Normal, Tumor histopathological image contains {};
 an annotated histopathological image representing a {}
-32
-Fundus JSIEC [81] Massive hard exudates, Vitreous particles, Possible glaucoma, Yellow-white spots-flecks, CRVO, BRVO, Large optic cup, Blur fundus without PDR, Bietti crystalline dystrophy, Rhegmatogenous RD, CSCR, Fibrosis, Congenital disc abnormality, Laser Spots, Vessel tortuosity, Maculopathy, RAO, Pathological myopia, DR2, Tessellated fundus, Retinitis pigmentosa, ERM, VKH disease, MH, Cotton-wool spots, Optic atrophy, Severe hypertensive retinopathy, DR3, DR1, Preretinal hemorrhage, Fundus neoplasm, Silicon oil in eye, Blur fundus with suspected PDR, Myelinated nerve fiber, Normal, Disc swelling and elevation, Peripheral retinal degeneration and break, Chorioretinal atrophy-coloboma, Dragged Disc {} presented in image 32
+Fundus JSIEC [74] Massive Hard Exudates, Vitreous Particles, Possible Glaucoma, Yellow-White Spots-Flecks, CRVO, BRVO, Large Optic Cup, Blur Fundus without PDR, Bietti Crystalline Dystrophy, Rhegmatogenous RD, CSCR, Fibrosis, Congenital Disc Abnormality, Laser Spots, Vessel Tortuosity, Maculopathy, RAO, Pathological Myopia, DR2, Tessellated Fundus, Retinitis Pigmentosa, ERM, VKH Disease, MH, Cotton-Wool Spots, Optic Atrophy, Severe Hypertensive Retinopathy, DR3, DR1, Preretinal Hemorrhage, Fundus Neoplasm, Silicon Oil in Eye, Blur Fundus with Suspected PDR, Myelinated Nerve Fiber, Normal, Disc Swelling and Elevation, Peripheral Retinal Degeneration and Break, Chorioretinal Atrophy-Coloboma, Dragged Disc {} presented in image
 HyperKvasir [75] Barrett’s, Barrett’s Short Segment, BBPS 0-1, BBPS 2-3, Cecum, Dyed Lifted Polyps, Dyed Resection Margins, Esophagitis A, Esophagitis B-D, Hemorrhoids, Terminal Ileum, Impacted Stool, Polyps, Pylorus, Retroflex Rectum, Retroflex Stomach, Ulcerative Colitis 0-1, Ulcerative Colitis 1, Ulcerative Colitis 1-2, Ulcerative Colitis 2, Ulcerative Colitis 2-3, Ulcerative Colitis 3, Z-Line this is an image of {};
 {} presented in image
-32
 Kvasir [72] Dyed Lifted Polyps, Dyed Resection Margins, Esophagitis, Normal Cecum, Normal Pylorus, Normal Z Line, Polyps, Ulcerative Colitis an endoscopic finding suggestive of {};
 a snapshot from an endoscopic procedure showing {}
-32
 RSNA [43] No Finding, Pneumonia the chest radiology image of {};
 the lung radiology illustrates that of {}
-8
-VinDr-SpineXR [73] No finding, Osteophytes, Disc space narrowing, Surgical implant, Foraminal stenosis, Other lesions, Vertebral collapse, Spondylolysthesis a spine x-ray image of {} 2
-NIH ChestX-ray14 [74] Atelectasis, Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumonia, Pneumothorax, Consolidation, Edema, Emphysema, Fibrosis, Pleural Thickening, Hernia, No Finding an image of {} 16
-CheXpert [58] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pneumothorax, Pleural Effusion, Pleural Other, Fracture, Support Devices, No Finding an image of {} 4
-ODIR [83] Normal, Diabetes, Glaucoma, Cataract, Age related Macular Degeneration, Hypertension, Pathological Myopia, Other diseases/abnormalities {} presented in retinography 32
-BUID [84] normal, malignant, benign a close-up ultrasound scan showing a {} in the breast 32
+VinDr-SpineXR [66] No Finding, Osteophytes, Disc Space Narrowing, Surgical Implant, Foraminal Stenosis, Other Lesions, Vertebral Collapse, Spondylolisthesis a spine x-ray image of {}
+NIH ChestX-ray14 [67] Atelectasis, Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumonia, Pneumothorax, Consolidation, Edema, Emphysema, Fibrosis, Pleural Thickening, Hernia, No Finding an image of {}
+CheXpert [50] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pneumothorax, Pleural Effusion, Pleural Other, Fracture, Support Devices, No Finding an image of {}
+ODIR [76] Normal, Diabetes, Glaucoma, Cataract, Age Related Macular Degeneration, Hypertension, Pathological Myopia, Other Diseases/Abnormalities {} presented in retinography
+BUID [77] Normal, Malignant, Benign a close-up ultrasound scan showing a {} in the breast
 PAD-UFES-20 [78] Basal Cell Carcinoma, Squamous Cell Carcinoma, Actinic Keratosis, Seborrheic Keratosis, Melanoma, Nevus a dermatological image of {};
 a diagnostic dermatological image illustrating a {}
-16
-OCTMNIST [86] choroidal neovascularization, diabetic macular edema, drusen, normal an image of {} 4
-Breast Tumor MRI [87] glioma, meningioma, pituitary, no tumor a breast mri image of {} 4
-Retinal OCT [91] Age-related macular degeneration, Choroidal neovascularization, Central serous retinopathy, Diabetic macular edema, Macular hole, Drusen, Diabetic retinopathy, Normal a retinal oct image of {} 4
-Brain Tumor CT [65] Healthy, Tumor an image of {} 4
-COVIDxCT [92] Normal, Pneumonia, COVID-19 an image of {} 4
-Brain Tumor MRI [65] Healthy, Tumor an image of {} 4
-DDSM [66] negative, positive the mammography image of {} 4
-Breast Cancer [69] Normal, Tumor the mammography image of {} 4
-Table 3: AUC scores for classification results across different modalities in the zero-shot setting.
-Bold indicates the best result and underline indicates the second best.
-95% CI is included in parentheses.
+OCTMNIST [79] Choroidal Neovascularization, Diabetic Macular Edema, Drusen, Normal an image of {}
+Breast Tumor MRI [80] Glioma, Meningioma, Pituitary, No Tumor a breast mri image of {}
+Retinal OCT [84] Age-Related Macular Degeneration, Choroidal Neovascularization, Central Serous Retinopathy, Diabetic Macular Edema, Macular Hole, Drusen, Diabetic Retinopathy, Normal a retinal oct image of {}
+Brain Tumor CT [57] Healthy, Tumor an image of {}
+COVIDxCT [85] Normal, Pneumonia, COVID-19 an image of {}
+Brain Tumor MRI [57] Healthy, Tumor an image of {}
+DDSM [58] Negative, Positive the mammography image of {}
+Breast Cancer [61] Normal, Tumor the mammography image of {}
+HMC-QU [62] Apical 4 Chamber, Apical 2 Chamber an image of {}
+Figure 5: Concepts from the MedConcept-23M dataset.
+Wordclouds of concepts in captions to qualitatively visualize the concepts in each medical image modality.
+Word size is proportional to its occurrence in MedConcept-23M.
+Common nouns and verbs are ignored.
+Table 3: AUC scores (%) for classification results across different modalities in the zero-shot setting.
+Bold indicates the best results and underline indicates the second best.The 95% CI is included in parentheses.
 Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
@@
 NIH ChestXray14 49.83 (49.72, 49.95) 57.69 (57.19, 58.18) 50.67 (50.60, 50.73) 51.93 (51.66, 52.18) 62.92 (62.37, 63.47)
-CheXpert 50.66 (50.59,50.74) 55.26 (55.06,55.47) 50.15 (50.13,50.17) 50.43 (50.38,50.49) 57.07 (56.86,57.28)
-Average 51.08 (49.76,52.37) 63.77 (62.42,65.02) 57.33 (56.05,58.48) 61.63 (60.43,62.70) 70.40 (69.18,71.50)
+CheXpert 50.66 (50.58, 50.74) 55.26 (55.06, 55.47) 50.15 (50.13, 50.17) 50.43 (50.38, 50.48) 57.06 (56.86, 57.28)
+Average 51.08 (49.76, 52.37) 63.77 (62.41, 65.02) 57.33 (56.05, 58.48) 61.63 (60.43, 62.70) 70.40 (69.17, 71.50)
 Fundus
@@
 Average 42.37 (40.16, 44.56) 53.90 (51.22, 56.43) 52.53 (50.30, 54.84) 52.35 (50.13, 54.58) 61.35 (59.08, 63.77)
-Dermoscopy
+Dermatology
 HAM10000 62.40 (59.99, 64.55) 72.14 (70.16, 74.08) 65.87 (63.52, 67.97) 67.54 (65.49, 69.48) 82.40 (80.91, 83.96)
@@
 BUID 54.00 (47.64, 60.75) 70.47 (64.56, 76.00) 69.86 (64.18, 75.20) 77.45 (71.80, 82.63) 79.75 (73.76, 85.67)
-Average 44.72 (40.99,48.40) 63.47 (60.27,66.57) 57.86 (54.45,61.18) 65.32 (61.45,69.05) 74.83 (71.13,78.56)
+HMC-DQ 49.87 (47.55, 52.27) 69.27 (67.17, 71.53) 58.45 (56.20, 60.58) 46.53 (44.32, 48.88) 75.43 (73.43, 77.38)
+Average 46.01 (43.22, 48.82) 64.92 (62.50, 67.35) 58.00 (55.47, 60.55) 60.62 (57.63, 63.53) 74.98 (72.14, 77.70)
 OCT
@@
 Average 55.34 (53.50, 57.17) 77.59 (76.53, 78.57) 79.04 (77.42, 80.56) 80.25 (78.64, 81.77) 90.98 (90.00, 91.82)
-Table 4: AUC scores for classification results across different modalities in the linear probes setting with 1% training data.
-Bold indicates the best result and underline indicates the second best.
-95% CI is included in parentheses.
+Table 4: AUC scores (%) for classification results across different modalities in the linear probing setting with 1% training data.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
@@
 NIH ChestXray14 60.43 (59.93, 60.96) 63.36 (62.73, 64.01) 67.38 (66.96, 67.84) 67.74 (67.20, 68.29) 71.99 (71.50, 72.46)
-CheXpert 71.51 (71.29,71.73) 77.31 (77.07,77.54) 80.77 (80.53,81.08) 81.87 (81.40,82.10) 84.96 (84.75,85.19)
-Average 64.91 (63.50,66.13) 67.73 (66.36,68.97) 67.20 (65.77,68.48) 67.24 (65.84,68.48) 71.48 (70.19,72.70)
+CheXpert 71.51 (71.30, 71.73) 77.31 (77.07, 77.53) 80.77 (80.54, 81.07) 81.87 (81.61, 82.09) 84.96 (84.76, 85.16)
+Average 64.91 (56.47, 74.46) 67.73 (59.51, 76.46) 67.20 (58.99, 75.73) 67.24 (59.05, 75.84) 71.48 (62.46, 79.56)
 Fundus
@@
 DRD 66.50 (66.02, 66.98) 68.00 (67.57, 68.49) 69.43 (68.95, 69.88) 67.67 (67.22, 68.12) 72.08 (71.66, 72.48)
-Average 54.76 (25.96,28.67) 55.42 (26.34,28.92) 56.22 (26.73,29.37) 55.39 (26.20,29.03) 57.44 (27.32,29.96)
+Average 54.76 (43.01, 66.50) 55.42 (42.83, 68.00) 56.22 (43.00, 69.43) 55.39 (43.11, 67.67) 57.44 (42.79, 72.08)
 Pathology
@@
 PCam200 79.47 (78.84, 80.14) 87.50 (87.00, 88.02) 88.44 (87.98, 88.92) 90.19 (89.77, 90.63) 95.69 (95.40, 95.96)
-Average 89.71 (89.36,90.04) 94.74 (94.50,94.96) 95.95 (95.77,96.12) 96.32 (96.14,96.50) 98.34 (98.23,98.45)
+Average 89.71 (81.58, 97.83) 94.74 (90.06, 98.37) 95.95 (91.17, 99.65) 96.32 (91.82, 99.18) 98.34 (96.65, 99.71)
 Endoscopy
@@
 Kvasir 92.10 (91.62, 92.58) 95.12 (94.81, 95.43) 96.13 (95.83, 96.41) 95.92 (95.66, 96.18) 97.48 (97.24, 97.71)
-Average 90.12 (44.76,45.35) 96.28 (47.97,48.29) 96.72 (48.21,48.51) 97.00 (48.37,48.64) 97.96 (48.84,49.10)
+Average 90.12 (88.13, 92.10) 96.28 (95.12, 97.43) 96.72 (96.13, 97.32) 97.00 (95.92, 98.07) 97.96 (97.48, 98.45)
 Mammography
@@
 DDSM 90.53 (89.82, 91.21) 86.17 (85.41, 86.91) 93.32 (92.79, 93.79) 88.44 (87.70, 89.18) 90.64 (89.98, 91.30)
-Average 64.88 (62.67,67.21) 64.93 (62.58,67.49) 68.41 (66.11,70.72) 66.48 (64.30,68.77) 67.76 (65.45,70.09)
-Dermoscopy
+Average 64.88 (50.08, 90.53) 64.93 (54.25, 86.17) 68.41 (54.85, 93.32) 66.48 (55.28, 88.44) 67.76 (55.35, 90.64)
+Dermatology
 HAM10000 75.44 (73.35, 77.50) 79.87 (78.05, 81.52) 73.79 (71.76, 75.79) 78.90 (77.30, 80.52) 83.35 (82.13, 84.63)
-Average 75.44 (36.67,38.75) 79.87 (39.03,40.76) 73.79 (35.88,37.90) 78.90 (38.65,40.26) 83.35 (41.06,42.31)
+Average 75.44 (73.35, 77.50) 79.87 (78.05, 81.52) 73.79 (71.76, 75.79) 78.90 (77.30, 80.52) 83.35 (82.13, 84.63)
 Ultrasound
 UBIBC 67.23 (63.00, 71.19) 73.01 (69.52, 76.57) 70.06 (66.24, 73.68) 75.58 (72.45, 78.82) 78.06 (74.79, 81.20)
-Average 67.23 (21.00,23.73) 73.01 (23.17,25.52) 70.06 (22.08,24.56) 75.58 (24.15,26.27) 78.06 (24.93,27.07)
+HMC-DQ 77.85 (75.90, 79.52) 89.76 (88.50, 91.03) 97.30 (96.76, 97.81) 89.58 (88.23, 90.86) 96.23 (95.48, 96.92)
+Average 72.54 (67.23, 77.85) 81.39 (73.01, 89.76) 83.68 (70.06, 97.30) 82.58 (75.58, 89.58) 87.15 (78.06, 96.23)
 OCT
@@
 Retinal OCT 81.27 (80.85, 81.67) 81.28 (80.91, 81.63) 84.73 (84.46, 84.98) 77.71 (77.21, 78.18) 84.24 (83.96, 84.51)
-Average 84.38 (83.74,85.02) 85.66 (85.09,86.21) 91.21 (90.92,91.49) 84.63 (84.05,85.21) 90.94 (90.65,91.21)
+Average 84.38 (81.27, 87.50) 85.66 (81.28, 90.03) 91.21 (84.73, 97.69) 84.63 (77.71, 91.55) 90.94 (84.24, 97.64)
 MRI
@@
 Brain Tumor MRI 96.94 (96.06, 97.79) 99.87 (99.76, 99.95) 99.62 (99.39, 99.79) 99.79 (99.60, 99.92) 99.79 (99.55, 99.95)
-Average 91.00 (90.31,91.67) 96.26 (95.91,96.59) 97.06 (96.73,97.39) 96.66 (96.33,97.01) 98.02 (97.73,98.31)
+Average 91.00 (85.05, 96.94) 96.26 (92.65, 99.87) 97.06 (94.51, 99.62) 96.66 (93.53, 99.79) 98.02 (96.25, 99.79)
 CT
@@
 COVIDxCT 69.56 (69.15, 69.97) 75.14 (74.75, 75.52) 92.98 (92.76, 93.20) 93.46 (93.24, 93.66) 95.31 (95.13, 95.50)
-Average 82.00 (81.01,82.84) 86.04 (85.52,86.55) 95.64 (95.16,96.01) 95.51 (95.06,95.96) 97.12 (96.78,97.42)
-Table 5: AUC scores for classification results across different modalities in the linear probes setting with 10% training data.
-Bold indicates the best result and underline indicates the second best.
-95% CI is included in parentheses.
+Average 82.00 (69.56, 94.44) 86.04 (75.14, 96.95) 95.64 (92.98, 98.30) 95.51 (93.46, 97.57) 97.12 (95.31, 98.94)
+Table 5: AUC scores (%) for classification results across different modalities in the linear probing setting with 10% training data.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
@@
 NIH ChestXray14 63.02 (62.48, 63.57) 68.44 (67.92, 68.92) 70.44 (69.95, 70.86) 71.35 (70.87, 71.82) 75.23 (74.78, 75.69)
-CheXpert 74.91 (74.66,75.18) 80.39 (80.20,80.70) 82.85 (82.60,83.06) 83.17 (82.99,83.36) 86.38 (86.12,86.58)
-Average 67.52 (65.92,68.92) 73.06 (71.97,74.06) 73.04 (71.66,74.15) 74.36 (73.07,75.49) 77.14 (76.12,78.01)
+CheXpert 74.91 (74.67, 75.18) 80.39 (80.20, 80.67) 82.85 (82.68, 83.05) 83.16 (82.99, 83.35) 86.38 (86.18, 86.58)
+Average 67.52 (59.42, 76.02) 73.06 (63.16, 81.90) 73.04 (63.42, 81.45) 74.36 (65.68, 83.05) 77.14 (66.72, 86.77)
 Fundus
@@
 ODIR 65.22 (64.28, 66.22) 68.40 (67.44, 69.43) 68.94 (67.98, 69.81) 68.23 (67.27, 69.23) 72.09 (71.16, 73.06)
-Average 62.98 (45.70,48.54) 65.12 (47.31,50.26) 65.83 (47.82,50.69) 64.14 (46.51,49.51) 67.95 (49.35,52.29)
+Average 62.98 (52.00, 71.71) 65.12 (54.24, 72.71) 65.83 (55.35, 73.21) 64.14 (54.38, 69.81) 67.95 (55.33, 76.44)
 Pathology
@@
 PCam200 83.33 (82.75, 83.88) 89.41 (88.96, 89.85) 91.31 (90.90, 91.70) 91.80 (91.39, 92.19) 96.01 (95.75, 96.27)
-Average 92.11 (91.80,92.42) 95.49 (95.27,95.69) 97.14 (97.00,97.27) 96.96 (96.81,97.11) 98.52 (98.42,98.64)
+Average 92.11 (85.64, 98.58) 95.49 (91.73, 99.00) 97.14 (92.94, 99.71) 96.96 (93.52, 99.32) 98.52 (96.95, 99.85)
 Endoscopy
@@
 Kvasir 95.85 (95.53, 96.17) 96.77 (96.49, 97.05) 97.96 (97.72, 98.17) 97.69 (97.47, 97.93) 98.76 (98.59, 98.93)
-Average 87.88 (65.32,66.49) 90.33 (67.19,68.29) 87.74 (65.24,66.37) 89.75 (66.74,67.88) 92.19 (68.68,69.63)
+Average 87.88 (72.95, 95.85) 90.33 (75.20, 99.02) 87.74 (66.15, 99.12) 89.75 (72.43, 99.13) 92.19 (78.28, 99.54)
 Mammography
@@
 DDSM 92.96 (92.47, 93.46) 92.64 (92.12, 93.17) 95.28 (94.82, 95.68) 91.93 (91.33, 92.49) 93.23 (92.63, 93.75)
-Average 67.90 (65.36,70.04) 69.10 (66.98,71.27) 70.72 (68.26,73.00) 68.11 (65.87,70.43) 70.70 (68.07,72.99)
-Dermoscopy
+Average 67.90 (54.61, 92.96) 69.10 (55.54, 92.64) 70.72 (57.93, 95.28) 68.11 (56.08, 91.93) 70.70 (59.39, 93.23)
+Dermatology
 HAM10000 83.48 (81.85, 85.13) 85.28 (83.84, 86.68) 82.21 (80.28, 83.92) 84.39 (83.04, 85.64) 87.33 (86.22, 88.42)
 PAD-UFES-20 79.24 (77.06, 81.44) 84.40 (82.51, 86.22) 76.90 (74.21, 79.36) 81.24 (79.08, 83.34) 83.67 (81.74, 85.58)
-Average 81.36 (80.02,82.72) 84.84 (83.69,86.05) 79.56 (77.94,81.05) 82.82 (81.57,84.12) 85.50 (84.37,86.70)
+Average 81.36 (79.24, 83.48) 84.84 (84.40, 85.28) 79.56 (76.90, 82.21) 82.82 (81.24, 84.39) 85.50 (83.67, 87.33)
 Ultrasound
@@
 BUID 79.54 (73.49, 85.26) 82.50 (76.92, 87.17) 80.19 (74.46, 85.49) 85.09 (79.95, 89.63) 85.32 (79.54, 90.47)
-Average 75.84 (48.20,52.83) 80.52 (51.60,55.64) 82.07 (52.65,56.74) 83.97 (53.98,57.85) 87.13 (55.95,59.98)
+HMC-DQ 94.25 (93.30, 95.09) 95.45 (94.68, 96.19) 98.04 (97.54, 98.48) 93.07 (92.09, 94.00) 97.94 (97.47, 98.39)
+Average 81.97 (72.13, 94.25) 85.50 (78.54, 95.45) 87.40 (80.19, 98.04) 87.00 (82.85, 93.07) 90.73 (85.32, 97.94)
 OCT
@@
 Retinal OCT 83.66 (83.34, 83.97) 83.92 (83.66, 84.19) 86.17 (85.98, 86.35) 82.63 (82.25, 82.98) 86.06 (85.86, 86.23)
-Average 88.97 (88.51,89.46) 89.44 (88.97,89.88) 92.42 (92.20,92.62) 89.22 (88.81,89.65) 92.62 (92.44,92.81)
+Average 88.97 (83.66, 94.28) 89.44 (83.92, 94.95) 92.42 (86.17, 98.66) 89.22 (82.63, 95.81) 92.62 (86.06, 99.19)
 MRI
@@
 Brain Tumor MRI 97.60 (96.70, 98.34) 99.95 (99.89, 99.99) 99.84 (99.68, 99.95) 99.81 (99.63, 99.95) 99.78 (99.46, 99.96)
-Average 93.70 (93.02,94.34) 98.50 (98.23,98.74) 98.98 (98.76,99.19) 98.08 (97.78,98.37) 99.12 (98.88,99.33)
+Average 93.70 (89.81, 97.60) 98.50 (97.06, 99.95) 98.98 (98.12, 99.84) 98.08 (96.36, 99.81) 99.12 (98.47, 99.78)
 CT
@@
 COVIDxCT 88.09 (87.81, 88.39) 91.40 (91.16, 91.65) 96.82 (96.67, 96.96) 95.30 (95.11, 95.49) 97.35 (97.22, 97.48)
-Average 91.76 (90.94,92.55) 94.98 (94.60,95.29) 97.70 (97.22,98.09) 96.99 (96.67,97.27) 98.40 (98.22,98.57)
-Table 6: AUC scores for classification results across different modalities in the linear probes setting with 100% training data.
-Bold indicates the best result and underline indicates the second best.
-95% CI is included in parentheses.
+Average 91.76 (88.09, 95.44) 94.98 (91.40, 98.55) 97.70 (96.82, 98.58) 96.99 (95.30, 98.68) 98.40 (97.35, 99.46)
+Table 6: AUC scores (%) for classification results across different modalities in the linear probing setting with 100% training data.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
@@
 NIH ChestXray14 61.87 (61.35, 62.42) 67.23 (66.61, 67.79) 71.07 (70.59, 71.56) 70.89 (70.39, 71.41) 74.51 (73.96, 75.08)
-CheXpert 77.93 (77.70,78.20) 82.83 (82.62,83.01) 83.99 (83.84,84.16) 83.73 (83.54,83.92) 87.07 (86.90,87.22)
-Average 70.73 (69.01,72.22) 76.70 (75.00,78.06) 77.29 (75.44,78.96) 77.76 (76.12,79.01) 80.62 (79.14,81.92)
+CheXpert 77.93 (77.70, 78.19) 82.83 (82.62, 83.00) 83.98 (83.84, 84.16) 83.73 (83.55, 83.92) 87.07 (86.90, 87.22)
+Average 70.73 (63.65, 78.41) 76.70 (68.30, 84.52) 77.29 (69.41, 84.34) 77.76 (69.37, 85.47) 80.62 (72.30, 88.36)
 Fundus
@@
 ODIR 72.23 (71.29, 73.14) 74.55 (73.49, 75.54) 76.46 (75.42, 77.52) 73.05 (71.92, 74.13) 77.89 (77.05, 78.88)
-Average 74.63 (71.52,77.53) 76.43 (73.24,79.35) 77.84 (74.66,80.87) 74.57 (71.45,77.52) 78.71 (75.73,81.79)
+Average 74.63 (68.48, 80.76) 76.43 (70.89, 82.18) 77.84 (72.35, 84.60) 74.57 (69.07, 81.55) 78.71 (72.54, 83.36)
 Pathology
@@
 PCam200 87.42 (86.90, 87.92) 92.02 (91.61, 92.42) 93.00 (92.62, 93.35) 92.57 (92.18, 92.95) 96.54 (96.28, 96.78)
-Average 94.79 (94.56,95.01) 97.05 (96.89,97.21) 97.86 (97.75,97.97) 97.50 (97.36,97.64) 98.82 (98.73,98.90)
+Average 94.79 (90.24, 99.32) 97.05 (93.84, 99.59) 97.86 (94.68, 99.86) 97.50 (94.27, 99.68) 98.82 (97.38, 99.94)
 Endoscopy
@@
 Kvasir 98.25 (98.03, 98.47) 98.83 (98.65, 99.01) 99.00 (98.81, 99.17) 98.87 (98.67, 99.05) 99.40 (99.25, 99.52)
-Average 91.60 (90.13,92.75) 92.50 (90.71,93.67) 91.57 (89.88,92.83) 92.21 (90.35,93.51) 93.58 (91.98,94.73)
+Average 91.60 (83.18, 98.08) 92.50 (83.72, 99.26) 91.57 (80.59, 99.36) 92.21 (82.68, 99.25) 93.58 (85.84, 99.61)
 Mammography
@@
 DDSM 95.95 (95.60, 96.32) 96.26 (95.88, 96.62) 96.91 (96.57, 97.21) 94.04 (93.52, 94.50) 96.32 (95.93, 96.70)
-Average 70.76 (68.20,73.15) 71.45 (69.01,73.81) 72.51 (69.98,74.95) 69.77 (67.32,72.11) 74.07 (71.50,76.37)
-Dermoscopy
+Average 70.76 (57.46, 95.95) 71.45 (58.13, 96.26) 72.51 (59.67, 96.91) 69.77 (57.20, 94.04) 74.07 (61.17, 96.32)
+Dermatology
 HAM10000 90.32 (89.11, 91.48) 92.63 (91.70, 93.47) 90.75 (89.66, 91.82) 88.97 (87.78, 90.12) 92.12 (91.22, 92.98)
 PAD-UFES-20 88.16 (86.33, 90.01) 90.09 (88.45, 91.80) 85.47 (83.30, 87.61) 87.33 (85.21, 89.41) 89.90 (88.21, 91.53)
-Average 89.24 (88.13,90.29) 91.36 (90.39,92.30) 88.11 (86.99,89.29) 88.15 (86.96,89.37) 91.01 (90.10,91.88)
+Average 89.24 (88.16, 90.32) 91.36 (90.09, 92.63) 88.11 (85.47, 90.75) 88.15 (87.33, 88.97) 91.01 (89.90, 92.12)
 Ultrasound
@@
 BUID 84.76 (79.68, 89.26) 86.60 (81.51, 91.44) 88.95 (83.99, 93.01) 87.22 (82.40, 91.57) 89.18 (83.99, 93.70)
-Average 75.40 (71.38,79.02) 78.27 (74.47,82.00) 81.03 (77.08,84.80) 78.01 (73.92,81.59) 82.47 (78.48,86.15)
+HMC-DQ 95.76 (95.00, 96.53) 98.65 (98.27, 99.00) 99.43 (99.19, 99.64) 95.50 (94.74, 96.23) 98.70 (98.31, 99.04)
+Average 80.49 (64.49, 92.74) 83.36 (67.02, 95.64) 85.63 (70.07, 96.81) 82.38 (66.57, 93.41) 86.53 (71.95, 96.32)
 OCT
@@
 Retinal OCT 85.31 (85.07, 85.55) 85.61 (85.39, 85.82) 86.64 (86.49, 86.77) 84.55 (84.26, 84.82) 86.73 (86.57, 86.87)
-Average 91.07 (90.69,91.44) 91.84 (91.57,92.08) 92.40 (92.16,92.62) 90.66 (90.23,91.06) 93.10 (92.96,93.23)
+Average 91.07 (85.31, 96.83) 91.84 (85.61, 98.06) 92.40 (86.64, 98.17) 90.66 (84.55, 96.77) 93.10 (86.73, 99.47)
 MRI
@@
 Brain Tumor MRI 98.34 (97.64, 98.86) 99.96 (99.91, 99.99) 99.91 (99.80, 99.98) 99.87 (99.72, 99.97) 99.87 (99.67, 99.99)
-Average 96.32 (95.77,96.83) 99.27 (99.08,99.45) 99.57 (99.43,99.69) 98.90 (98.67,99.12) 99.52 (99.36,99.66)
+Average 96.32 (94.30, 98.34) 99.27 (98.58, 99.96) 99.57 (99.23, 99.91) 98.90 (97.94, 99.87) 99.52 (99.17, 99.87)
 CT
@@
 COVIDxCT 93.87 (93.67, 94.07) 96.46 (96.31, 96.62) 97.59 (97.47, 97.71) 96.14 (95.98, 96.31) 98.11 (97.99, 98.22)
-Average 95.46 (94.79,96.03) 97.79 (97.56,98.01) 98.22 (97.82,98.56) 97.64 (97.40,97.84) 98.86 (98.71,98.99)
-Table 7: AUC scores for classification results across different modalities in the fully fine-tuning setting.
-Bold indicates the best result and underline indicates the second best.
-95% CI is included in parentheses.
+Average 95.46 (93.87, 97.05) 97.79 (96.46, 99.13) 98.22 (97.59, 98.85) 97.64 (96.14, 99.13) 98.86 (98.11, 99.61)
+Table 7: AUC scores (%) for classification results across different modalities in the fully fine-tuning setting.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Dataset CLIP SigLIP-400M PMC-CLIP BiomedCLIP ConceptCLIP
@@
 NIH ChestXray14 71.21 (70.68, 71.75) 78.91 (78.43, 79.35) 73.17 (72.68, 73.62) 70.67 (70.18, 71.17) 78.99 (78.51, 79.43)
-CheXpert 89.29 (89.14,89.42) 89.34 (89.17,89.49) 87.85 (87.67,88.06) 87.56 (87.38,87.75) 89.16 (88.97,89.35)
-Average 75.06 (73.43,76.46) 79.47 (77.85,80.81) 76.62 (75.21,77.89) 78.26 (76.77,79.69) 82.76 (81.29,83.99)
+CheXpert 89.29 (89.14, 89.42) 89.34 (89.17, 89.49) 87.86 (87.67, 88.08) 87.56 (87.40, 87.75) 89.16 (88.97, 89.35)
+Average 75.06 (73.42, 76.45) 79.47 (77.84, 80.81) 76.62 (75.20, 77.90) 78.26 (76.78, 79.67) 82.76 (81.28, 84.01)
 Fundus
-RFMiD2 60.81 (51.87,68.51) 58.34 (50.65,64.88) 63.39 (54.47,71.84) 63.42 (54.52,71.52) 69.09 (60.30,77.22)
+RFMiD2 60.81 (51.87, 68.51) 58.34 (50.65, 64.88) 63.39 (54.47, 71.84) 62.54 (54.55, 69.84) 69.09 (60.30, 77.22)
 DRD 71.44 (70.84, 72.00) 83.51 (83.15, 83.89) 82.55 (82.15, 82.94) 77.31 (76.84, 77.74) 80.35 (79.95, 80.72)
@@
 ODIR 76.77 (75.72, 77.86) 81.24 (80.29, 82.14) 81.07 (80.07, 82.06) 79.08 (78.02, 80.05) 83.37 (82.52, 84.28)
-Average 74.50 (71.18,77.17) 77.71 (74.71,80.47) 78.82 (75.75,81.74) 77.13 (73.99,80.13) 80.33 (77.28,83.03)
+Average 74.50 (71.18, 77.17) 77.71 (74.71, 80.47) 78.82 (75.75, 81.74) 76.91 (73.86, 79.76) 80.33 (77.28, 83.03)
 Pathology
@@
 Average 73.54 (71.40, 75.88) 81.21 (79.00, 83.25) 79.46 (77.19, 81.49) 78.04 (75.68, 80.04) 81.71 (79.44, 83.47)
-Dermoscopy
+Dermatology
 HAM10000 93.60 (92.07, 94.88) 96.97 (95.83, 97.86) 95.39 (94.57, 96.14) 94.79 (93.52, 95.86) 96.47 (95.75, 97.14)
@@
 BUID 95.00 (91.95, 97.58) 94.21 (90.90, 96.68) 93.44 (89.84, 96.68) 95.91 (93.60, 97.73) 94.97 (92.37, 97.19)
-Average 88.52 (84.28,92.28) 87.01 (83.11,90.69) 88.52 (84.39,92.52) 88.80 (84.37,92.55) 89.01 (85.18,92.95)
+HMC-DQ 97.44 (96.91, 97.94) 99.35 (99.09, 99.58) 99.63 (99.39, 99.82) 99.87 (99.73, 99.95) 99.74 (99.59, 99.86)
+Average 90.75 (87.55, 93.63) 90.09 (87.20, 92.87) 91.30 (88.20, 94.30) 91.57 (88.25, 94.38) 91.70 (88.82, 94.65)
 OCT
@@
 Average 98.56 (98.49, 98.63) 98.50 (98.42, 98.57) 99.40 (99.33, 99.46) 98.56 (98.50, 98.62) 99.65 (99.62, 99.68)
-Table 8: Performance of Recall@1,5,10 metrics of different models on the PMC-9K dataset. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
+Table 8: Recall@1, 5, 10 (%) of different models for cross-modal retrieval on the PMC-9K dataset.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Image-to-Text
@@
 CLIP 5.63 (5.18, 6.10) 12.47 (11.80, 13.15) 17.14 (16.38, 17.92)
-SigLIP-B 3.38 (3.04,3.77) 8.25 (7.69,8.79) 11.66 (10.98,12.29)
 SigLIP-400M 2.86 (2.54, 3.23) 7.04 (6.53, 7.59) 9.68 (9.09, 10.37)
@@
 CLIP 6.31 (5.79, 6.83) 13.11 (12.44, 13.86) 18.39 (17.63, 19.24)
-SigLIP-B 3.29 (2.94,3.64) 8.06 (7.54,8.57) 11.68 (11.04,12.31)
 SigLIP-400M 2.65 (2.32, 2.97) 6.96 (6.47, 7.47) 9.82 (9.22, 10.40)
@@
 ConceptCLIP 83.24 (82.51, 83.96) 94.50 (94.05, 94.99) 96.86 (96.50, 97.21)
-Table 9: Performance of Recall@1,50,200 metrics of different models on the QUILT-1M dataset. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
+Table 9: Recall@1, 50, 200 (%) of different models for cross-modal retrieval on the QUILT-1M dataset.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Image-to-Text
@@
 CLIP 0.62 (0.42, 0.58) 4.14 (3.88, 4.56) 8.37 (7.95, 8.84)
-SigLIP-B 0.32 (0.21,0.32) 3.24 (2.90,3.51) 7.56 (7.10,8.00)
 SigLIP-400M 0.28 (0.19, 0.30) 3.82 (3.53, 4.15) 7.98 (7.57, 8.46)
@@
 CLIP 0.74 (0.48, 0.65) 5.18 (4.80, 5.49) 9.40 (8.92, 9.90)
-SigLIP-B 0.42 (0.27,0.40) 4.07 (3.68,4.35) 8.76 (8.26,9.22)
 SigLIP-400M 0.36 (0.22, 0.35) 4.05 (3.72, 4.39) 9.24 (8.76, 9.80)
@@
 ConceptCLIP 1.95 (1.38, 1.67) 17.07 (16.44, 17.73) 32.50 (31.79, 33.45)
-Table 10: Results of models on the medical visual question answering task. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
+Table 10: Results (%) on the medical visual question answering task. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
 Metric CLIP BiomedCLIP PMC-CLIP SigLIP-400M ConceptCLIP
@@
 Overall Accuracy 68.71 (64.52, 72.73) 68.21 (63.63, 72.51) 70.50 (66.73, 74.73) 66.73 (62.53, 70.96) 70.70 (66.73, 74.73)
-Table 11: Results of models on the medical report generation task. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
+Table 11: Results (%) on the medical report generation task. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
 Metric CLIP BiomedCLIP PMC-CLIP SigLIP-400M ConceptCLIP
@@
 METEOR 15.41 (15.49, 15.84) 15.19 (15.03, 15.36) 14.02 (13.88, 14.18) 14.19 (14.03, 14.36) 15.66 (15.49, 15.84)
-CIDEr 76.80 (76.32,82.23) 73.79 (71.25,76.45) 63.21 (60.96,65.54) 74.11 (71.57,76.62) 79.26 (76.32,82.23)
 Micro Precision 28.93 (30.80, 33.35) 26.08 (24.92, 27.30) 19.61 (18.44, 20.79) 15.05 (14.04, 16.19) 32.09 (30.80, 33.35)
@@
 IU X-Ray
-BLEU-1 49.20 46.06 20.63 45.91 49.74
-BLEU-2 32.32 29.51 9.90 29.10 32.57
-BLEU-3 23.26 21.07 5.54 20.58 23.60
-BLEU-4 17.37 15.74 3.19 15.44 17.82
-ROUGE-L 37.64 37.40 16.98 35.95 39.06
-METEOR 20.46 21.27 12.69 20.04 21.16
-CIDEr 46.58 44.88 0.62 47.02 42.56
-Micro Precision 54.70 55.71 19.52 51.36 51.84
-Micro Recall 50.48 54.53 26.40 39.88 57.03
-Micro F1 52.51 55.11 22.44 44.90 54.31
-Table 12: Results of the models on various datasets for Whole-Slide image tasks.
-We use C-index for survival prediction tasks, for other tasks, we use AUC scores.
-Bold indicates the best result and underline indicates the second best.
-Mean±std is presented.
-Dataset CLIP BiomedCLIP PMC-CLIP SigLIP-400M PLIP PathGen-CLIP ConceptCLIP
+BLEU-1 44.11 (43.48, 45.64) 43.59 (42.59, 44.61) 20.31 (19.59, 21.00) 41.97 (40.89, 43.04) 44.58 (43.48, 45.64)
+BLEU-2 27.95 (27.31, 29.82) 26.79 (25.64, 28.01) 8.75 (8.22, 9.30) 25.43 (24.23, 26.59) 28.60 (27.31, 29.82)
+BLEU-3 19.08 (18.13, 20.66) 17.93 (16.86, 19.06) 3.97 (3.57, 4.37) 16.24 (14.94, 17.64) 19.38 (18.13, 20.66)
+BLEU-4 12.57 (12.07, 14.51) 12.11 (11.09, 13.19) 1.61 (1.36, 1.87) 10.72 (9.59, 11.95) 13.30 (12.07, 14.51)
+ROUGE-L 37.65 (37.39, 39.68) 37.37 (36.36, 38.45) 16.98 (16.39, 17.54) 35.97 (34.89, 37.06) 38.52 (37.39, 39.68)
+METEOR 21.46 (20.81, 22.13) 22.06 (21.85, 23.31) 12.89 (12.42, 13.33) 21.14 (20.38, 21.93) 22.60 (21.85, 23.31)
+Micro Precision 44.32 (40.17, 48.31) 47.74 (40.93, 48.81) 22.31 (18.94, 25.63) 35.01 (31.02, 38.81) 44.84 (40.93, 48.81)
+Micro Recall 44.33 (40.17, 48.31) 47.24 (40.65, 48.36) 22.48 (19.24, 25.93) 34.75 (30.93, 38.39) 44.49 (40.65, 48.36)
+Micro F1 44.26 (40.28, 48.36) 47.35 (40.68, 48.73) 22.29 (18.95, 25.62) 34.82 (31.02, 38.59) 44.52 (40.68, 48.73)
+Table 12: Results (%) on various datasets for pathology whole-slide image tasks.
+We use the C-index for survival prediction tasks.
+For other tasks, we use AUC scores.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
+Dataset CLIP BiomedCLIP PMC-CLIP SigLIP-400M ConceptCLIP
 Cancer Diagnosis
-BRACS-3 80.51±3.51 90.0±2.35 87.34±2.89 87.84±2.52 89.57±2.48 90.2±2.49 91.65±2.28
-BRACS-7 72.68±3.27 81.43±2.68 77.29±3.07 80.85±2.78 82.57±2.54 84.97±2.11 85.04±2.18
-BRCA 89.94±2.18 91.03±2.48 87.44±3.04 91.16±2.41 86.37±3.49 93.72±1.93 92.98±2.14
-NSCLC 90.84±1.69 92.89±1.43 93.21±1.42 94.05±1.27 93.47±1.47 95.5±1.12 95.01±1.15
-Camelyon 86.69±3.25 89.86±2.75 93.98±1.86 90.15±2.81 92.35±2.46 94.69±1.83 95.17±1.76
-Molecular Subtyping
-BRCA 70.23±2.93 72.15±2.96 71.75±3.09 70.92±2.64 73.93±3.05 74.04±2.81 74.36±3.08
+BRACS-3 80.51 (73.63, 87.39) 90.00 (85.39, 94.61) 87.34 (81.68, 92.99) 87.84 (82.90, 92.78) 91.65 (87.18, 96.12)
+BRACS-7 72.68 (66.27, 79.09) 81.43 (76.18, 86.68) 77.29 (70.27, 84.31) 80.85 (75.40, 86.30) 85.04 (80.77, 89.31)
+NSCLC 90.84 (87.53, 94.15) 92.89 (89.09, 96.69) 93.21 (90.43, 96.00) 94.05 (91.56, 96.54) 95.01 (92.76, 97.26)
+Camelyon 86.69 (80.32, 93.06) 89.86 (83.47, 96.25) 93.98 (90.33, 97.63) 90.15 (84.63, 95.67) 95.17 (91.72, 98.62)
+Mutation Prediction
+BRCA-TP53 60.83 (50.52, 71.14) 76.23 (67.21, 85.25) 66.89 (56.89, 76.89) 77.95 (69.19, 86.71) 80.15 (72.05, 88.25)
+BRCA-TTN 42.74 (29.76, 55.72) 39.75 (23.26, 56.24) 40.03 (24.98, 55.08) 42.35 (26.77, 57.93) 52.14 (35.18, 69.10)
+LUAD-TP53 45.61 (33.46, 57.76) 66.42 (55.35, 77.49) 60.96 (48.67, 73.25) 58.55 (46.69, 70.41) 68.02 (56.85, 79.19)
 Survival Prediction
-BRCA 59.16±5.72 62.5±5.76 57.22±5.78 62.49±5.15 59.45±5.75 66.35±4.41 64.59±5.24
-LUAD 63.48±5.04 62.74±4.5 58.12±4.76 62.78±4.82 60.27±4.79 59.99±4.81 59.75±5.28
-LUSC 58.4±4.25 55.98±4.63 55.5±4.27 56.86±4.82 54.83±4.43 56.88±4.43 59.66±4.77
+BRCA 59.16 (47.95, 70.37) 62.50 (51.21, 73.79) 57.22 (46.00, 68.44) 62.49 (52.40, 72.58) 64.59 (54.32, 74.86)
+LUSC 58.40 (50.07, 66.73) 55.98 (46.90, 65.06) 55.50 (47.13, 63.87) 56.86 (47.40, 66.32) 59.66 (50.31, 69.01)
 Table 13: Performance of medical vision-language models on zero-shot medical concept annotation tasks (AUC %).
-“w/o Local Info.” denotes that the local information of image patches is not used in zero-shot concept annnotation.
+“w/o Local Info.” denotes that the local information of image regions is not used in zero-shot concept annnotation.
 Bold indicates the best result and underline indicates the second best.
 95% CI is included in parentheses.
-Model Derm7pt (Dermoscopy) SkinCon (Dermoscopy) WBCAtt (Hematology) BrEaST (Ultrasound) LUNA16 (CT) Average
+Model Derm7pt (Dermatology) SkinCon (Dermatology) WBCAtt (Hematology) BrEaST (Ultrasound) LUNA16 (CT) Average
 CLIP 53.59 (51.66, 55.32) 63.88 (62.41, 65.47) 56.27 (55.61, 56.96) 45.30 (41.47, 48.97) 51.64 (49.89, 53.30) 54.14
 SigLIP-400M 57.80 (55.70, 59.81) 67.75 (66.15, 69.62) 56.77 (56.14, 57.39) 51.98 (48.17, 55.68) 52.54 (50.80, 54.03) 57.37
-MONET 66.16 (64.24, 68.13) 67.30 (65.71, 68.90) - - - -
 PMC-CLIP 66.91 (65.09, 68.81) 61.19 (59.35, 63.06) 54.93 (54.13, 55.76) 56.73 (52.88, 60.42) 54.84 (53.10, 56.54) 58.92
 BiomedCLIP 65.28 (63.17, 67.33) 68.88 (67.20, 70.57) 52.44 (51.71, 53.20) 54.25 (50.28, 58.08) 54.99 (53.47, 56.52) 59.17
-ConceptCLIP w/o Local Info. 67.62 70.62 58.97 64.92 58.42 64.11
 ConceptCLIP 68.56 (66.53, 70.45) 72.20 (70.70, 73.76) 60.59 (59.91, 61.34) 66.24 (61.81, 70.07) 59.04 (57.43, 60.70) 65.33
@@
 ConceptCLIP 80.20 (75.37, 84.79) 99.49 (99.33, 99.62) 83.23 (70.90, 92.95) 68.63 (62.71, 74.42) 82.89
-Table 15: Ablation study on the effectiveness of the PC-Align loss and the local information (“Local Info.”).
-Bold indicates the best result and underline indicates the second best.
-95% CI is included in parentheses.
+Table 15: Ablation study on the effectiveness of the RC-Align loss and the local information (“Local Info.”) using the AUC (%) score.
+Bold indicates the best results and underline indicates the second best.
+The 95% CI is included in parentheses.
 Model SIIM-ACR Covid-CXR2 VinDr-Mammo Brain Tumor CT
-ConceptCLIP w/o PC-Align 80.86 (77.87,83.83) 80.07 (78.16,81.77) 47.72 (43.73,51.18) 84.90 (82.42,87.17)
+ConceptCLIP w/o RC-Align 80.86 (77.87, 83.83) 80.07 (78.16, 81.77) 47.72 (43.73, 51.18) 84.90 (82.42, 87.17)
 ConceptCLIP w/o Local Info. 81.24 (78.54, 84.04) 79.67 (77.98, 81.51) 50.75 (47.41, 54.19) 90.43 (88.35, 92.28)
 ConceptCLIP 83.05 (80.40, 85.63) 81.77 (80.19, 83.35) 51.78 (48.25, 55.05) 92.60 (90.76, 94.20)
-Table 16: Medical image analysis dataset splits and evaluation metrics.
-Task Dataset Split Metrics
+Table 16: Medical image analysis dataset splits (training/validation/testing) and evaluation metrics.
+Task Dataset Train/Valid/Test Split Metrics
 Binary Classification SIIM-ACR [51] $2587$ / $862$ / $863$ AUC
@@
 Breast Cancer [61] $2372$ / $675$ / $336$
+HMC-QU [62] $7710$ / $956$ / $2267$
 Multi-Label Classification RFMiD2 [63] $342$ / $113$ / $149$ AUC
@@
 BRACS-7 [86] $382$ / $109$ / $54$
-BRCA [94] $716$ / $102$ / $307$
 NSCLC [87] $664$ / $100$ / $289$
 Camelyon [56, 88] $630$ / $91$ / $180$
-Molecular Subtyping BRCA [94] $716$ / $102$ / $307$ AUC
+Mutation Prediction BRCA-TP53 [87] $438$ / $68$ / $128$ AUC
+BRCA-TTN [87] $438$ / $68$ / $128$
+LUAD-TP53 [87] $313$ / $44$ / $91$
 Survival Prediction BRCA [87] $716$ / $102$ / $307$ C-Index
-LUAD [94] $318$ / $45$ / $92$
 LUSC [87] $316$ / $45$ / $91$
+Table 17: Medical concept annotation datasets.
+Dataset Description and setting Link
+Derm7pt [12] A dermoscopic image dataset containing 1,011 images with clinical concepts for melanoma skin lesions in dermatology.
+The test set is adopted for evaluation.
+Only the dermoscopic images are considered in this paper for concept annotations.
+The seven clinical concepts include “Pigment network”, “Dots and globules”, “Pigmentation”, “Streaks”, “Regression structures”, “Blue-whitish veil”, and “Vascular structures”.
+SkinCon [10] A skin disease dataset densely annotated by experts for fine-grained model debugging and analysis.
+We choose 22 concepts that have at least 50 images representing the concept from the Fitzpatrick 17k (F17k) [89] subset.
+350 images are sampled for evaluation.
+The clinical concepts include “Papule”, “Plaque”, “Pustule”, “Bulla”, “Patch”, “Nodule”, “Ulcer”, “Crust”, “Erosion”, “Atrophy”, “Exudate”, “Telangiectasia”, “Scale”, “Scar”, “Friable”, “Dome-shaped”, “Brown (Hyperpigmentation)”, “White (Hypopigmentation)”, “Purple”, “Yellow”, “Black”, and “Erythema”.
+WBCAtt [11] This dataset is a densely annotated dataset for white blood cell (WBC) recognition, containing 11 morphological attributes for 10,298 cell images.
+The test set which contains 3,099 images is used for evaluation.
+There are 11 concepts, including “Cell size”, “Cell shape”, “Nucleus shape”, “Nuclear cytoplasmic ratio”,“Chromatin density”, “Cytoplasm vacuole”, “Cytoplasm texture”, “Cytoplasm colour”, “Granule type”, “Granule colour”, and “Granularity”.
+BrEaST [90] A breast ultrasound dataset containing 256 images with fine-grained concepts for breast lesion analysis.
+Seven BI-RADS descriptors are adopted as the evaluated concepts, including “Shape”, “Magrin”, “Echogenicity”, “Posterior features”, “Halo”, “Calcifications”, and “Skin thickening”.
+LUNA16 [13] A curated version of LIDC-IDRI [91] dataset, which contains lung images with clinical concepts.
+1,185 slices of lesions from the chest CTs with six clinical concepts are used for evaluation.
+The concepts include “Calcification”, “Lobulation”, “Margin”, “Sphericity”, “Spiculation”, and “Texture”.
+Table 18: State-of-the-art vision-language pre-training models.
+Model Name Description Link
+Models in the General Domain
+CLIP [34] CLIP [34] is a vision-language model pre-trained on large-scale image-text pairs via contrastive learning in the general domain, which showcases strong ability in zero-shot classification.
+We take the version of CLIP-ViT-B-16 pre-trained on LAION-2B [92] under the framework of OpenCLIP [93].
+SigLIP-400M [28] SigLIP [28] is a variant of the CLIP model, which is pre-trained via contrastive learning using the sigmoid loss.
+In our experiments, we use the version of ViT-SO400M-14-SigLIP (SigLIP-400M).
+Models in the Medical Domain
+PMC-CLIP [3] PMC-CLIP [3] is a vision-language model tailored for medicine.
+The authors collected PubMed open access articles and conducted a series of data processing operations, producing a dataset with 1.6 million medical image-text pairs.
+The dataset is used to pre-train a CLIP model, thus resulting in the PMC-CLIP model.
+We use the origin checkpoint that the authors provided.
+BiomedCLIP [2] BiomedCLIP [2] is a medical vision-language model pre-trained on PMC-15M, a dataset with 15 million medical image-text pairs sourced from PMC-OA.
+We take the released checkpoint for evaluation.